### v1

# Poincare field theory for massless particles

Abstract:
Our main proposition is that field equations for all spins can be obtained from Casimir eigenvalue equations for Poincare group. We have already confirm that statement for massive scalar, spinor and vector fields in Ref.[1]. In the present article we are going to confirm this statement for massless vector, second and fourth rang tensor fields. In particular we will obtain Maxwell equations and Einstein equation in weak field approximation.
As is well known, Wigner define a particles as irreducible representation of Poincare group [2, 3]. But, as Weinberg noted in [4] irreducible representations for massless vector field with helicities ±$\pm 1$ do not exist. In the present article we will conform this statement for a wide class of massless fields. They are not Lorentz invariant since their Lorentz transformations have additional term in the form of gauge transformations. Such fields can appear in the theory only in the form which do not depend on corresponding gauge parameters. These forms are our equations of motion and they are by definition gauge invariant. So, the massless case is significantly different from massive one.
1 Introduction
In the previous article [1] we introduced equations of motion for massive particles with arbitrary spin as eigenvalue equations for Casimir operators of Poincare group. Here we will develop appropriate approach for massless particles. Although in massless case there is one Casimir operator (helicity), covariant form of its definition contains Pauli-Lubanski vector $W_{a}$ which has three independent components ($P^{a}W_{a}=0$). So, we have three equations and their compatibility produces significant differences from massive case.
As in the massive case the fields are functions which transforms with some representation of Poincare group. For arbitrary representations we will introduce notation $\Psi^{A}(x)$ where $A$ contains the set of corresponding indices.
We will work in the frame of standard momentum which remarkably simplifies all considerations. First we will solve eigenvalue problem for component of spin operator $(S_{12})^{A}{}_{B}=\sigma(W_{0})^{A}{}_{B}=(W_{3})^{A}{}_{B}$. Corresponding eigenvalues we can interpret as helicities $\lambda_{i}$ and eigenvectors $\Psi^{A}_{i}$ as potential irreducible representations of little group. But in some important cases components of Pauli-Lubanski vector do not commute. So, there is no common eigenvectors for all components. Even more, components $W_{1}$ and $W_{2}$ in the frame of standard momentum do not annihilate eigenvectors $\Psi^{A}_{i}$, which means that $\Psi^{A}_{i}$ are not representations of Poincare group.
This feature has unexpected consequences discovered by Weinberg [4] that Lorentz transformations of some physically relevant fields (like vector and symmetric second rang tensor fields) have additional term in the form of gauge transformation. So, for these fields we can obtain actions as gauge invariant expressions of eigenvectors $\Psi^{A}_{i}$.
Therefore, gauge dependent fields are not Lorentz invariant and so they are not representation of Poincare group. But corresponding field strength and Lagrangian (gauge invariant expression of gauge fields) are representations of Poincare group. So, for gauge fields we will introduce name incomplete representation of Poincare group (up to gauge transformations) and for corresponding irreducible part incomplete irreducible representation of Poincare group.
As a consequences, Wigner’s definition of particles [2, 3] as irreducible representations of Poincare group, which valid in massive case, in massless case must have an improvement. We can define physical massless fields as a class of equivalence, where two fields are equivalent if they relate by gauge transformation. The action is expression which depend on class of equivalence.
2 Principle field equations
In this section we will introduce principle field equations. They are defined only on the basis of symmetry group of the theory and they are formulated with the help of Casimir invariants. In this article we start with massless Poincare group and introduce principle field equations which valid for arbitrary helicity.
2.1 Representation for Pauli-Lubanski vector
Casimir invariants are expressions that commute with all group elements. In massless case, where square of momentum vanishes $P^{2}=0$, there are two Casimir operators for Poincare group, the helicity $\lambda$ and sign of $P^{0}$. The covariant definition of helicity is
$W_{a}=\lambda P_{a}\,,\qquad W_{a}=\frac{1}{2}\varepsilon_{abcd}S^{bc}P^{d}\,,$ (1)
where $W_{a}$ is Pauli-Lubanski vector and $S_{ab}$ is spin parts of Lorentz generators. Since $P^{a}W_{a}=0$ we can conclude from (1) that $P^{2}=0$, as it should be for massless case.
Two comments should be made. First, it is possible to use non covariant definition of helicity $\lambda=S_{i}n_{i}$ as projection of space part of spin generator $S_{i}=\frac{1}{2}\,\varepsilon_{ijk}S_{jk}$ to momentum axis $n_{i}=\frac{p_{i}}{|p_{i}|^{2}}$. It produces the same spectrum for $\lambda$ as relation (1). But in that case we are losing the crucial possibility to insist on Lorentz invariance which will product gauge transformations. Second, in definition of Pauli-Lubanski vector orbital part of four dimensional rotation generatr $L_{ab}=x_{a}P_{b}-x_{b}P_{a}$ does not appear since $\varepsilon_{abcd}L^{bc}P^{d}=0$.
For all types of fields representation of momentum is $(P_{a})^{A}{}_{B}\to\delta^{A}_{B}\,i\partial_{a}$. In article [1] we proved relation
$(S_{ab})^{AB}{}_{CD}=(S_{ab})^{A}{}_{C}\delta_{D}^{B}+\delta^{A}_{C}(S_{ab})^{B}{}_{D}\,,$ (2)
which allows us to find representations for arbitrary spin operator using fundamental representations. This initial representation for Dirac spinors and vector fields are
$(S_{ab})^{\alpha}{}_{\beta}=\frac{i}{4}[\gamma_{a},\gamma_{b}]^{\alpha}{}_{\beta}\,,\qquad(S_{ab})^{c}{}_{d}=i\Big{(}\delta^{c}_{a}\eta_{bd}-\delta^{c}_{b}\eta_{ad}\Big{)}\,.$ (3)
For example, the expression for vectors in terms of spinors $V^{a}(x)={\bar{\psi}}(x)\gamma^{a}\psi(x)$ and first relation (3) produces the second one, which can also been obtained by direct calculation.
Using representations for momenta and spin generators and the fact that they commute, we can find representation for Pauli-Lubanski vector.
2.2 Standard momentum $k^{a}$, space inversion and little group
It is useful to work in frame of standard momentum $k^{a}$. In massless case, when Lorentz invariant function of momentum vanish, $p^{2}=0$, we can chose $k^{a}=(1,0,0,-1)$. Then we can express any momentum $p^{a}$ as Lorentz transformation of $k^{a}$
$p^{a}=L^{a}{}_{b}(p)k^{b}\,.$ (4)
Up to now we considered proper and orthochronous Lorentz group $\Lambda\in SO(1,3)^{\uparrow}$ which contains the identity matrix and preserves the direction of time and parity ( $\det\Lambda=1$ and $\Lambda^{0}{}_{0}\geq 1$ ). But we will also need space inversion ( $\det\Lambda=-1$ and $\Lambda^{0}{}_{0}\geq 1$ ) which for vector fields is implemented by the operator
${\cal P}^{a}{}_{b}=\delta^{a}_{0}\delta_{b}^{0}-\delta^{a}_{i}\delta_{b}^{i}\,.\qquad(i=1,2,3)$ (5)
In order to follow behavior under space inversion we will introduce new parameter $\sigma$ with inversion rules ${\cal P}\sigma=-\sigma$, and obviously $\sigma^{2}=1$. Then rotations are invariant under parity $S_{ij}\to S_{ij}$ and boosts change the sign $S_{0i}\to\sigma S_{0i}$ while $k^{a}\to(1,0,0,-\sigma)$.
Since $P^{a}W_{a}=0$, Pauli-Lubanski vector has three components and for standard momentum they are
$W_{0}=\sigma S_{12}\,,\qquad W_{3}=S_{12}\,,\qquad$
$W_{1}\equiv\Pi_{2}=S_{02}-S_{32}\,,\qquad-W_{2}\equiv\Pi_{1}=S_{01}-S_{31}\,.$ (6)
All three generators annihilate standard momentum. So group element $W^{a}{}_{b}$, leaves $k^{a}$ invariant $W^{a}{}_{b}k^{b}=k^{a}$, which is definition of little group in massless Poincare case.
Commutation relations of little group generators have a form
$[\Pi_{1},\Pi_{2}]=0\,,\qquad[S_{12},\Pi_{1}]=i\Pi_{2}\,,\qquad[S_{12},\Pi_{2}]=-i\Pi_{1}\,.$ (7)
Therefore, as it is well known, in massless case the little group of Poincare group is group $E(2)$, which consists of translations and rotations in two-dimensional Euclidean space.
Note that unlike the massive case here little group contains Abelian subgroup with generators $\Pi_{1}$ and $\Pi_{2}$. Representation of little group must be invariant under this Abelian subgroup. We will see that violation of this feature has important repercussion that gauge invariance is consequence of Lorentz transformation [4].
2.3 Principle field equations
With the help of expression (1) and representation of momenta and Pauli-Lubanski vector we can postulate principle field equations for massless fields $\Psi^{A}(x)$ with helicity $\lambda$
$(W_{a})^{A}{}_{B}\Psi^{B}(x)=\lambda(P_{a})^{A}{}_{B}\Psi^{B}(x)\,.$ (8)
We claim that this equations contains all massless free field equations with arbitrary helicity, known in classical field theory.
We will investigate principle field equations in momentum representation in the frame of standard momentum. Then differential equations (8) become algebraic equations. After solving this algebraic equations we can go back to $p^{a}$ dependent solutions using (4) and then to coordinate representation.
Using (8) and (2.2) we can obtain principle field equations for standard momentum
$(S_{12})^{A}{}_{B}\Psi^{B}(k)=\lambda\sigma\Psi^{A}(k)\,,$
$(\Pi_{1})^{A}{}_{B}\Psi^{B}(k)=0\,,\qquad(\Pi_{2})^{A}{}_{B}\Psi^{B}(k)=0\,.$ (9)
Let us stress that all massless free field equations considered in classical field theory [4, 5, 6, 7, 8] can be obtained from Eqs.(2.3). In the present article we will confirm this statement for most important massless fields.
2.4 Consequences of parity conservation
The operators $(S_{12})^{A}{}_{B}$, $(\Pi_{1})^{A}{}_{B}$ and $(\Pi_{2})^{A}{}_{B}$ are representation of little group generators and $\lambda\sigma$ is a eigenvalue of operator $(S_{12})^{A}{}_{B}$. Under space inversion operator $(S_{12})^{A}{}_{B}$ is invariant which means that helicity change the sign
${\cal P}\lambda=-\lambda\,.$ (10)
Therefore, if the theory is invariant under space inversion then for any non zero helicity it must exist opposite one.
Helicity is invariant under proper orthochronous Poincare group. Particles of opposite helicities $+\lambda$ and $-\lambda$, which are related by transformation of space inversion, corresponds to different species of particles. A theory that conserves parity must treat both helicities states equally. So, both massless states with opposite helicities we can consider as two polarizations of the same particle. This particle is irreducible representation of wider group which combined Poincare group and parity.
The most important examples are electromagnetic and gravitation interactions which are invariant under space inversion. In particular, both massless particles with helicities $\lambda=1$ and $\lambda=-1$ are called photons and both massless particles with helicities $\lambda=2$ and $\lambda=-2$ are called gravitons. General state of massless particle which includes both helicities $\lambda$ and $-\lambda$ is
$\Psi^{A}(x)=\alpha\Psi^{A}_{+\lambda}(x)+\beta\Psi^{A}_{-\lambda}(x)\,,\qquad|\alpha|^{2}+|\beta|^{2}=1\,,$ (11)
and in momentum representation
$\Psi^{A}(p)=\alpha\Psi^{A}_{+\lambda}(p)+\beta\Psi^{A}_{-\lambda}(p)\,.$ (12)
2.5 Real fields in coordinate and momentum representation
In most cases we will work with real fields, such that in coordinate and momentum representation
$(\Psi^{A})^{\ast}(x)=\Psi^{A}(x)\,,\qquad(\Psi^{A})^{\ast}(p)=\Psi^{A}(-p)\,.$ (13)
So for real fields
$\Psi^{A}(x)=\int d^{4}p\Big{(}\Psi^{A}(p)+\Psi^{A}(-p)\Big{)}e^{-ipx}\,.$ (14)
Similarly, for imaginary field we have
$(\Psi^{A})^{\ast}(x)=-\Psi^{A}(x)\,,\qquad(\Psi^{A})^{\ast}(p)=-\Psi^{A}(-p)\,.$ (15)
We will need real part of the expression which in momentum representation has a form $\Psi_{n}^{B}(p)=p^{a_{1}}p^{a_{2}}\cdots p^{a_{n}}\Psi^{B}(p)$
$\Psi_{n}^{B}(x)=\int d^{4}p\Big{[}\Psi_{n}^{B}(p)+\Psi_{n}^{B}(-p)\Big{]}e^{-ipx}=(i)^{n}\partial^{a_{1}}\partial^{a_{2}}\cdots\partial^{a_{n}}\int d^{4}p\Big{[}\Psi^{B}(p)+(-1)^{n}\Psi^{B}(-p)\Big{]}e^{-ipx}\,.\qquad$ (16)
For $n=1$ we have
$\Psi_{1}^{B}(x)=\int d^{4}p\Big{[}p^{a}\Psi^{B}(p)-p^{a}\Psi^{B}(-p)\Big{]}e^{-ipx}=\partial^{a}\Omega_{1}^{B}(x)\,,\qquad\qquad$ (17)
and for $n=2$
$\Psi_{2}^{B}(x)=\int d^{4}p\Big{[}p^{a}p^{b}\Psi^{B}(p)+p^{a}p^{b}\Psi^{B}(-p)\Big{]}e^{-ipx}=\partial^{a}\partial^{b}\Omega_{2}^{B}(x)\,,$ (18)
where
$\Omega_{1}^{B}(x)=i\int d^{4}p\Big{[}\Psi^{B}(p)-\Psi^{B}(-p)\Big{]}e^{-ipx}\,,\qquad\Omega_{2}^{B}(x)=-\int d^{4}p\Big{[}\Psi^{B}(p)+\Psi^{B}(-p)\Big{]}e^{-ipx}\,.\qquad\qquad$ (19)
According to (15) and (13) both $\Omega_{1}^{B}(x)$ and $\Omega_{2}^{B}(x)$ are real functions.
2.6 Solution of $(S_{12})^{A}{}_{B}$ equation
We are going to solve eigenproblem of operator $(S_{12})^{A}{}_{B}$ which will produces spectrum of helicities $\lambda_{i}$ $(i=1,2,\cdots,n)$. Then we will find projection operators $(P_{i})^{A}{}_{B}$ and eigenfunctions $\Psi_{i}^{A}$ corresponding to these eigenvalues.
In order that first equation (2.3) has nontrivial solutions its characteristic polynomial must vanish
$P_{S_{12}}(\lambda\sigma)=\det\Big{(}(S_{12})^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}\Big{)}=0\,.$ (20)
The zeros of characteristic polynomial are eigenvalues $\lambda_{i}$.
Next we can construct projection operators $(P_{i})^{A}{}_{B}$ on subspaces $\Xi_{i}$ (we will call it eigenspace) with dimension $d_{i}=dim\Xi_{i}$ corresponding to helicities $\lambda_{i}$
$(P_{i})^{A}{}_{B}=\frac{\Big{[}\prod_{j\neq i}^{n}\Big{(}S_{12}-\lambda_{j}\sigma\delta\Big{)}\Big{]}^{A}{}_{B}}{\prod_{j\neq i}^{n}\Big{(}\lambda_{i}\sigma-\lambda_{j}\sigma\Big{)}}\,,$ (21)
and obtain corresponding eigenfunctions
$\Psi_{i}^{A}(k)=(P_{i})^{A}{}_{B}\Psi^{B}(k)\,.$ (22)
From (21) and (22) it follows that space inversion connect projectors and eigenfunctions with opposite helicities
${\cal P}\,(P_{\lambda_{i}})^{A}{}_{B}=(P_{-\lambda_{i}})^{A}{}_{B}\,,\qquad{\cal P}\,\Psi_{\lambda_{i}}^{A}(k)=\Psi_{-\lambda_{i}}^{A}(k)\,.$ (23)
Generally, (see section 8 for details) with the help of projection operators, for tensor fields of rang $n$ we can rewrite equation (20) in the form
$\det(S_{12}-\lambda\sigma\delta)=\prod_{i=-n}^{n}(i-\lambda\sigma)^{\tilde{d}_{i}}\det P_{i}=0\,,\qquad$ (24)
and using $\sigma^{2}=1$ and $\tilde{d}_{i}=\tilde{d}_{-i}$
$\det(S_{12}-\lambda\sigma\delta)=(-\lambda\sigma)^{\tilde{d}_{0}}\prod_{i=1}^{n}(i^{2}-\lambda^{2})^{\tilde{d}_{i}}\,\,\prod_{i=-n}^{n}\det P_{i}=0\,.\qquad$ (25)
Exponents in multipliers $\tilde{d}_{i}$ are algebraic multiplicity while $d_{i}$ as dimensions of eigenspaces $\Xi_{i}$ are geometric multiplicity. Generally, $d_{i}\leq\tilde{d}_{i}$, but in our case $d_{i}=\tilde{d}_{i}$ so that from now on we will use a unique notation $d_{i}$.
Note that for tensor fields of rang $n$ relation $\sum_{i=-n}^{n}d_{i}=d_{0}+2\sum_{i=1}^{n}d_{i}=4^{n}$ is satisfied and spectrum of helicities is
$\lambda_{0}=0\,,\qquad\lambda_{\pm i}=\pm i\,,\qquad i=(1,2,\cdots,n)\,.$ (26)
Let us stress that gauge invariant fields $\Psi_{i}^{A}(k)$ (see the next subsection for explanation) are representations of Poincare group, but in general case they are not irreducible representations. To get irreducible representations we must separate fields $\Psi_{i}^{A}(k)$ using symmetry properties. Note that dimension of subspace $d_{i}=dim\Xi_{i}$ corresponding to helicity $\lambda_{i}$ is in fact number of degrees of freedom of fields $\Psi_{i}^{A}(k)$. Similarly, dimension of subspace of irreducible representation is number of degrees of freedom of corresponding fields. We are going to check it in all examples.
2.7 Lorentz transformations induce gauge transformations
We solved first equation (2.3) and it produces complete set of eigenvalues and eigenfunctions. But we have additional two conditions. It is of great importance to investigate the compatibility of all equations (2.3).
2.7.1 Gauge dependent fields
In equations for standard momentum (2.3) operators $\Pi_{1}$ and $\Pi_{2}$ should annihilate field $\Psi^{A}(k)$. But explicit calculation shows that this does not happen in some physically relevant cases. Since equations (2.3) are condition for Lorentz invariance, violation of these conditions breaks Lorentz invariance. To measure this violation we introduce expression
$\delta\Psi_{i}^{A}(\varepsilon_{1},\varepsilon_{2})(k)=i\Big{(}\varepsilon_{1}\Pi_{1}+\varepsilon_{2}\Pi_{2}\Big{)}^{A}{}_{B}\Psi_{i}^{B}(k)\,,\qquad$ (27)
where $\varepsilon_{1}$ and $\varepsilon_{2}$ are some parameters.
Field equations must be Lorentz invariant and consequently they should not depend on $\delta\Psi_{i}^{A}(\varepsilon_{1},\varepsilon_{2})(k)$. It means that
$\delta\Psi_{i}^{A}(\varepsilon_{1},\varepsilon_{2})(k)={\cal G}_{i}\Big{(}k,\Omega_{\pm}\Big{)}\,,\qquad\Omega_{\pm}=\Omega_{\pm}(\varepsilon_{\pm})$ (28)
are gauge transformations of fields $\Psi_{i}^{A}(k)$ with gauge parameters $\Omega_{\pm}(\varepsilon_{\pm})$. So, as Weinberg noted Lorentz transformations induce gauge transformations [4].
We can perform transition from standard momentum $k^{a}$ to arbitrary one $p^{a}$ and using relation $p^{a}=i\partial^{a}$ we can obtain gauge transformation in coordinate space
$\delta\Psi_{i}^{A}(\Omega_{\pm},x)={\cal G}_{i}\Big{(}i\partial,\Omega_{\pm}(x)\Big{)}\,.$ (29)
Finally, we can chose the action
$S(\Psi_{i})=\int d^{4}x{\cal L}\Big{(}\Psi_{i}^{A}(x)\Big{)}\,,$ (30)
as a gauge invariant expressions of $\Psi_{i}^{A}(x)$, which means that following condition must be satisfied
$S\Big{(}\Psi_{i}+\delta\Psi_{i}\Big{)}=S\Big{(}\Psi_{i}\Big{)}\,.$ (31)
Fields which violate Lorentz invariance are gauge dependent fields and we will simply call them gauge fields. They are not representation of Poincare group. If we want to weaken the statement they are representation of Poincare group up to gauge transformations, since corresponding field strength and Lagrangian are representations of Poincare group. Therefore, for these fields we will introduce name incomplete representation of Poincare group.
2.7.2 Gauge invariant fields
Gauge invariant fields are representation of Poincare group. So, they are physical and we can use them without restrictions. We can obtain their form as solutions of equations (2.3). That leads us to field equations of gauge invariant fields as some constraints on the obtained solution.
Next we will consider the most important physical cases of massless fields: scalar field, vector field with helicities $\lambda=\pm 1$ (photon) and symmetric second rank tensor with helicities $\lambda=\pm 2$ (metric). We will also consider second rank tensor with helicities $\lambda=\pm 1$ (field strength) and fourth rank tensor with helicities $\lambda=\pm 2$ (Riemann curvature).
3 Scalar field
The simplest example is scalar field $\varphi(x)$. In that case spin operator $S_{ab}$ vanishes. In particular $S_{12}=0$ which means that helicity is zero, $\lambda=0$. From $\Pi_{1}=0=\Pi_{2}$ we obtain $\delta\varphi(\varepsilon_{1},\varepsilon_{2})(x)=0$. Therefore, massless scalar field $\varphi(x)$ is gauge invariant, and we can take it as physical field without restrictions. It is one dimensional irreducible representation of Poinace group and consequently it has one degree of freedom. Equation of motion is massless Klein-Gordon equation $\partial^{2}\varphi(x)=0$.
4 Vector field and Maxwell equations
The case of massless vector field is most important. Besides representing electromagnetic interaction we will see that solutions for massless tensor fields of arbitrary rang comes down to the case of vector field.
4.1 Eigenvalues and projection operators
For vector field we have $A,B\to a,b$, $\Psi^{A}\to V^{a}$ and
$(S_{ab})^{A}{}_{B}\to\Big{(}S_{ab}\Big{)}^{c}{}_{d}=i\Big{(}\delta^{c}_{a}\,\eta_{bd}-\delta^{c}_{b}\,\eta_{ad}\Big{)}\,,\qquad\Rightarrow\qquad(S_{12})^{a}{}_{b}=i\Big{(}\delta^{a}_{1}\eta_{2b}-\delta^{a}_{2}\eta_{1b}\Big{)}\,.$ (1)
The consistency condition requires that characteristic polynomial vanishes. It means
$\det\Big{(}{\cal S}_{12}-\lambda\sigma\Big{)}^{a}{}_{b}=\lambda^{2}(\lambda-1)(\lambda+1)=0\,,$ (2)
with solutions for eigenvalues
$\lambda_{0}=0\,,\qquad\lambda_{\pm 1}=\pm 1\,.\qquad$ (3)
Note that eigenspace $\Xi_{0}$ corresponding to $\lambda_{0}=0$ is two dimensional $d_{0}=dim\Xi_{0}=2$ and consequently corresponding eigenvecor carries two degrees of freedom. The eigenspaces $\Xi_{\pm 1}$ corresponding to $\lambda_{\pm 1}=1$ are one dimensional $d_{\pm 1}=dim\Xi_{\pm 1}=1$ and consequently each of corresponding eigenvecors carries one degree of freedom.
We already had general expression for projection operators (21) which for vector fields, where $n=3$, takes the form
$(P_{i})^{a}{}_{b}=\frac{\prod_{j\neq i}^{3}\Big{(}(S_{12})^{a}{}_{b}-\lambda_{j}\sigma\delta^{a}_{b}\Big{)}}{\prod_{j\neq i}^{3}\Big{(}\lambda_{i}\sigma-\lambda_{j}\sigma\Big{)}}\,,$ (4)
or explicitly
$(P_{0})^{a}{}_{b}(k)=\delta^{a}_{b}-(S_{12}^{2})^{a}{}_{b}=\delta^{a}_{b}-\delta^{a}_{\alpha}\delta_{b}^{\alpha}\,,\qquad(\alpha=1,2)$
$(P_{\pm 1})^{a}{}_{b}(k)=\frac{1}{2}\Big{(}(S_{12}^{2})^{a}{}_{b}\pm\sigma(S_{12})^{a}{}_{b}\Big{)}=\frac{1}{2}\Big{[}\delta^{a}_{\alpha}\delta_{b}^{\alpha}\pm i\sigma\Big{(}\delta^{a}_{1}\eta_{2b}-\delta^{a}_{2}\eta_{1b}\Big{)}\Big{]}\,.$ (5)
4.2 Basic vectors
For further benefits we introduce basic vectors $k^{a}$ and $q^{a}$ in two dimensional eigenspace $\Xi_{0}$. Then complete set of basic vectors $e^{a}_{i}=\{k^{a},q^{a},\breve{p}_{+}^{a},\breve{p}_{-}^{a}\}$ where $i=\{0+,0-,+1,-1\}$ are
$k^{a}=\delta^{a}_{0}-\sigma\delta^{a}_{3}\,,\quad q^{a}=\delta^{a}_{0}+\sigma\delta^{a}_{3}\,,\quad\breve{p}_{\pm}^{a}=\delta^{a}_{1}\pm i\sigma\delta^{a}_{2}\,.\qquad$ (6)
It is easy to check that
$(S_{12})^{a}{}_{b}e^{b}_{i}=\lambda_{i}e^{a}_{i}\,,$ (7)
where $\lambda_{i}$ take the values from (3). Consequently, basic vector $e^{a}_{i}$ carries helicity $\lambda_{i}$. In particular, basic vectors $\breve{p}_{\pm}^{a}$ carry helicities ±$\pm 1$ and both basic vectors $k^{a}$ and $q^{a}$ carry helicities 00. If $n_{\pm}$ are numbers of vectors $\breve{p}_{\pm}^{a}$ in the representation of field $\Psi^{A}$, then its helicity is
$\lambda=n_{+}+n_{-}\,.$ (8)
Note that tensors with highest helicity (for example tensor of rang $n$ with helicities $\lambda_{\pm n}=\pm n$) have a form $\breve{p}_{\pm}^{a_{1}}\breve{p}_{\pm}^{a_{2}}\cdots\breve{p}_{\pm}^{a_{n}}\mathbf{T}$. Consequently, tensors with highest helicity are one dimensional and they are symmetric in all indices. We will confirm that in all examples.
Projection operators in terms of basic vectors take the form
$(P_{0+})^{a}{}_{b}(k)=\frac{1}{2}q^{a}k_{b}\,,\qquad(P_{0-})^{a}{}_{b}(k)=\frac{1}{2}k^{a}q_{b}\,,\qquad(P_{\pm 1})^{a}{}_{b}(k)=-\frac{1}{2}\breve{p}_{\pm}^{a}(\breve{p}_{\mp})_{b}\,.$ (9)
We can check that above expressions are really projectors using product of basic vectors where only nontrivial products are
$k^{a}q_{a}=2\,,\qquad\breve{p}_{\pm}^{a}(\breve{p}_{\mp})_{a}=-2\,.$ (10)
Note that all basic vectors are light-like.
4.3 Eigenfunctions of operator $(S_{12})^{a}{}_{b}$
The eigenfunctions of operator $(S_{12})^{a}{}_{b}$ are $V_{i}^{a}(k)=(P_{i})^{a}{}_{b}V^{b}$ where $i=\{0+,0-,+1,-1\}$, or explicitly
$V_{0+}^{a}(k)=\frac{1}{2}q^{a}k_{b}V^{b}\equiv q^{a}\mathbf{V}_{k}\,,\qquad V_{0-}^{a}(k)=\frac{1}{2}k^{a}q_{b}V^{b}\equiv k^{a}\mathbf{V}_{q}\,,\qquad$
$V_{\pm 1}^{a}(k)=-\frac{1}{2}\breve{p}_{\pm}^{a}(\breve{p}_{\mp})_{b}V^{b}\equiv\breve{p}_{\pm}^{a}\mathbf{V}_{\mp}\,.\qquad$ (11)
It is easy to check that relation (8) is satisfied since scalars like $\mathbf{V}_{\mp}=(\breve{p}_{\mp})_{b}V^{b}$ carry helicity 00.
The basic vectors determines the characteristics of eigenvectors. They have the same helicity and as we will see the same gauge transformations and the same parity. Here all eigenfunctions $V_{i}^{a}(k)$ are potential representations of Poincare group, but only gauge invariant ones are really representations of Poincare group.
Under space inversion vectors $k^{a}$ and $q^{a}$ are invariant while
${\cal P}\breve{p}_{\pm}^{a}=-\breve{p}_{\mp}^{a}\,.$ (12)
It confirms that general statements (23) valid for vector fields
${\cal P}(P_{\pm 1})^{a}{}_{b}(k)=(P_{\mp 1})^{a}{}_{b}(k)\,,\qquad{\cal P}V_{\pm 1}^{a}(k)=V_{\mp 1}^{a}(k)\,.$ (13)
The vector field $V^{a}$ has four components. There are four corresponding eigenfunctions $V^{a}_{i}$ with four components. But they are not independent. In fact, if we multiply all components $V^{a}_{i}$ with all basic vectors we obtain only four nonzero expressions
$k_{a}V_{0+}^{a}=k_{b}V^{b}=2\mathbf{V}_{k}\,,\qquad q_{a}V_{0-}^{a}=q_{b}V^{b}=2\mathbf{V}_{q}\,,\qquad(\breve{p}_{\mp})_{a}V_{\pm 1}^{a}=(\breve{p}_{\mp})_{b}V^{b}=-2\mathbf{V}_{\mp}\,.\qquad$ (14)
4.4 Gauge transformations of massless vector fields
Gauge transformations of eigenfunctions of operator $(S_{12})^{a}{}_{b}$ for massless vector fields
$\delta V_{i}^{a}(\varepsilon_{1},\varepsilon_{2})(k)=i\Big{(}\varepsilon_{1}\Pi_{1}+\varepsilon_{2}\Pi_{2}\Big{)}^{a}{}_{b}V_{i}^{b}(k)\,,\qquad$ (15)
are particular case of Eq.(27). Explicitly, using expressions
$(\Pi_{1})^{a}{}_{b}=i\Big{(}\delta^{a}_{0}\eta_{1b}-\delta^{a}_{1}\eta_{0b}+\delta^{a}_{1}\eta_{3b}-\delta^{a}_{3}\eta_{1b}\Big{)}\,,\qquad(\Pi_{2})^{a}{}_{b}=i\Big{(}\delta^{a}_{0}\eta_{2b}-\delta^{a}_{2}\eta_{0b}+\delta^{a}_{2}\eta_{3b}-\delta^{a}_{3}\eta_{2b}\Big{)}\,,\qquad$ (16)
we get
$\delta V_{i}^{a}(\varepsilon_{1},\varepsilon_{2})=-\Big{[}\varepsilon_{1}\Big{(}\delta^{a}_{0}\eta_{1b}-\delta^{a}_{1}\eta_{0b}+\delta^{a}_{1}\eta_{3b}-\delta^{a}_{3}\eta_{1b}\Big{)}+\varepsilon_{2}\Big{(}\delta^{a}_{0}\eta_{2b}-\delta^{a}_{2}\eta_{0b}+\delta^{a}_{2}\eta_{3b}-\delta^{a}_{3}\eta_{2b}\Big{)}\Big{]}V_{i}^{b}$
$=\left|{\begin{array}[]{cccc}0&\varepsilon_{1}&\varepsilon_{2}&0\\ \varepsilon_{1}&0&0&\varepsilon_{1}\\ \varepsilon_{2}&0&0&\varepsilon_{2}\\ 0&-\varepsilon_{1}&-\varepsilon_{2}&0\\ \end{array}}\right|\left|{\begin{array}[]{c}V^{0}_{i}\\ V^{1}_{i}\\ V^{2}_{i}\\ V^{3}_{i}\\ \end{array}}\right|\,.\qquad\qquad\qquad$ (25)
In particular, the gauge transformations of basic vectors have simple form
$\delta k^{a}=0\,,\qquad\delta q^{a}=\varepsilon_{+}\breve{p}_{-}^{a}+\varepsilon_{-}\breve{p}_{+}^{a}\,,\qquad\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}\,.\qquad(\varepsilon_{\pm}=\varepsilon_{1}\pm i\sigma\varepsilon_{2})\qquad$ (26)
The four independent scalars (14) are gauge invariant. So, gauge transformations of eigenfunctions $V_{i}^{a}(k)=(P_{i})^{a}{}_{b}V^{b}(k)$ are determined by gauge transformations of basic vectors.
Let us stress that only component $V_{0-}^{a}(k)$ is gauge invariant, $\delta V_{0-}^{b}(k)=0$, and consequently only this part is irreducible representation of Poincae group. The other components $V_{0+}^{a}(k)$ and $V_{\pm 1}^{b}(k)$ are not gauge invariant and they transform as
$\delta V_{0+}^{a}(k)=\omega_{+}\breve{p}_{-}^{a}+\omega_{-}\breve{p}_{+}^{a}\,,\qquad\omega_{\pm}(k)=\frac{1}{2}k_{a}V^{a}\varepsilon_{\pm}\,,$
$\delta V_{\pm 1}^{a}(k)=k^{a}\Omega_{\pm}\,,\qquad\Omega_{\pm}(k)=-\frac{1}{2}(\breve{p}_{\mp})_{b}V^{b}\varepsilon_{\pm}\,.$ (27)
From (26) we can conclude that under space inversion ${\cal P}\varepsilon_{\pm}=-\varepsilon_{\mp}$. It produces ${\cal P}\omega_{\pm}=-\omega_{\mp}$ and ${\cal P}\Omega_{\pm}=\Omega_{\mp}$. It is in agreement with relations ${\cal P}V_{0+}^{a}(k)=V_{0+}^{a}(k)$ and ${\cal P}V_{\pm}^{a}(k)=V_{\mp}^{a}(k)$.
4.5 Massless vector field with helicity $\lambda=0$
Component $V_{0+}^{a}(k)$ is not gauge invariant but from its gauge transformation (4.4) we can conclude that scalar $\varphi_{1}=k_{a}V_{0+}^{a}(k)$ is gauge invariant. Component $V_{0-}^{a}(k)$ is already gauge invariant and unique scalar which we can construct from this vector is $\varphi_{2}=q_{a}V_{0-}^{a}(k)$. These two scalars are first two invariants in (14). Since they are gauge invariant there are no restrictions on scalar fields $\varphi_{1}$ and $\varphi_{2}$ and we can use them as physical fields. These two fields fill two dimensional eigenspace with $\Xi_{0}$ and they have together two degrees of freedom. They satisfy massless Klein-Gordon equation.
4.6 Action for massless vector field with helicities $\lambda=\pm 1$ produces Maxwell equations
The electromagnetic interaction is symmetric under space inversion. Therefore, instead of two components $V_{+1}^{a}(k)$ and $V_{-1}^{a}(k)$ which are incomplete irreducible representation of proper Poincare group we will introduce one vector field $A^{a}(k)$ which is incomplete irreducible representation of wider group which includes space inversion. Using (11) we obtain
$A^{a}(k)=\alpha V^{a}_{+1}(k)+\beta V^{a}_{-1}(k)\,.\qquad$ (28)
We also introduce one parameter $\Omega(k)=\alpha\Omega_{+}(k)+\beta\Omega_{-}(k)$ instead of two components $\Omega_{+}(k)$ and $\Omega_{-}(k)$ related by space inversion.
As we have already seen representations $V_{+1}^{a}(k)$ and $V_{-1}^{a}(k)$ are one dimensional which means that each of them has one degree of freedom. So, field $A^{a}(k)$ has two degrees of freedom and as we will see it describes the photon.
Now, we are ready to introduce action as gauge invariant expressions of $A^{a}(k)$. First we perform transition from standard momentum $k^{a}$ to arbitrary one $p^{a}$ and then to coordinate dependent fields using $p^{a}=i\partial^{a}$. For real field $A^{a}(x)$, according to (14) and (28) we have
$A^{a}(x)=\int d^{4}p\Big{(}A^{a}(p)+A^{a}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}V^{a}_{+1}(p)+V^{a}_{+1}(-p)\Big{)}+\beta\Big{(}V^{a}_{-1}(p)+V^{a}_{-1}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (29)
With the help of second line in (4.4) we obtain
$\delta A^{a}(x)=\int d^{4}p\,p^{a}\Big{[}\alpha\Big{(}\Omega_{+1}(p)-\Omega_{+1}(-p)\Big{)}+\beta\Big{(}\Omega_{-1}(p)-\Omega_{-1}(-p)\Big{)}\Big{]}e^{-ipx}\,,$ (30)
and using (17) and (19)
$\delta A^{a}(x)=\partial^{a}\Big{(}\alpha\Omega_{+1}(x)+\beta\Omega_{-1}(x)\Big{)}=\partial^{a}\Omega(x)\,.$ (31)
Action should depend only on the gauge invariant combination of vector fields, in fact of field strength $F_{ab}=\partial_{a}A_{b}-\partial_{b}A_{a}$. Since it must be scalar we can take
$I_{0}=-\frac{1}{4}\int d^{4}xF_{ab}F^{ab}\,.$ (32)
It is not obvious that field equation, $\partial^{2}A^{a}-\partial^{a}\partial_{b}A^{b}=0$, which follows from (32) has two degrees of freedom. The gauge transformation (31) has one parameter. We can use Lorentz gauge fixing $\partial_{a}A^{a}=0$ which kills one component of vector field and after that equation of motion takes the form $\partial^{2}A^{a}=0$. But now both gauge fixing and field equation are invariant under additional special gauge transformation $\delta A^{a}=\partial^{a}\Omega_{0}$, such that $\partial^{2}\Omega_{0}=0$. Corresponding gauge fixing kills additional degree of freedom and consequently electromagnetic field $A^{a}(x)$ has two degrees of freedom.
Interaction with other fields can be described by the action
$I_{int}\Big{(}A^{a}\Big{)}=\int d^{4}xA^{a}J_{a}\,,$ (33)
with requirement that it is gauge invariant $I_{int}\Big{(}A^{a}+\delta A^{a}\Big{)}=I_{int}\Big{(}A^{a}\Big{)}$. After partial integration we obtain $\int d^{4}x\Omega\,\partial^{a}J_{a}=0$ and for arbitrary $\Omega$ it produces condition on current to be conserved $\partial^{a}J_{a}=0$.
The complete action is
$I=I_{0}+I_{int}=\int d^{4}x\Big{(}-\frac{1}{4}F_{ab}F^{ab}+A^{a}J_{a}\Big{)}\,.$ (34)
Variation with respect to $A^{a}$ produces Maxwell equations.
5 Second rang tensor
In this section we will consider general features of second rang tensor.
5.1 Projection operators
In the case of second rang tensor we have $A\to(ab)$ and $\Psi^{A}\to T^{ab}$. We will keep notations $(P_{i})^{ab}{}_{cd}$ for projectors and $(S_{12})^{ab}{}_{cd}$ for spin operator of second rang tensor. Since we will use abbreviated notation we introduce new notations for vector operators, $(P_{0})^{a}{}_{b}=(\pi_{0})^{a}{}_{b}$ and $(P_{\pm 1})^{a}{}_{b}=(\pi_{\pm})^{a}{}_{b}$ for projectors and $(S_{12})^{a}{}_{b}=\rho^{a}{}_{b}$ for spin operator.
According to (2) in abbreviated notation representation of spin operator for second rang tensor takes a form
$(S_{12})^{ab}{}_{cd}=\rho^{a}{}_{c}\delta^{b}_{d}+\delta^{a}_{c}\rho^{b}{}_{d}=\Big{(}\rho\delta+\delta\rho\Big{)}^{ab}{}_{cd}\,,\qquad$ (1)
where $\rho^{a}{}_{b}$ has been defined in (1). Note that multiplication in abbreviated notation means that we multiply first term with first term and second term with second term $(A\ast B)(C\ast D)=(AC\ast BD)$.
Using relations $\rho=\pi_{+}-\pi_{-}$ and $\delta=\pi_{0}+\pi_{+}+\pi_{-}$ we obtain (for simplicity we will omit indices)
$S_{12}=P_{1}+2P_{2}-P_{-1}-2P_{-2}\,,\qquad\delta\delta=P_{1}+P_{2}+P_{-1}+P_{-2}+P_{0}\,,\qquad$ (2)
where
$P_{0}=\pi_{0}\pi_{0}+\pi_{+}\pi_{-}+\pi_{-}\pi_{+}\,,\qquad P_{\pm 1}=\pi_{0}\pi_{\pm}+\pi_{\pm}\pi_{0}\,,\qquad P_{\pm 2}=\pi_{\pm}\pi_{\pm}\,.\qquad$ (3)
Using the fact that $\pi_{0},\pi_{+}$ and $\pi_{-}$ are projectors we can conclude that $(P_{i})^{ab}{}_{cd}$ $(i=0,\pm 1,\pm 2)$ are projectors also.
5.2 Solutions of $S_{12}$ equation
For second rang tensor characteristic matrix takes the form
$S_{12}-\lambda\sigma\delta\delta=-\lambda\sigma P_{0}+(1-\lambda\sigma)P_{1}+(2-\lambda\sigma)P_{2}-(1+\lambda\sigma)P_{-1}-(2+\lambda\sigma)P_{-2}\,.\qquad$ (4)
Since $(P_{i})^{ab}{}_{cd}$ are projectors, characteristic polynomial is already factored and we have
$\det\Big{(}S_{12}-\lambda\sigma\delta\delta\Big{)}=\Big{(}-\lambda\sigma\Big{)}^{d_{0}}\Big{(}1-\lambda\sigma\Big{)}^{d_{1}}\Big{(}1+\lambda\sigma\Big{)}^{d_{-1}}\Big{(}2-\lambda\sigma\Big{)}^{d_{2}}\Big{(}2+\lambda\sigma\Big{)}^{d_{-2}}\prod_{i=-2}^{2}\det P_{i}=0\,,\qquad\qquad$ (5)
where $d_{i}$ are dimensions of eigenspaces $\Xi_{i}$. Dimension of subspace projected by $\pi_{0}$ is $2$ while dimensions of subspaces projected by $\pi_{+}$ and $\pi_{-}$ are $1$, so that in our case $d_{0}=6\,,\,d_{\pm 1}=4\,,\,d_{\pm 2}=1$. It is in agreement with general relation since $6+2(4+1)=4^{2}$.
Then with the help of relation $\sigma^{2}=1$ we obtain
$\det\Big{(}S_{12}-\lambda\sigma\delta\delta\Big{)}=\Big{(}-\lambda\Big{)}^{6}\Big{(}1-\lambda^{2}\Big{)}^{4}\Big{(}4-\lambda^{2}\Big{)}\prod_{i=-2}^{2}\det P_{i}=0\,.\qquad$ (6)
Using the fact that $\det P_{i}\neq 0$ we can find that spectrum of helicities is
$\lambda_{0}=0\,,\qquad\lambda_{\pm 1}=\pm 1\,,\qquad\lambda_{\pm 2}=\pm 2\,.\qquad$ (7)
Consequently, the eigenfunctions of operator $S_{12}$ in full notation have the form
$T_{i}^{ab}(k)=(P_{i})^{ab}{}_{cd}T^{cd}(k)\,,\qquad(i=0,\pm 1,\pm 2)$ (8)
where projection operators $(P_{i})^{ab}{}_{cd}$ are defined in (3).
6 Second rang tensor with helicities $\lambda=\pm 2$ and general relativity in weak field approximation
Second rang tensor with helicities $\lambda=\pm 2$ is of particular importance since its symmetric part (metric) contribute to general relativity.
6.1 Gauge transformation for highest helicities $\lambda=\pm 2$ of second rang tensors
Each component of the second rang tensor with highest helicities $\lambda=\pm 2\,$
$T_{\pm 2}^{ab}(k)=(P_{\pm 2})^{ab}{}_{cd}T^{cd}(k)=(\pi_{\pm 1})^{a}{}_{c}(\pi_{\pm 1})^{b}{}_{d}T^{cd}(k)=\frac{1}{4}(\breve{p}_{\mp})_{c}(\breve{p}_{\mp})_{d}T^{cd}(k)\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}\,,\qquad$ (1)
is one dimensional representation and consequently it has one degree of freedom. Note that by definition it is symmetric tensor and relation (8) is satisfy.
With the help of (2) we can express little group generators for massless second rang tensors $(\Pi_{1})^{ab}{}_{cd}$ and $(\Pi_{2})^{ab}{}_{cd}$ in terms of corresponding generators for vectors $(\Pi_{1})^{a}{}_{b}$ and $(\Pi_{2})^{a}{}_{b}$
$(\Pi_{1})^{ab}{}_{cd}=(\Pi_{1})^{a}{}_{c}\delta^{b}_{d}+\delta^{a}_{c}(\Pi_{1})^{b}{}_{d}\,,\qquad(\Pi_{2})^{ab}{}_{cd}=(\Pi_{2})^{a}{}_{c}\delta^{b}_{d}+\delta^{a}_{c}(\Pi_{2})^{b}{}_{d}\,.$ (2)
In that case instead of general expression for gauge transformation (27) we have
$\delta T_{i}^{ab}(\varepsilon_{1},\varepsilon_{2})(k)=i\Big{(}\varepsilon_{1}\Pi_{1}+\varepsilon_{2}\Pi_{2}\Big{)}^{ab}{}_{cd}\,T_{i}^{cd}(k)\,.\qquad$ (3)
Using expressions for highes helicities (1) and gauge transformation for basic vector $\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}$ we obtain
$\delta T_{\pm 2}^{ab}(k)=k^{a}\Omega_{\pm}^{b}+k^{b}\Omega_{\pm}^{a}\,,\qquad\Omega_{\pm}^{a}(k)=\frac{1}{4}(\breve{p}_{\mp})_{c}(\breve{p}_{\mp})_{d}T^{cd}\breve{p}_{\pm}^{a}\varepsilon_{\pm}\,.\qquad$ (4)
Note that again ${\cal P}T_{\pm 2}^{ab}(k)=T_{\mp 2}^{ab}(k)$ and ${\cal P}\Omega_{\pm}^{a}=\Omega_{\mp}^{a}$.
Here we came to a similar conclusion for highest helicities of second rang tensor as in the case of vector field: it is not physical since it is not Lorentz invariant.
6.2 General relativity in weak field approximation
As in the case of vector fields, components of massless second rang tensors of opposite helicities $\lambda=2$ and $\lambda=-2$ are related by transformation of space inversion. Since gravitation interaction is symmetric under space inversion it is common to treat both components of incomplete irreducible representation of proper Poincare group as a single particle called graviton. So, instead of two polarizations $T_{+2}^{ab}(k)$ and $T_{-2}^{ab}(k)$, both with one degree of freedom, we will introduce one symmetric tensor $h^{ba}(k)=h^{ab}(k)$ with two degrees of freedom, which is incomplete irreducible representation of Poincare group with space inversion. Using (11) we obtain
$h^{ab}(k)=\alpha T^{ab}_{+2}(k)+\beta T^{ab}_{-2}(k)\,.\qquad$ (5)
We also introduce one parameter $\Omega^{a}(k)=\alpha\Omega_{+}^{a}(k)+\beta\Omega_{-}^{a}(k)$ instead of two components $\Omega_{+}^{a}(k)$ and $\Omega_{-}^{a}(k)$ related by space inversion.
If we go from standard momentum frame to arbitrary frame and then to coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, we can obtain gauge transformation in coordinate representation. For real field, according to (14) and (5) we have
$h^{ab}(x)=\int d^{4}p\Big{(}h^{ab}(p)+h^{ab}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}T^{ab}_{+2}(p)+T^{ab}_{+2}(-p)\Big{)}+\beta\Big{(}T^{ab}_{-2}(p)+T^{ab}_{-2}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (6)
With the help of (4) we obtain gauge transformation
$\delta h^{ab}(x)=\int d^{4}p\,\Big{\{}p^{a}\Big{[}\alpha\Big{(}\Omega_{+}^{b}(p)-\Omega_{+}^{b}(-p)\Big{)}+\beta\Big{(}\Omega_{-}^{b}(p)-\Omega_{-}^{b}(-p)\Big{)}\Big{]}$
$+p^{b}\Big{[}\alpha\Big{(}\Omega_{+}^{a}(p)-\Omega_{+}^{a}(-p)\Big{)}+\beta\Big{(}\Omega_{-}^{a}(p)-\Omega_{-}^{a}(-p)\Big{)}\Big{]}\Big{\}}e^{-ipx}\,,$ (7)
and using (17) and (19)
$\delta h^{ab}(x)=\partial^{a}\Big{(}\alpha\Omega_{+}^{b}(x)+\beta\Omega_{-}^{b}(x)\Big{)}+\partial^{b}\Big{(}\alpha\Omega_{+}^{a}(x)+\beta\Omega_{-}^{a}(x)\Big{)}=\partial^{a}\Omega^{b}(x)+\partial^{b}\Omega^{a}(x)\,.\qquad$ (8)
In order to get field equation we need gauge invariant function of tensor $h^{ab}(x)$. There are several solutions: Riemann, Ricci and scalar curvatures as well as Einstein tensor in weak field approximation. They can be obtained as small perturbation around flat metric using relation $g_{ab}=\eta_{ab}+h_{ab}$, see Refs.[7, 8].
Since equation of motion with respect to $h^{ab}$ has two indices we can chose Ricci or Einstein tensor. We are going later to include interaction and so we will take Einstein tensor for free field equation (for expression for Einstein tensor see Sec.8)
$G^{ab}(x)=R^{ab}-\frac{1}{2}\eta^{ab}R=\partial_{c}\partial^{b}h^{ac}-\partial^{2}h^{ab}-\partial^{a}\partial^{b}h+\partial^{a}\partial_{c}h^{cb}-\eta^{ab}\Big{(}\partial_{c}\partial_{d}h^{cd}-\partial^{2}h\Big{)}=0\,.\qquad$ (9)
Similarly as in the case of vector field it can be shown (see for example Ref.[8]) that field $h^{ab}$ which satisfy this equation has two degrees of freedom. The symmetric tensor $h^{ab}$ has $10$ components. To fix gauge invariance (8) we can use $4$ gauge fixings conditions $\partial_{a}h^{ab}=\frac{1}{2}\partial^{b}h^{a}{}_{a}$ which kills $4$ component of tensor field and after that equation of motion takes the form $\partial^{2}\Big{(}h^{ab}-\frac{1}{2}\eta^{ab}h\Big{)}=0$. But now both gauge fixing and field equation are invariant under additional special gauge transformation $\delta h^{ab}(x)=\partial^{a}\Omega_{0}^{b}(x)+\partial^{b}\Omega_{0}^{a}(x)$, such that $\partial^{2}\Omega_{0}^{a}=0$. Corresponding gauge fixing kills additional $4$ degree of freedom and consequently gravitation field $h^{ab}(x)$ has two degrees of freedom.
Field equation (9) can be obtained from the action
$I_{0}(h^{ab})=\int d^{4}x\Big{(}\frac{1}{2}\partial_{c}h_{ab}\partial^{c}h^{ab}-\partial_{c}h_{ab}\partial^{b}h^{ac}+\partial_{c}h^{ac}\partial_{a}h-\frac{1}{2}\partial_{a}h\partial^{a}h\Big{)}\,.$ (10)
Interaction with matter field $\Phi^{A}$ we can describe with the action quadratic in $h^{ab}$
$I_{int}(h^{ab})=\int d^{4}x{\cal L}(h^{ab},\Phi^{A})\,.$ (11)
Note that the integration measure is the flat Minkowski measure $d^{4}x$, as Lagrangian is already quadratic in $h^{ab}$.
From requirement that $I_{int}(h^{ab})$ is gauge invariant $I_{int}(h^{ab}+\delta h^{ab})=I_{int}(h^{ab})$ we have
$\int d^{4}x\delta h^{ab}\frac{\partial{\cal L}}{\partial h^{ab}}=\int d^{4}x\Big{(}\partial^{a}\Omega^{b}+\partial^{b}\Omega^{a}\Big{)}\Theta_{ab}=0\,.\qquad\Theta_{ab}=\frac{\partial{\cal L}}{\partial h^{ab}}$ (12)
Using partial integration and fact that $\Theta_{ab}$ is symmetric tensor we obtain that $\Theta_{ab}$ is the conserved energy-momentum tensor, $\partial_{a}\Theta^{ab}=0$.
The complete action is $I(h^{ab})=I_{0}(h^{ab})+I_{int}(h^{ab})$. Its variation with respect to $h^{ab}$ produces complete Einstein equations in weak field approximation
$G_{ab}(h^{ab})=\Theta_{ab}\,.$ (13)
7 Second rang tensor with helicities $\lambda=\pm 1$ produces Maxwell equations
Beside highes helicities $\lambda=\pm 2$, second rang tensor has also components with helicities $\lambda=\pm 1$. We expect that these components describe Maxwell equations in terms of field strength.
7.1 Second rang tensor with helicities $\lambda=\pm 1$
The components of second rang tensor with helicities $\lambda=\pm 1$ have a form
$T_{\pm 1}^{ab}(k)=(P_{\pm 1})^{ab}{}_{cd}T^{cd}(k)=\Big{(}(\pi_{0})^{a}{}_{c}(\pi_{\pm})^{b}{}_{d}+(\pi_{\pm})^{a}{}_{c}(\pi_{0})^{b}{}_{d}\Big{)}T^{cd}(k)\,.$ (1)
Since $\pi_{0}$ project to two dimensional space and $\pi_{+}$ and $\pi_{-}$ to one dimensional space, both components $T_{\pm 1}^{ab}(k)$ are four dimensional representations and they have four degrees of freedom.
To avoid influence of nontrivial gauge transformations of $(\pi_{0+})^{a}{}_{b}$ we will save only $(\pi_{0-})^{a}{}_{b}$ part (gauge invariant part of $(\pi_{0})^{a}{}_{b}$) which means that we will take $(\pi_{0})^{a}{}_{b}\to(\pi_{0-})^{a}{}_{b}$. With that choice we obtain
$T_{\pm 1(-)}^{ab}(k)=(P_{\pm 1(-)})^{ab}{}_{cd}T^{cd}(k)=\Big{(}(\pi_{0-})^{a}{}_{c}(\pi_{\pm})^{b}{}_{d}+(\pi_{\pm})^{a}{}_{c}(\pi_{0-})^{b}{}_{d}\Big{)}T^{cd}(k)\,.$ (2)
Note that now each component $T_{+1(-)}^{ab}(k)$ and $T_{-1(-)}^{ab}(k)$ has two degrees of freedom.
Using expressions for projectors we get
$T_{\pm 1(-)}^{ab}(k)=-\frac{1}{4}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{cd}(k)\Big{)}k^{a}\breve{p}_{\pm}^{b}-\frac{1}{4}\Big{(}(\breve{p}_{\mp})_{c}q_{d}T^{cd}(k)\Big{)}\breve{p}_{\pm}^{a}k^{b}\,.$ (3)
Since $k_{a}k^{a}=0$ and $k_{a}\breve{p}_{\pm}^{a}=0$ there are constraints $k_{a}T_{\pm 1(-)}^{ab}(k)=0$. The fields $T_{\pm 1(-)}^{ab}(k)$ with helicities $\lambda=\pm 1$ contain one basic vector $\breve{p}_{\pm}^{a}$ and consequently relation (8) is satisfy.
For symmetric tensor we obtain
$(T^{S}_{\pm 1(-)})^{(ab)}(k)=-\frac{1}{4}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{(cd)}(k)\Big{)}\Big{(}k^{a}\breve{p}_{\pm}^{b}+k^{b}\breve{p}_{\pm}^{a}\Big{)}\equiv\mathbf{T}_{\mp 2}^{S}\Big{(}k^{a}\breve{p}_{\pm}^{b}+k^{b}\breve{p}_{\pm}^{a}\Big{)}\,,\qquad$ (4)
or
$(T^{S}_{\pm 1(-)})^{(ab)}(k)=k^{a}\varphi_{\pm}^{b}(k)+k^{b}\varphi_{\pm}^{a}(k)\,,\qquad\varphi_{\pm}^{a}=\mathbf{T}_{\mp 2}^{S}\breve{p}_{\pm}^{a}\,,$ (5)
and for antisymmetric one
$(T^{A}_{\pm 1(-)})^{[ab]}(k)=-\frac{1}{4}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{[cd]}(k)\Big{)}\Big{(}k^{a}\breve{p}_{\pm}^{b}-k^{b}\breve{p}_{\pm}^{a}\Big{)}\equiv\mathbf{T}_{\mp 2}^{A}\Big{(}k^{a}\breve{p}_{\pm}^{b}-k^{b}\breve{p}_{\pm}^{a}\Big{)}\,,\qquad$ (6)
or
$(T^{A}_{\pm 1(-)})^{[ab]}(k)=k^{a}A_{\pm}^{b}(k)-k^{b}A_{\pm}^{a}(k)\,,\qquad A_{\pm}^{a}=\mathbf{T}_{\mp 2}^{A}\breve{p}_{\pm}^{a}\,.$ (7)
Here $\varphi_{\pm}^{a}$ and $A_{\pm}^{a}$ are new fields defined above. Note that both $\varphi_{\pm}^{a}$ and $A_{\pm}^{a}$ are proportional to $\breve{p}_{\pm}^{a}$, as well as fields $V_{\pm}^{a}$ in Eq.(4.3). According to scalar product of basic vectors $k_{a}\breve{p}_{\pm}^{a}=0$ we can conclude that $k_{a}\varphi_{\pm}^{a}=0$ and $k_{a}A_{\pm}^{a}=0$, which means that these solutions are in Lorentz gauge. We will see that these fields describe electromagnetic interaction.
Finally, each component $(T^{S})_{+1(-)}^{ab}(k)$, $(T^{S})_{-1(-)}^{ab}(k)$, $(T^{A})_{+1(-)}^{ab}(k)$ and $(T^{A})_{-1(-)}^{ab}(k)$ have one degree of freedom.
7.2 Gauge transformation for second rang tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Using gauge transformations for basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=k^{a}\varepsilon_{\pm}$ we can obtain gauge transformations for symmetric and antisymmetric parts of second rang tensor with helicities $\lambda=\pm 1$
$\delta(T^{S}_{\pm 1(-)})^{(ab)}(k)=k^{a}k^{b}\omega_{\pm}\,,\qquad\omega_{\pm}(k)=-\frac{1}{2}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{(cd)}(k)\Big{)}\varepsilon_{\pm}\,,\qquad$
$\delta(T^{A}_{\pm 1(-)})^{[ab]}(k)=0\,.\qquad\qquad\qquad\qquad$ (8)
So, symmetric tensor is gauge dependent and antisymmetric one is gauge invariant.
7.2.1 Symmetric tensor with helicities $\lambda=\pm 1$
Note that under space inversion ${\cal P}(T^{S}_{\pm 1(-)})^{(ab)}(k)=(T^{S}_{\mp 1(-)})^{(ab)}(k)$, ${\cal P}\varphi^{a}_{\pm}(k)=\varphi^{a}_{\mp}(k)$ and ${\cal P}\omega_{\pm}(k)=\omega_{\mp}(k)$. So, according to (11) instead of two components $(T^{S}_{+1(-)})^{(ab)}(k)$ and $(T^{S}_{-1(-)})^{(ab)}(k)$ we will introduce one symmetric tensor
$\varphi^{ab}(k)=\alpha(T^{S}_{+1(-)})^{(ab)}(k)+\beta(T^{S}_{-1(-)})^{(ab)}(k)\,,\qquad$ (9)
and instead of two components $\varphi^{a}_{\pm}$ we will introduce one vector field
$\varphi^{a}(k)=\alpha\varphi^{a}_{+}(k)+\beta\varphi^{a}_{-}(k)\,,\qquad$ (10)
which are incomplete irreducible representation of Poincare group with space inversion. The representation of the field $\varphi^{ab}(k)$ as well as of $\varphi^{a}(k)$ is two dimensional and consequently they have two degrees of freedom. Since the parameters $\omega_{\pm}(k)$ are also connected by parity transformation we can introduce one parameter $\omega(k)=\alpha\omega_{+}(k)+\beta\omega_{-}(k)$.
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, we obtain expression for symmetric tensor $\varphi^{ab}(x)$ in terms of vector field $\varphi^{a}(x)$ and theirs gauge transformations. For real field, according to (14) and (9) we have
$\varphi^{ab}(x)=\int d^{4}p\Big{(}\varphi^{ab}(p)+\varphi^{ab}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}(T^{S}_{+1(-)})^{(ab)}(p)+(T^{S}_{+1(-)})^{(ab)}(-p)\Big{)}+\beta\Big{(}(T^{S}_{-1(-)})^{(ab)}(p)+(T^{S}_{-1(-)})^{(ab)}(-p)\Big{)}\Big{]}e^{-ipx}\,.\qquad$ (11)
Then with the help of (5) we obtain
$\varphi^{ab}(x)=\int d^{4}p\,\Big{\{}p^{a}\Big{[}\alpha\Big{(}\varphi_{+}^{b}(p)-\varphi_{+}^{b}(-p)\Big{)}+\beta\Big{(}\varphi_{-}^{b}(p)-\varphi_{-}^{b}(-p)\Big{)}\Big{]}$
$+p^{b}\Big{[}\alpha\Big{(}\varphi_{+}^{a}(p)-\varphi_{+}^{a}(-p)\Big{)}+\beta\Big{(}\varphi_{-}^{a}(p)-\varphi_{-}^{a}(-p)\Big{)}\Big{]}\Big{\}}e^{-ipx}\,,$ (12)
and using (17) and (19)
$\varphi^{ab}(x)=\partial^{a}\Big{(}\alpha\varphi_{+}^{b}(x)+\beta\varphi_{-}^{b}(x)\Big{)}+\partial^{b}\Big{(}\alpha\varphi_{+}^{a}(x)+\beta\varphi_{-}^{a}(x)\Big{)}=\partial^{a}\varphi^{b}(x)+\partial^{b}\varphi^{a}(x)\,.\qquad$ (13)
To obtain gauge transformation we use (7.2) and (7.2.1)
$\delta\varphi^{ab}(x)(x)=\int d^{4}p\,p^{a}p^{b}\Big{[}\alpha\Big{(}\omega_{+}(p)-\omega_{+}(-p)\Big{)}+\beta\Big{(}\omega_{-}(p)-\omega_{-}(-p)\Big{)}\Big{]}e^{-ipx}\,,$ (14)
and with the help of (18) and (19)
$\delta\varphi^{ab}(x)(x)=\partial^{a}\partial^{b}\Big{(}\alpha\omega_{+}(x)+\beta\omega_{-}(x)\Big{)}=\partial^{a}\partial^{b}\omega(x)\,.\qquad$ (15)
So, gauge transformation of field $\varphi^{a}(x)$ takes the form
$\delta\varphi^{a}(x)=\partial^{a}\omega_{1}(x)\,,\qquad\omega_{1}(x)=\frac{1}{2}\omega(x)\,.$ (16)
The field $\varphi^{a}(x)$ behaves like electromagnetic field. It is vector field, it has the same helicity, the same gauge transformation and the same number of degrees of freedom. So, this field should not be considered as new field.
7.2.2 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Antisymmetric part is gauge invariant and consequently Lorentz invariant and we can use it as physical field. Instead of two components $(T^{A}_{+1(-)})^{[ab]}(k)$ and $(T^{A}_{-1(-)})^{[ab]}(k)$, according to (11) we will introduce one antisymmetric tensor
$F^{ab}(k)=\alpha(T^{A}_{+1(-)})^{[ab]}(k)+\beta(T^{A}_{-1(-)})^{[ab]}(k)\,,\qquad$ (17)
which is irreducible representation of Poincare group with space inversion. Similarly, from (7) we have
$F^{ab}(k)=k^{a}A^{b}(k)-k^{b}A^{a}(k)\,,\qquad A^{a}(k)=\alpha A_{+}^{a}(k)+\beta A_{-}^{a}(k)\,.$ (18)
Using (6) and scalar product of basic vectors we can conclude that there are two constraints on field $F^{ab}(k)$
$k_{a}F^{ab}(k)=0\,,\qquad\varepsilon_{abcd}k^{b}F^{cd}(k)=0\,.\qquad$ (19)
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, in the same way as before we get
$\partial_{a}F^{ab}(x)=0\,,\qquad\varepsilon_{abcd}\partial^{b}F^{cd}(x)=0\,,\qquad$ (20)
which are just vacuum Maxwell equations. As well as in the previous case the representation of the field $F^{ab}(x)$ is two dimensional and consequently it has two degrees of freedom. We could have expected that, since it describes Maxwell equations.
To express coupling of electromagnetic fields with matter (to a conserved current) we must solve second equation which produces
$F_{ab}(x)=\partial_{a}A_{b}(x)-\partial_{b}A_{a}(x)\,,\qquad$ (21)
which also follows from (18).
Note that $A_{a}(x)$ is gauge field since $F_{ab}(x)$ is invariant under transformation $\delta A_{a}(x)=\partial_{a}\Omega$. Then interaction with a matter can be express as in (33).
8 Fourth rang tensor with helicities $\lambda=\pm 2$ produces general relativity in weak field approximation
As symmetric second rang tensor with helicities $\lambda=\pm 1$ describe electromagnetic interaction through field strength, we can expect that fourth rang tensor with helicities $\lambda=\pm 2$ and appropriate symmetries describe gravitation interaction through Riemann tensor. In the case of fourth rang tensor we have $A\to(abcd)$ and $\Psi^{A}\to T^{abcd}$. In this section to simplify notation we will keep index $A$ instead of four indices $abcd$. When we include symmetry properties we will go back to the notation with indices $abcd$.
8.1 Fourth rang tensor
We are going to solve equation
$\Big{(}S^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}\Big{)}T^{B}=0\,,$ (1)
where $T^{A}$ is fourth rang tensor, $T^{A}=T^{abcd}$. The problem comes down to that of vector fields. According to (2) in abbreviated notation spin operator takes the form
$(S_{12})^{A}{}_{B}=\Big{(}\rho\delta^{3}+\delta\rho\delta^{2}+\delta^{2}\rho\delta+\delta^{3}\rho\Big{)}^{A}{}_{B}\,.\qquad$ (2)
Using expressions $\rho=\pi_{+}-\pi_{-}$ and $\delta=\pi_{0}+\pi_{+}+\pi_{-}$ after multiplications we obtain
$(S_{12})^{A}{}_{B}=\sum_{i=-4}^{4}i\Big{(}P_{i}\Big{)}^{A}{}_{B}\,,\qquad$ (3)
and
$\delta^{A}_{B}=(\delta^{4})^{A}{}_{B}=\Big{(}(\pi_{0}+\pi_{+}+\pi_{-})^{4}\Big{)}^{A}{}_{B}=\sum_{i=-4}^{4}\Big{(}P_{i}\Big{)}^{A}{}_{B}\,,$ (4)
where the multiplication factor $i$ in (3) is consequence of combinatorics.
Here $(P_{i})^{A}{}_{B}$ is sum of all terms $\pi_{i_{1}}\pi_{i_{2}}\pi_{i_{3}}\pi_{i_{4}}$ such that $\sum_{m=1}^{4}i_{m}=i$. We count $\pi_{+}$ as $\pi_{1}$ and $\pi_{-}$ as $\pi_{-1}$. Let us stress that order of multipliers is important, so that for example $\pi_{+}\pi_{0}\pi_{+}\pi_{0}$ is different from $\pi_{0}\pi_{+}\pi_{+}\pi_{0}$. Note that as a consequences of these rules, and taking into account that $\pi_{0},\pi_{+}$ and $\pi_{-}$ are projectors, all expressions $P_{i}$ where $i=0,\pm 1,\pm 2,\pm 3,\pm 4$ are projectors also, $P_{i}P_{j}=\delta_{ij}P_{i}$.
To obtain spectrum of helicities we are going to solve consistency condition that characteristic polynomial vanishes, $\det(S_{12}-\lambda\sigma\delta)^{A}{}_{B}=0$. Using (3) and (4) we obtain expression for characteristic polynomial
$(S_{12})^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}=\sum_{i=-4}^{4}(i-\lambda\sigma)(P_{i})^{A}{}_{B}\,.\qquad$ (5)
Since $(P_{i})^{A}{}_{B}$ are projection operators, expression $S_{12}-\lambda\sigma\delta$ is already factored and we have
$\det(S_{12}-\lambda\sigma\delta)=\prod_{i=-4}^{4}(i-\lambda\sigma)^{d_{i}}\,\det P_{i}=0\,,\qquad$ (6)
where $d_{i}$ are dimensions of eigenspaces $\Xi_{i}$. In our case
$d_{4}=1\,,\qquad d_{3}=8\,,\qquad d_{2}=28\,,\qquad d_{1}=56\,,\qquad d_{0}=70\,,$ (7)
with general relation $d_{i}=d_{-i}$. Note that $2(1+8+28+56)+70=4^{4}$.
Then with the help of relation $\sigma^{2}=1$ we obtain
$\det(S_{12}-\lambda\sigma\delta)=(-\lambda)^{d_{0}}\prod_{i=1}^{4}(i^{2}-\lambda^{2})^{d_{i}}\,\prod_{i=-4}^{4}\det P_{i}=0\,.\qquad$ (8)
Using the fact that $\det P_{i}\neq 0$ we can conclude that
$(-\lambda)^{d_{0}}\prod_{i=1}^{4}(i^{2}-\lambda^{2})^{d_{i}}=0\,,\qquad$ (9)
which produces spectrum of helicities for four rang tensor
$\lambda_{0}=0\,,\qquad\lambda_{\pm i}=\pm i\,.\qquad i=(1,2,\cdots,4)$ (10)
The eigenfunctions, which are our candidates for irreducible representations, for four rang massless tensors are
$T_{i}^{A}(k)=(P_{i})^{A}{}_{B}T^{B}(k)\,.\qquad(i=0,\pm 1,\pm 2,\cdots,\pm 4)$ (11)
Note that approach of this subsection can be easily generalized to tensors of arbitrary rang.
8.2 General relativity in terms of Riemann tensor
In this subsection we will find irreducible representations of fourth rang tensor with helicity $\lambda=\pm 2$. First we obtain irreducible representations of proper Poincare group, $T_{\pm 2(-)}^{abcd}(k)$ and then of wider group which combined Poincare group and parity, $R^{abcd}(k)$. The last one is linearized Riemann tensor. Finally, we obtain Einstein equations in vacuum.
8.2.1 Irreducible representations of proper Poincare group
The fourth rang tensor with helicities $\lambda=\pm 2$ has a form
$T_{\pm 2}^{abcd}(k)=(P_{\pm 2})^{abcd}{}_{efgh}T^{efgh}(k)$
$=\Big{[}\sum_{\Sigma_{1}}(\pi_{+})^{a}{}_{e}(\pi_{+})^{b}{}_{f}(\pi_{0})^{c}{}_{g}(\pi_{0})^{d}{}_{h}+\sum_{\Sigma_{2}}(\pi_{+})^{a}{}_{e}(\pi_{+})^{b}{}_{f}(\pi_{+})^{c}{}_{g}(\pi_{-})^{d}{}_{h}\Big{]}T^{efgh}(k)\,,\qquad$ (12)
where $\Sigma_{1}$ and $\Sigma_{2}$ means sum over all possible permutations of projectors.
As well as in the case of second rang tensor with helicities $\lambda=\pm 1$, to describe particles with helicities $\lambda=\pm 2$ with the help of fourth rang tensor, in every place we will substitute $\pi_{0}$ with its gauge invariant part $\pi_{0-}$. Then fourth rang tensor with helicities $\lambda=\pm 2$ has two terms: one with $2$ vectors $\breve{p}_{\pm}^{a}$ and $2$ vectors $k^{a}$ and the other with $3$ vectors $\breve{p}_{\pm}^{a}$ and one vector $\breve{p}_{\mp}^{a}$
$T_{\pm 2(-)}^{abcd}(k)=\sum_{\Sigma_{1}}\mathbf{T}_{\mp 4}^{\Sigma_{1}}\,\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}k^{c}k^{d}+\sum_{\Sigma_{2}}\mathbf{T}_{\mp 4}^{\Sigma_{2}}\,\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}\breve{p}_{\pm}^{c}\breve{p}_{\mp}^{d}\,,\qquad$ (13)
where $\Sigma_{1}$ and $\Sigma_{2}$ means sum over all possible permutations of basic vectors. The coefficients have a form
$\mathbf{T}_{\mp 4}^{\Sigma_{1}}=\frac{1}{2^{4}}(\breve{p}_{\mp})_{e}\,(\breve{p}_{\mp})_{f}\,q_{g}q_{h}\,T^{efgh}\,.\qquad\mathbf{T}_{\mp 4}^{\Sigma_{2}}=\frac{1}{2^{4}}(\breve{p}_{\mp})_{e}\,(\breve{p}_{\mp})_{f}\,(\breve{p}_{\mp})_{g}(\breve{p}_{\pm})_{h}\,T^{efgh}\,.\qquad$ (14)
If we count numbers of basic vectors $\breve{p}_{\pm}^{a}$ we can conclude that relation (8) $n_{\pm}+n_{\mp}=\lambda$ valid since ±$\pm 2+0=\pm 2$ and ±$\pm 3+(\mp 1)=\pm 2$.
We are interested in tensor antisymmetric in the first pair ($ab$) as well as in the second pair of indices ($cd$). Then only terms of the form
$\mathbf{T}_{\mp 4}^{\Sigma_{1}}\to\mathbf{T}_{\mp 4}^{R}=\frac{1}{2^{4}}\Big{(}q_{e}(\breve{p}_{\mp})_{f}-q_{f}(\breve{p}_{\mp})_{e}\Big{)}\,\Big{(}q_{g}(\breve{p}_{\mp})_{h}-q_{h}(\breve{p}_{\mp})_{g}\Big{)}\,\,T^{efgh}\,,\qquad$ (15)
survive and we obtain
$T_{\pm 2(-)}^{[ab][cd]}(k)=\mathbf{T}_{\mp 4}^{R}\,\Big{(}k^{a}\breve{p}_{\pm}^{b}-k^{b}\breve{p}_{\pm}^{a}\Big{)}\Big{(}k^{c}\breve{p}_{\pm}^{d}-k^{d}\breve{p}_{\pm}^{c}\Big{)}\,.\qquad$ (16)
This tensor is symmetric under exchange of the two pairs of indices ($ab$) and ($cd$). We can rewrite it in terms of second rang symmetric tensor $h_{\pm}^{ab}=\mathbf{T}_{\mp 4}^{R}\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}$ as
$T_{\pm 2(-)}^{[ab][cd]}(k)=k^{a}k^{c}h_{\pm}^{bd}-k^{a}k^{d}h_{\pm}^{bc}-k^{b}k^{c}h_{\pm}^{ad}+k^{b}k^{d}h_{\pm}^{ac}\,.\qquad$ (17)
It is easy to check symmetry properties of tensor $T_{\pm 2(-)}^{[ab][cd]}(k)$: cyclic permutation symmetry and Bianchi identity
$T_{\pm 2(-)}^{a[bcd]}=0\,,\qquad T_{\pm 2(-)}^{ab[cd}k^{e]}=0\,.$ (18)
With the help of expressions for scalar product of the basic vectors and (16) we obtain
$T_{\pm 2(-)}^{ac}(k)=\eta_{bd}T_{\pm 2(-)}^{abcd}(k)=0\,,\qquad T_{\pm 2(-)}(k)=\eta_{ab}T_{\pm 2(-)}^{ab}(k)=0\,.$ (19)
Using the gauge transformations of the basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}$ we have
$\delta T_{\pm 2(-)}^{[ab][cd]}(k)=0\,,$ (20)
which means that fields $T_{\pm 2(-)}^{[ab][cd]}(k)$ are gauge invariant and we can use them as a physical fields. They are irreducible representations of proper Poincare group.
8.2.2 Curvatures as irreducible representations of Poincare group with parity
As explained earlier, instead of two components $T_{+2(-)}^{[ab][cd]}(k)$ and $T_{-2(-)}^{[ab][cd]}(k)$ we will introduce tensor $R^{abcd}(k)$ which is irreducible representation of Poincare group with space inversion. With the help of (11) we have
$R^{abcd}(k)=\alpha T_{+2(-)}^{[ab][cd]}(k)+\beta T_{+2(-)}^{[ab][cd]}(k)=k^{a}k^{c}h^{bd}(k)-k^{a}k^{d}h^{bc}(k)-k^{b}k^{c}h^{ad}(k)+k^{b}k^{d}h^{ac}(k)\,,\qquad$
$h^{ab}(k)=\alpha h_{+}^{ab}(k)+\beta h_{-}^{ab}(k)=\alpha\mathbf{T}_{-4}^{R}(k)\breve{p}_{+}^{a}\breve{p}_{+}^{b}+\beta\mathbf{T}_{+4}^{R}(k)\breve{p}_{-}^{a}\breve{p}_{-}^{b}\,.\qquad\qquad\qquad$ (21)
The coordinate representation for real fields is
$R^{abcd}(x)=\int d^{4}p\Big{(}R^{abcd}(p)+R^{abcd}(-p)\Big{)}e^{-ipx}\,,\qquad$ (22)
and with the help of (18)
$R^{abcd}(x)=\partial^{a}\partial^{c}h^{bd}(x)-\partial^{a}\partial^{d}h^{bc}(x)-\partial^{b}\partial^{c}h^{ad}(x)+\partial^{b}\partial^{d}h^{ac}(x)\,.\qquad$ (23)
where according to (19)
$h^{ab}(x)=-\int d^{4}p\Big{[}h^{ab}(p)+h^{ab}(-p)\Big{]}e^{-ipx}\,.\qquad$ (24)
In the expression (23) we can recognize Riemann tensor see Ref.[7], and in the expression (24) metric tensor in weak field approximation.
Then we can calculate linearized Ricci tensor
$R^{bd}(x)=\eta_{ac}R^{abcd}(x)=\partial^{2}h^{bd}(x)-\partial_{a}\partial^{d}h^{ba}(x)-\partial^{b}\partial_{a}h^{ad}(x)+\partial^{b}\partial^{d}h(x)\,,\qquad$ (25)
linearized scalar curvature
$R(x)=\eta_{bd}R^{bd}=\partial^{2}h-\partial_{a}\partial_{b}h^{ba}-\partial_{a}\partial_{b}h^{ab}+\partial^{2}h=2\Big{(}\partial^{2}h-\partial_{a}\partial_{b}h^{ba}\Big{)}\,,\qquad$ (26)
and linearized Einstein tensor
$G^{bd}(x)=R^{bd}-\frac{1}{2}\eta^{bd}R=\partial^{2}h^{bd}-\partial_{a}\partial^{d}h^{ba}-\partial^{b}\partial_{a}h^{ad}+\partial^{b}\partial^{d}h-\eta^{bd}\Big{(}\partial^{2}h-\partial_{a}\partial_{b}h^{ba}\Big{)}\,.\qquad$ (27)
8.2.3 Einstein equations in vacuum
Using expressions (18), (20) and (8.2.2), and substituting $k^{a}\to p^{a}\to i\partial^{a}$ we can conclude that tensor $R^{abcd}(x)$ has standard symmetry properties of Riemann tensor: cyclic permutation symmetry and Bianchi identity
$R^{a[bcd]}(x)=0\,,\qquad R^{ab[cd}k^{e]}(x)=0\,,$ (28)
and that it is gauge invariant
$\delta R^{abcd}(x)=0\,.\qquad$ (29)
So, it can be used as physical field.
From (19) we obtain that Ricci and scalar curvatures vanish
$R^{ac}(x)=\eta_{bd}R^{abcd}(x)=0\,,\qquad R(x)=\eta_{ab}R^{ab}(x)=0\,,$ (30)
which produces Einstein equations in vacuum
$G^{ac}(x)=R^{ac}(x)-\frac{1}{2}\eta^{ac}R(x)=0\,.$ (31)
To express coupling of gravitational fields with matter (to a conserved energy-momentum tensor) we must solve Bianchi identity in terms of symmetric tensor $h^{ab}(x)$. The solution is Eq. (23) and then Einstein tensor takes the form (27). Consequently, interaction with matter can be express as in (11).
8.2.4 Dimensions of the representations and degrees of freedom
Let us count dimensions of these representations. As we have already shown dimension of each of representation $T_{\pm 2}^{abcd}(k)$ defined in Eq.(8.2.1) is $d_{2}=28$, where we used the fact that subspace corresponding to $\pi_{0}$ is two dimensional. The first term is $6\cdot 2\cdot 2=24$ dimensional where $6$ is number of combination of $4$ things taken $2$ at a time and $2\cdot 2$ is dimension of subspace corresponding to two operators $\pi_{0}$. The second term is $4$ dimensional so that $24+4=28$. After substitution two dimensional projector $\pi_{0}$ with one dimensional $\pi_{0-}$ we obtain that $T_{\pm 2(-)}^{abcd}(k)$, defined in Eq.(13) is $6\cdot 1\cdot 1+4=10$ dimensional representation, which could be expected for symmetric second rang tensor.
Using symmetry properties the second term in (13) disappears and we are left with $6$ dimensional representation, describing with Eq.(16). Since in that equation we are missing terms with $k^{a}k^{b}$ and $k^{c}k^{d}$, we are left with $4$ dimensional representation. In Eq.(16) both multipliers are antisymmetric. The other three possibilities (that both multipliers are symmetric and that one is symmetric and the other is antisymmetric) complements dimension of such kind of representations to $4$. It means that in our particular case each of representation $T_{+2(-)}^{[ab][cd]}(k)$ and $T_{-2(-)}^{[ab][cd]}(k)$ is one dimensional. Consequently, they are irreducible representations of proper Poincare group with one degree of freedom.
According to discussion above, Riemann tensor $R^{abcd}(k)$ which is irreducible representation of Poincare group with space inversion, has two degrees of freedom.
9 Conclusion
Our basic assumption is that for arbitrary symmetry group all field equations follow from corresponding Casimir eigenvalue equations. In the present article we confirmed that statement for the case of massless Poincare group.
Our systematic approach has verified Weinberg observation [4] that Lorentz invariance is violated with terms in the form of local gauge transformations. The source of this effect is non commutativity of components of Pauli-Lubanski vector. We check this statements for vector fields and for second and fourth rank tensor fields.
Let us stress that helicity is crucial for determining field equations for irreducible components of some field $\Psi^{A}$ and to find the number of theirs degrees of freedom. Note that we can describe the same theory using different fields. For example, as it shows Eq.(8) all representations with helicities $\lambda=\pm 1$ (vector field, symmetric and antisymmetric second rang tensor) have one more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe electromagnetic interaction. Also, all representations with helicities $\lambda=\pm 2$ (symmetric second rang tensor and fourth rang tensor with appropriate symmetries) have two more vector $\breve{p}^{a}_{\pm}$ than $\breve{p}^{a}_{\mp}$ and describe gravitation interaction. Consequently, contrary to the common belief that only highes helicity of certain tensor is relevant, we claim that all helicities are relevant. Different fields with the same helicity can describe the same theory, by sometimes it not easy to write field equation in terms of that fields.
It should be emphasize the important logical sequence in the present article Covariant definition of helicity $\to$ little group invariance $\to$ local gauge transformations $\to$ equations of motion for massless fields. In other words, the necessity of existence of local gauge transformations for massless fields originates from requirement for Lorentz invariance.