### v3

# Poincare field theory for massless particles

Abstract:
Our main proposition is that the field equations for all spins can be derived from the Casimir eigenvalue equations of the Poincaré group. We have already confirmed this claim for massive scalar, spinor, and vector fields in Ref.[1]. In this article, we extend this analysis to massless fields—specifically, the vector field, as well as second and fourth rank tensor fields and finally for arbitrary rank tensor fields. In particular, we will derive the Maxwell equations, the Einstein equations in the weak-field approximation as well as Fronsdal equation.
As is well known, defined particles as irreducible representations of the Poincaré group [2, 3]. However, as Weinberg noted in [4], irreducible representations for massless vector fields with helicities ±$\pm 1$ do not exist. In this article, we confirm this statement for a broad class of massless fields. Unlike their massive counterparts, these fields are not Lorentz invariant because their Lorentz transformations include an additional gauge transformation term. Consequently, such fields can only appear in the theory in forms that are independent of the corresponding gauge parameters. These gauge invariant forms constitute our equations of motion. Thus, the massless case differs fundamentally from the massive one.
Finally, we demonstrate that our approach not only reproduces key results from well-known studies on highest helicity fields but also extends to other helicities.
1 Introduction
In our previous work [1], we derived the equations of motion for massive particles of arbitrary spin by formulating them as eigenvalue equations for the Casimir operators of the Poincaré group. In this paper, we generalize this framework to massless particles, where new subtleties arise due to gauge invariance.
While only one Casimir operator (helicity) exists for massless particles, its covariant definition involves the Pauli-Lubanski vector $W_{a}$, which possesses three independent components (due to the constraint $P^{a}W_{a}=0$). Thus, we obtain three equations, and their compatibility conditions lead to significant distinctions from the massive case.
As in the massive scenario, the fields are functions transforming under a representation of the Poincaré group. For a general representation, we denote these fields as $\Psi^{A}(x)$, where $A$ encompasses the relevant set of indices.
We will work in the standard momentum frame, which significantly simplifies our analysis. First, we solve the eigenvalue problem for the spin operator component $(S_{12})^{A}{}_{B}=\sigma(W_{0})^{A}{}_{B}=(W_{3})^{A}{}_{B}$. The resulting eigenvalues can be interpreted as helicities $\lambda_{i}$ and the corresponding eigenvectors $\Psi^{A}_{i}$ serve as candidates for irreducible representations. However, in certain important cases, the components of the Pauli-Lubanski vector fail to commute. Consequently, no common eigenvectors exist for all components. Moreover, in the standard momentum frame, the operators $W_{1}$ and $W_{2}$ do not annihilate eigenvectors $\Psi^{A}_{i}$, implying that $\Psi^{A}_{i}$ cannot represent genuine irreducible representations of the Poincaré group.
This feature has unexpected consequences, as discovered by Weinberg [4], where the Lorentz transformations of certain physically relevant fields (such as vector and symmetric second-rank tensor fields) acquire an additional term in the form of a gauge transformation. Consequently, for these fields, we can construct actions as gauge-invariant expressions involving the eigenvectors $\Psi^{A}_{i}$.
Thus, gauge-dependent fields are not Lorentz invariant and therefore do not constitute proper representations of the Poincaré group. However, the corresponding field strengths and Lagrangians—being gauge-invariant expressions of the gauge fields—do form representations of the Poincaré group. For this reason, we introduce the term incomplete representation of the Poincaré group (up to gauge transformations) for gauge fields, and incomplete irreducible representation of the Poincaré group for their corresponding irreducible components.
As a consequence, Wigner’s definition of particles [2, 3] as irreducible representations of the Poincaré group, while valid for massive cases, requires refinement in the massless case. Instead, we can define physical massless fields as equivalence classes, where two fields are considered equivalent if they are related by a gauge transformation. The action is then an expression that depends only on the equivalence class.
The main objective of this work is to demonstrate that principle field equations give rise to all known field equations. To this end, we first derive the standard equations for helicities $1$ and $2$. There is also considerable interest in massless fields with higher spin. The theory of particles with highest spin was originally developed by Fierz and Pauli [5]. In Ref. [2] and [3] , it was shown that the one-particle states correspond to irreducible, unitary representations of the Poincaré group.
A minimal set of auxiliary traceless, symmetric tensor fields was later proposed in Ref.[6]. Since having an action is essential for field quantization, the equations of motion and corresponding action were derived in Ref.[7]. Another notable approach, based on a hierarchy of generalized Christoffel symbols, was introduced by de Wit and Freedman [8]. All the above studies rely on the Fierz-Pauli constraints [5] which serve as their foundation. These constraints require that the field be a fully symmetric tensor of rank $n$, while also being traceless, divergence-free, and satisfying the Klein-Gordon equation.
In this work, we present a more fundamental approach to analyzing massless tensor fields. We develop a systematic procedure to determine all possible helicity states and the corresponding tensor structures for any given helicity.
As a special case, for a rank $n$ tensor field, the highest helicity state corresponds to a fully symmetric, traceless, and divergence-free tensor that satisfies the Klein-Gordon equation. Thus, our highest helicity states naturally satisfy the Fierz-Pauli constraints [5]. Beyond the highest helicity, we also investigate non-maximal helicity states and provide their physical interpretation (see Sections 7 and 8 and 12).
2 Principle field equations
In this section, we introduce the fundamental field equations of the theory. These equations are constructed solely from the symmetry group properties and are formulated using the Casimir invariants.
In the present work, we begin with the massless Poincaré group and derive fundamental field equations that hold for arbitrary helicity states.
2.1 Representation for Pauli-Lubanski vector
Casimir invariants are expressions that commute with all group elements. In massless case, where square of momentum vanishes $P^{2}=0$, there are two Casimir operators for Poincare group, the helicity $\lambda$ and sign of $P^{0}$. The covariant definition of helicity is
$W_{a}=\lambda P_{a}\,,\qquad W_{a}=\frac{1}{2}\varepsilon_{abcd}S^{bc}P^{d}\,,$ (1)
where $W_{a}$ is Pauli-Lubanski vector and $S_{ab}$ is spin parts of Lorentz generators. Since $P^{a}W_{a}=0$ we can conclude from (1) that $P^{2}=0$, as it should be for massless case.
Two comments should be made. First, it is possible to use non covariant definition of helicity $\lambda=S_{i}n_{i}$ as projection of space part of spin generator $S_{i}=\frac{1}{2}\,\varepsilon_{ijk}S_{jk}$ to momentum axis $n_{i}=\frac{p_{i}}{|p_{i}|^{2}}$. It produces the same spectrum for $\lambda$ as relation (1). But in that case we are losing the crucial possibility to insist on Lorentz invariance which will product gauge transformations. Second, in definition of Pauli-Lubanski vector orbital part of four dimensional rotation generatr $L_{ab}=x_{a}P_{b}-x_{b}P_{a}$ does not appear since $\varepsilon_{abcd}L^{bc}P^{d}=0$.
For all types of fields representation of momentum is $(P_{a})^{A}{}_{B}\to\delta^{A}_{B}\,i\partial_{a}$. In article [1] we proved relation
$(S_{ab})^{AB}{}_{CD}=(S_{ab})^{A}{}_{C}\delta_{D}^{B}+\delta^{A}_{C}(S_{ab})^{B}{}_{D}\,,$ (2)
which allows us to find representations for arbitrary spin operator using fundamental representations. This initial representation for Dirac spinors and vector fields are
$(S_{ab})^{\alpha}{}_{\beta}=\frac{i}{4}[\gamma_{a},\gamma_{b}]^{\alpha}{}_{\beta}\,,\qquad(S_{ab})^{c}{}_{d}=i\Big{(}\delta^{c}_{a}\eta_{bd}-\delta^{c}_{b}\eta_{ad}\Big{)}\,.$ (3)
For example, the expression for vectors in terms of spinors $V^{a}(x)={\bar{\psi}}(x)\gamma^{a}\psi(x)$ and first relation (3) produces the second one, which can also been obtained by direct calculation.
Using representations for momenta and for spin generators as well as the fact that they commute, we can find representation for Pauli-Lubanski vector.
2.2 Standard momentum $k^{a}$, space inversion and little group
It is useful to work in frame of standard momentum $k^{a}$. In massless case, when Lorentz invariant function of momentum vanish, $p^{2}=0$, we can chose standard momentum in the form $k^{a}=(1,0,0,-1)$. Then we can express any momentum $p^{a}$ as Lorentz transformation of $k^{a}$
$p^{a}=L^{a}{}_{b}(p)k^{b}\,.$ (4)
Up to now we considered proper and orthochronous Lorentz group $\Lambda\in SO(1,3)^{\uparrow}$ which contains the identity matrix and preserves the direction of time and parity ( $\det\Lambda=1$ and $\Lambda^{0}{}_{0}\geq 1$ ). But we will also need space inversion ( $\det\Lambda=-1$ and $\Lambda^{0}{}_{0}\geq 1$ ) which for vector fields is implemented by the operator
${\cal P}^{a}{}_{b}=\delta^{a}_{0}\delta_{b}^{0}-\delta^{a}_{i}\delta_{b}^{i}\,.\qquad(i=1,2,3)$ (5)
In order to follow behavior under space inversion we will introduce new parameter $\sigma$ with inversion rules ${\cal P}\sigma=-\sigma$, and obviously $\sigma^{2}=1$. Then rotations are invariant under parity $S_{ij}\to S_{ij}$ and boosts change the sign $S_{0i}\to\sigma S_{0i}$ while $k^{a}\to(1,0,0,-\sigma)$.
Since $P^{a}W_{a}=0$, Pauli-Lubanski vector has three independent components and for standard momentum they are
$W_{0}=\sigma S_{12}\,,\qquad W_{3}=S_{12}\,,\qquad$
$W_{1}\equiv\Pi_{2}=S_{02}-S_{32}\,,\qquad-W_{2}\equiv\Pi_{1}=S_{01}-S_{31}\,.$ (6)
All three generators annihilate standard momentum. So group element $W^{a}{}_{b}$, leaves $k^{a}$ invariant $W^{a}{}_{b}k^{b}=k^{a}$, which is definition of little group in massless Poincare case.
Commutation relations of little group generators have the form
$[\Pi_{1},\Pi_{2}]=0\,,\qquad[S_{12},\Pi_{1}]=i\Pi_{2}\,,\qquad[S_{12},\Pi_{2}]=-i\Pi_{1}\,.$ (7)
Therefore, as it is well known, in massless case the little group of Poincare group is group $E(2)$, which consists of translations and rotations in two-dimensional Euclidean space.
Note that unlike the massive case here little group contains Abelian subgroup with generators $\Pi_{1}$ and $\Pi_{2}$. Representation of little group must be invariant under this Abelian subgroup. We will see that violation of this feature has important repercussion that gauge invariance is consequence of Lorentz transformation [4].
2.3 Principle field equations
Using the expression (1) along with the momentum and Pauli-Lubanski vector we postulate principle field equations for massless fields $\Psi^{A}(x)$ with helicity $\lambda$
$(W_{a})^{A}{}_{B}\Psi^{B}(x)=\lambda(P_{a})^{A}{}_{B}\Psi^{B}(x)\,.$ (8)
We demonstrate that all massless free field equations with arbitrary helicity in classical field theory [4, 9, 10, 11, 12] can be derived from this fundamental equation. In this work, we explicitly verify this claim for the most physically significant massless fields.
We will investigate principle field equations in momentum representation in the frame of standard momentum. Then differential equations (8) become algebraic equations. After solving this algebraic equations we can go back to $p^{a}$ dependent solutions using (4) and then to coordinate representation.
Using (8) and (2.2) we can obtain principle field equations for standard momentum
$(S_{12})^{A}{}_{B}\Psi^{B}(k)=\lambda\sigma\Psi^{A}(k)\,,$
$(\Pi_{1})^{A}{}_{B}\Psi^{B}(k)=0\,,\qquad(\Pi_{2})^{A}{}_{B}\Psi^{B}(k)=0\,.$ (9)
2.4 Consequences of parity conservation
The operators $(S_{12})^{A}{}_{B}$, $(\Pi_{1})^{A}{}_{B}$ and $(\Pi_{2})^{A}{}_{B}$ are representation of little group generators and $\lambda\sigma$ is a eigenvalue of operator $(S_{12})^{A}{}_{B}$. Under space inversion operator $(S_{12})^{A}{}_{B}$ is invariant which means that helicity change the sign
${\cal P}\lambda=-\lambda\,.$ (10)
Therefore, if the theory is invariant under space inversion then for any non zero helicity it must exist opposite one.
Helicity is invariant under proper orthochronous Poincare group. Particles of opposite helicities $+\lambda$ and $-\lambda$, which are related by transformation of space inversion, corresponds to different species of particles. A theory that conserves parity must treat both helicities states equally. So, both massless states with opposite helicities we can consider as two polarizations of the same particle. This particle is irreducible representation of wider group which combined Poincare group and parity.
The most important examples are electromagnetic and gravitation interactions which are invariant under space inversion. In particular, both massless particles with helicities $\lambda=1$ and $\lambda=-1$ are called photons and both massless particles with helicities $\lambda=2$ and $\lambda=-2$ are called gravitons. General state of massless particle which includes both helicities $\lambda$ and $-\lambda$ is
$\Psi^{A}(x)=\alpha\Psi^{A}_{+\lambda}(x)+\beta\Psi^{A}_{-\lambda}(x)\,,\qquad|\alpha|^{2}+|\beta|^{2}=1\,,$ (11)
and in momentum representation
$\Psi^{A}(p)=\alpha\Psi^{A}_{+\lambda}(p)+\beta\Psi^{A}_{-\lambda}(p)\,.$ (12)
2.5 Real fields in coordinate and momentum representation
In most cases we will work with real fields, such that in coordinate and momentum representation
$(\Psi^{A})^{\ast}(x)=\Psi^{A}(x)\,,\qquad(\Psi^{A})^{\ast}(p)=\Psi^{A}(-p)\,.$ (13)
So for real fields
$\Psi^{A}(x)=\int d^{4}p\Big{(}\Psi^{A}(p)+\Psi^{A}(-p)\Big{)}e^{-ipx}\,.$ (14)
Similarly, for imaginary field we have
$(\Psi^{A})^{\ast}(x)=-\Psi^{A}(x)\,,\qquad(\Psi^{A})^{\ast}(p)=-\Psi^{A}(-p)\,.$ (15)
We will need real part of the expression which in momentum representation has a form $\Psi_{n}^{B}(p)=p^{a_{1}}p^{a_{2}}\cdots p^{a_{n}}\Psi^{B}(p)$
$\Psi_{n}^{B}(x)=\int d^{4}p\Big{[}\Psi_{n}^{B}(p)+\Psi_{n}^{B}(-p)\Big{]}e^{-ipx}=(i)^{n}\partial^{a_{1}}\partial^{a_{2}}\cdots\partial^{a_{n}}\int d^{4}p\Big{[}\Psi^{B}(p)+(-1)^{n}\Psi^{B}(-p)\Big{]}e^{-ipx}\,.\qquad$ (16)
For $n=1$ we have
$\Psi_{1}^{B}(x)=\int d^{4}p\Big{[}p^{a}\Psi^{B}(p)-p^{a}\Psi^{B}(-p)\Big{]}e^{-ipx}=\partial^{a}\Omega_{1}^{B}(x)\,,\qquad\qquad$ (17)
and for $n=2$
$\Psi_{2}^{B}(x)=\int d^{4}p\Big{[}p^{a}p^{b}\Psi^{B}(p)+p^{a}p^{b}\Psi^{B}(-p)\Big{]}e^{-ipx}=\partial^{a}\partial^{b}\Omega_{2}^{B}(x)\,,$ (18)
where
$\Omega_{1}^{B}(x)=i\int d^{4}p\Big{[}\Psi^{B}(p)-\Psi^{B}(-p)\Big{]}e^{-ipx}\,,\qquad\Omega_{2}^{B}(x)=-\int d^{4}p\Big{[}\Psi^{B}(p)+\Psi^{B}(-p)\Big{]}e^{-ipx}\,.\qquad\qquad$ (19)
According to (15) and (13) both $\Omega_{1}^{B}(x)$ and $\Omega_{2}^{B}(x)$ are real functions.
2.6 Solution of $(S_{12})^{A}{}_{B}$ equation
We are going to solve eigenproblem of operator $(S_{12})^{A}{}_{B}$ which will produces spectrum of helicities $\lambda_{i}$ $(i=1,2,\cdots,n)$. Then we will find projection operators $(P_{i})^{A}{}_{B}$ and eigenfunctions $\Psi_{i}^{A}$ corresponding to these eigenvalues.
In order that first equation (2.3) has nontrivial solutions its characteristic polynomial must vanish
$P_{S_{12}}(\lambda\sigma)=\det\Big{(}(S_{12})^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}\Big{)}=0\,.$ (20)
The zeros of characteristic polynomial are eigenvalues $\lambda_{i}$.
Next, we can construct projection operators $(P_{i})^{A}{}_{B}$ on subspaces $\Xi_{i}$ (we will call it eigenspace) with dimension $d_{i}=dim\Xi_{i}$ corresponding to helicities $\lambda_{i}$
$(P_{i})^{A}{}_{B}=\frac{\Big{[}\prod_{j\neq i}^{n}\Big{(}S_{12}-\lambda_{j}\sigma\delta\Big{)}\Big{]}^{A}{}_{B}}{\prod_{j\neq i}^{n}\Big{(}\lambda_{i}\sigma-\lambda_{j}\sigma\Big{)}}\,,$ (21)
and obtain corresponding eigenfunctions
$\Psi_{i}^{A}(k)=(P_{i})^{A}{}_{B}\Psi^{B}(k)\,.$ (22)
From (21) and (22) it follows that space inversion connect projectors and eigenfunctions with opposite helicities
${\cal P}\,(P_{\lambda_{i}})^{A}{}_{B}=(P_{-\lambda_{i}})^{A}{}_{B}\,,\qquad{\cal P}\,\Psi_{\lambda_{i}}^{A}(k)=\Psi_{-\lambda_{i}}^{A}(k)\,.$ (23)
Generally, (see section 9 for details) with the help of projection operators, for tensor fields of rank $n$ we can rewrite equation (20) in the form
$\det(S_{12}-\lambda\sigma\delta)=\prod_{i=-n}^{n}(i-\lambda\sigma)^{\tilde{d}_{i}}\det P_{i}=0\,,\qquad$ (24)
and using $\sigma^{2}=1$ and $\tilde{d}_{i}=\tilde{d}_{-i}$
$\det(S_{12}-\lambda\sigma\delta)=(-\lambda\sigma)^{\tilde{d}_{0}}\prod_{i=1}^{n}(i^{2}-\lambda^{2})^{\tilde{d}_{i}}\,\,\prod_{i=-n}^{n}\det P_{i}=0\,.\qquad$ (25)
Exponents in multipliers $\tilde{d}_{i}$ are algebraic multiplicity while $d_{i}$ as dimensions of eigenspaces $\Xi_{i}$ are geometric multiplicity. Generally, $d_{i}\leq\tilde{d}_{i}$, but in our case $d_{i}=\tilde{d}_{i}$ so that from now on we will use a unique notation $d_{i}$.
Note that for tensor fields of rank $n$ relation $\sum_{i=-n}^{n}d_{i}=d_{0}+2\sum_{i=1}^{n}d_{i}=4^{n}$ is satisfied and spectrum of helicities is
$\lambda_{0}=0\,,\qquad\lambda_{\pm i}=\pm i\,,\qquad i=(1,2,\cdots,n)\,.$ (26)
Let us stress that gauge invariant fields $\Psi_{i}^{A}(k)$ (see the next subsection for explanation) are representations of Poincare group, but in general case they are not irreducible representations. To get irreducible representations we must separate fields $\Psi_{i}^{A}(k)$ using symmetry properties. Note that dimension of subspace $d_{i}=dim\Xi_{i}$ corresponding to helicity $\lambda_{i}$ is in fact number of degrees of freedom of fields $\Psi_{i}^{A}(k)$. Similarly, dimension of subspace of irreducible representation is number of degrees of freedom of corresponding fields. We are going to check it in all examples.
2.7 Lorentz transformations induce gauge transformations
We solved first equation (2.3) and it produces complete set of eigenvalues and eigenfunctions. But we have additional two conditions. It is of great importance to investigate the compatibility of all equations (2.3).
2.7.1 Gauge dependent fields
In equations for standard momentum (2.3) operators $\Pi_{1}$ and $\Pi_{2}$ should annihilate field $\Psi^{A}(k)$. But explicit calculation shows that this does not happen in some physically relevant cases. Since equations (2.3) are condition for Lorentz invariance, violation of these conditions breaks Lorentz invariance. To measure this violation we introduce expression
$\delta\Psi_{i}^{A}(\varepsilon_{1},\varepsilon_{2})(k)=i\Big{(}\varepsilon_{1}\Pi_{1}+\varepsilon_{2}\Pi_{2}\Big{)}^{A}{}_{B}\Psi_{i}^{B}(k)\,,\qquad$ (27)
where $\varepsilon_{1}$ and $\varepsilon_{2}$ are some parameters.
Field equations must be Lorentz invariant and consequently they should not depend on $\delta\Psi_{i}^{A}(\varepsilon_{1},\varepsilon_{2})(k)$. It means that
$\delta\Psi_{i}^{A}(\varepsilon_{1},\varepsilon_{2})(k)={\cal G}_{i}\Big{(}k,\Omega_{\pm}\Big{)}\,,\qquad\Omega_{\pm}=\Omega_{\pm}(\varepsilon_{\pm})$ (28)
are gauge transformations of fields $\Psi_{i}^{A}(k)$ with gauge parameters $\Omega_{\pm}(\varepsilon_{\pm})$. So, as Weinberg noted Lorentz transformations induce gauge transformations [4].
We can perform transition from standard momentum $k^{a}$ to arbitrary one $p^{a}$ and using relation $p^{a}=i\partial^{a}$ we can obtain gauge transformation in coordinate space
$\delta\Psi_{i}^{A}(\Omega_{\pm},x)={\cal G}_{i}\Big{(}i\partial,\Omega_{\pm}(x)\Big{)}\,.$ (29)
Finally, we can chose the action
$S(\Psi_{i})=\int d^{4}x{\cal L}\Big{(}\Psi_{i}^{A}(x)\Big{)}\,,$ (30)
as a gauge invariant expressions of $\Psi_{i}^{A}(x)$, which means that following condition must be satisfied
$S\Big{(}\Psi_{i}+\delta\Psi_{i}\Big{)}=S\Big{(}\Psi_{i}\Big{)}\,.$ (31)
Fields which violate Lorentz invariance are gauge dependent fields and we will simply call them gauge fields. They are not representation of Poincare group. If we want to weaken the statement they are representation of Poincare group up to gauge transformations, since corresponding field strength and Lagrangian are representations of Poincare group. Therefore, for these fields we will introduce name incomplete representation of Poincare group.
2.7.2 Gauge invariant fields
Gauge invariant fields are representation of Poincare group. So, they are physical and we can use them without restrictions. We can obtain their form as solutions of equations (2.3). That leads us to field equations of gauge invariant fields as some constraints on the obtained solution.
Next we will consider the most important physical cases of massless fields: scalar field, vector field with helicities $\lambda=\pm 1$ (photon) and symmetric second rank tensor with helicities $\lambda=\pm 2$ (metric). We will also consider second rank tensor with helicities $\lambda=\pm 1$ (field strength) and fourth rank tensor with helicities $\lambda=\pm 2$ (Riemann curvature). We will also comment gauge fields with higher helicities.
3 Scalar field
The simplest example is scalar field $\varphi(x)$. In that case spin operator $S_{ab}$ vanishes. In particular $S_{12}=0$ which means that helicity is zero, $\lambda=0$. From $\Pi_{1}=0=\Pi_{2}$ we obtain $\delta\varphi(\varepsilon_{1},\varepsilon_{2})(x)=0$. Therefore, massless scalar field $\varphi(x)$ is gauge invariant, and we can take it as physical field without restrictions. It is one dimensional irreducible representation of Poinace group and consequently it has one degree of freedom. Equation of motion is massless Klein-Gordon equation $\partial^{2}\varphi(x)=0$.
4 Vector field and Maxwell equations
The case of massless vector field is most important. Besides representing electromagnetic interaction we will see that solutions for massless tensor fields of arbitrary rang comes down to the case of vector field.
4.1 Eigenvalues and projection operators
For vector field we have $A,B\to a,b$, $\Psi^{A}\to V^{a}$ and
$(S_{ab})^{A}{}_{B}\to\Big{(}S_{ab}\Big{)}^{c}{}_{d}=i\Big{(}\delta^{c}_{a}\,\eta_{bd}-\delta^{c}_{b}\,\eta_{ad}\Big{)}\,,\qquad\Rightarrow\qquad(S_{12})^{a}{}_{b}=i\Big{(}\delta^{a}_{1}\eta_{2b}-\delta^{a}_{2}\eta_{1b}\Big{)}\,.$ (1)
The consistency condition requires that characteristic polynomial vanishes. It means
$\det\Big{(}{\cal S}_{12}-\lambda\sigma\Big{)}^{a}{}_{b}=\lambda^{2}(\lambda-1)(\lambda+1)=0\,,$ (2)
with solutions for eigenvalues
$\lambda_{0}=0\,,\qquad\lambda_{\pm 1}=\pm 1\,.\qquad$ (3)
We note that eigenspace $\Xi_{0}$ corresponding to the eigenvalue $\lambda_{0}=0$ is two dimensional $d_{0}=dim\Xi_{0}=2$ implying that its associated eigenvector carries two degrees of freedom. Similarly, the eigenspaces $\Xi_{\pm 1}$ corresponding to eigenvalues $\lambda_{\pm 1}=1$ respectively are both one dimensional $d_{\pm 1}=dim\Xi_{\pm 1}=1$. Consequently, each of their associated eigenvectors carries a single degree of freedom.
We already had general expression for projection operators (21) which for vector fields, where $n=3$, takes the form
$(P_{i})^{a}{}_{b}=\frac{\prod_{j\neq i}^{3}\Big{(}(S_{12})^{a}{}_{b}-\lambda_{j}\sigma\delta^{a}_{b}\Big{)}}{\prod_{j\neq i}^{3}\Big{(}\lambda_{i}\sigma-\lambda_{j}\sigma\Big{)}}\,,$ (4)
or explicitly
$(P_{0})^{a}{}_{b}(k)=\delta^{a}_{b}-(S_{12}^{2})^{a}{}_{b}=\delta^{a}_{b}-\delta^{a}_{\alpha}\delta_{b}^{\alpha}\,,\qquad(\alpha=1,2)$
$(P_{\pm 1})^{a}{}_{b}(k)=\frac{1}{2}\Big{(}(S_{12}^{2})^{a}{}_{b}\pm\sigma(S_{12})^{a}{}_{b}\Big{)}=\frac{1}{2}\Big{[}\delta^{a}_{\alpha}\delta_{b}^{\alpha}\pm i\sigma\Big{(}\delta^{a}_{1}\eta_{2b}-\delta^{a}_{2}\eta_{1b}\Big{)}\Big{]}\,.$ (5)
4.2 Basic vectors
For further benefits we introduce basic vectors $k^{a}$ and $q^{a}$ in two dimensional eigenspace $\Xi_{0}$. Then complete set of basic vectors $e^{a}_{i}=\{k^{a},q^{a},\breve{p}_{+}^{a},\breve{p}_{-}^{a}\}$ where $i=\{0+,0-,+1,-1\}$ are
$k^{a}=\delta^{a}_{0}-\sigma\delta^{a}_{3}\,,\quad q^{a}=\delta^{a}_{0}+\sigma\delta^{a}_{3}\,,\quad\breve{p}_{\pm}^{a}=\delta^{a}_{1}\pm i\sigma\delta^{a}_{2}\,.\qquad$ (6)
It is easy to check that
$(S_{12})^{a}{}_{b}e^{b}_{i}=\lambda_{i}e^{a}_{i}\,,$ (7)
where $\lambda_{i}$ take the values from (3). Consequently, basic vector $e^{a}_{i}$ carries helicity $\lambda_{i}$. In particular, basic vectors $\breve{p}_{\pm}^{a}$ carry helicities ±$\pm 1$ and both basic vectors $k^{a}$ and $q^{a}$ carry helicities 00. If $n_{\pm}$ are numbers of vectors $\breve{p}_{\pm}^{a}$ in the representation of field $\Psi^{A}$, then its helicity is
$\lambda=n_{+}-n_{-}\,.$ (8)
[…]
Projection operators in terms of basic vectors take the form
$(P_{0+})^{a}{}_{b}(k)=\frac{1}{2}q^{a}k_{b}\,,\qquad(P_{0-})^{a}{}_{b}(k)=\frac{1}{2}k^{a}q_{b}\,,\qquad(P_{\pm 1})^{a}{}_{b}(k)=-\frac{1}{2}\breve{p}_{\pm}^{a}(\breve{p}_{\mp})_{b}\,.$ (9)
We can check that above expressions are really projectors using product of basic vectors where only nontrivial products are
$k^{a}q_{a}=2\,,\qquad\breve{p}_{\pm}^{a}(\breve{p}_{\mp})_{a}=-2\,.$ (10)
Note that all basic vectors are light-like.
4.3 Eigenfunctions of operator $(S_{12})^{a}{}_{b}$
The eigenfunctions of operator $(S_{12})^{a}{}_{b}$ are $V_{i}^{a}(k)=(P_{i})^{a}{}_{b}V^{b}$ where $i=\{0+,0-,+1,-1\}$, or explicitly
$V_{0+}^{a}(k)=\frac{1}{2}q^{a}k_{b}V^{b}\equiv q^{a}\mathbf{V}_{k}\,,\qquad V_{0-}^{a}(k)=\frac{1}{2}k^{a}q_{b}V^{b}\equiv k^{a}\mathbf{V}_{q}\,,\qquad$
$V_{\pm 1}^{a}(k)=-\frac{1}{2}\breve{p}_{\pm}^{a}(\breve{p}_{\mp})_{b}V^{b}\equiv\breve{p}_{\pm}^{a}\mathbf{V}_{\mp}\,.\qquad$ (11)
It is easy to check that relation (8) is satisfied since scalars like $\mathbf{V}_{\mp}=(\breve{p}_{\mp})_{b}V^{b}$ carry helicity 00.
The basic vectors determines the characteristics of eigenvectors. They have the same helicity and as we will see the same gauge transformations and the same parity. Here all eigenfunctions $V_{i}^{a}(k)$ are potential representations of Poincare group, but only gauge invariant ones are really representations of Poincare group.
Under space inversion vectors $k^{a}$ and $q^{a}$ are invariant while
${\cal P}\breve{p}_{\pm}^{a}=-\breve{p}_{\mp}^{a}\,.$ (12)
It confirms that general statements (23) valid for vector fields
${\cal P}(P_{\pm 1})^{a}{}_{b}(k)=(P_{\mp 1})^{a}{}_{b}(k)\,,\qquad{\cal P}V_{\pm 1}^{a}(k)=V_{\mp 1}^{a}(k)\,.$ (13)
The vector field $V^{a}$ has four components. There are four corresponding eigenfunctions $V^{a}_{i}$ with four components. But they are not independent. In fact, if we multiply all components $V^{a}_{i}$ with all basic vectors we obtain only four nonzero expressions
$k_{a}V_{0+}^{a}=k_{b}V^{b}=2\mathbf{V}_{k}\,,\qquad q_{a}V_{0-}^{a}=q_{b}V^{b}=2\mathbf{V}_{q}\,,\qquad(\breve{p}_{\mp})_{a}V_{\pm 1}^{a}=(\breve{p}_{\mp})_{b}V^{b}=-2\mathbf{V}_{\mp}\,.\qquad$ (14)
4.4 Gauge transformations of massless vector fields
Gauge transformations of eigenfunctions of operator $(S_{12})^{a}{}_{b}$ for massless vector fields
$\delta V_{i}^{a}(\varepsilon_{1},\varepsilon_{2})(k)=i\Big{(}\varepsilon_{1}\Pi_{1}+\varepsilon_{2}\Pi_{2}\Big{)}^{a}{}_{b}V_{i}^{b}(k)\,,\qquad$ (15)
are particular case of Eq.(27). Explicitly, using expressions
$(\Pi_{1})^{a}{}_{b}=i\Big{(}\delta^{a}_{0}\eta_{1b}-\delta^{a}_{1}\eta_{0b}+\delta^{a}_{1}\eta_{3b}-\delta^{a}_{3}\eta_{1b}\Big{)}\,,\qquad(\Pi_{2})^{a}{}_{b}=i\Big{(}\delta^{a}_{0}\eta_{2b}-\delta^{a}_{2}\eta_{0b}+\delta^{a}_{2}\eta_{3b}-\delta^{a}_{3}\eta_{2b}\Big{)}\,,\qquad$ (16)
we get
$\delta V_{i}^{a}(\varepsilon_{1},\varepsilon_{2})=-\Big{[}\varepsilon_{1}\Big{(}\delta^{a}_{0}\eta_{1b}-\delta^{a}_{1}\eta_{0b}+\delta^{a}_{1}\eta_{3b}-\delta^{a}_{3}\eta_{1b}\Big{)}+\varepsilon_{2}\Big{(}\delta^{a}_{0}\eta_{2b}-\delta^{a}_{2}\eta_{0b}+\delta^{a}_{2}\eta_{3b}-\delta^{a}_{3}\eta_{2b}\Big{)}\Big{]}V_{i}^{b}$
$=\left|{\begin{array}[]{cccc}0&\varepsilon_{1}&\varepsilon_{2}&0\\ \varepsilon_{1}&0&0&\varepsilon_{1}\\ \varepsilon_{2}&0&0&\varepsilon_{2}\\ 0&-\varepsilon_{1}&-\varepsilon_{2}&0\\ \end{array}}\right|\left|{\begin{array}[]{c}V^{0}_{i}\\ V^{1}_{i}\\ V^{2}_{i}\\ V^{3}_{i}\\ \end{array}}\right|\,.\qquad\qquad\qquad$ (25)
In particular, the gauge transformations of basic vectors have simple form
$\delta k^{a}=0\,,\qquad\delta q^{a}=\varepsilon_{+}\breve{p}_{-}^{a}+\varepsilon_{-}\breve{p}_{+}^{a}\,,\qquad\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}\,.\qquad(\varepsilon_{\pm}=\varepsilon_{1}\pm i\sigma\varepsilon_{2})\qquad$ (26)
The four independent scalars (14) are gauge invariant. So, gauge transformations of eigenfunctions $V_{i}^{a}(k)=(P_{i})^{a}{}_{b}V^{b}(k)$ are determined by gauge transformations of basic vectors.
Let us stress that only component $V_{0-}^{a}(k)$ is gauge invariant, $\delta V_{0-}^{b}(k)=0$, and consequently only this part is irreducible representation of Poincae group. The other components $V_{0+}^{a}(k)$ and $V_{\pm 1}^{b}(k)$ are not gauge invariant and they transform as
$\delta V_{0+}^{a}(k)=\omega_{+}\breve{p}_{-}^{a}+\omega_{-}\breve{p}_{+}^{a}\,,\qquad\omega_{\pm}(k)=\frac{1}{2}k_{a}V^{a}\varepsilon_{\pm}\,,$
$\delta V_{\pm 1}^{a}(k)=k^{a}\Omega_{\pm}\,,\qquad\Omega_{\pm}(k)=-\frac{1}{2}(\breve{p}_{\mp})_{b}V^{b}\varepsilon_{\pm}\,.$ (27)
From (26) we can conclude that under space inversion ${\cal P}\varepsilon_{\pm}=-\varepsilon_{\mp}$. It produces ${\cal P}\omega_{\pm}=-\omega_{\mp}$ and ${\cal P}\Omega_{\pm}=\Omega_{\mp}$. It is in agreement with relations ${\cal P}V_{0+}^{a}(k)=V_{0+}^{a}(k)$ and ${\cal P}V_{\pm}^{a}(k)=V_{\mp}^{a}(k)$.
4.5 Massless vector field with helicity $\lambda=0$
Component $V_{0+}^{a}(k)$ is not gauge invariant but from its gauge transformation (4.4) we can conclude that scalar $\varphi_{1}=k_{a}V_{0+}^{a}(k)$ is gauge invariant. Component $V_{0-}^{a}(k)$ is already gauge invariant and unique scalar which we can construct from this vector is $\varphi_{2}=q_{a}V_{0-}^{a}(k)$. These two scalars are first two invariants in (14). Since they are gauge invariant there are no restrictions on scalar fields $\varphi_{1}$ and $\varphi_{2}$ and we can use them as physical fields. These two fields fill two dimensional eigenspace with $\Xi_{0}$ and they have together two degrees of freedom. They satisfy massless Klein-Gordon equation.
4.6 Action for massless vector field with helicities $\lambda=\pm 1$ produces Maxwell equations
The electromagnetic interaction is symmetric under space inversion. Therefore, instead of two components $V_{+1}^{a}(k)$ and $V_{-1}^{a}(k)$ which are incomplete irreducible representation of proper Poincare group we will introduce one vector field $A^{a}(k)$ which is incomplete irreducible representation of wider group which includes space inversion. Using (11) we obtain
$A^{a}(k)=\alpha V^{a}_{+1}(k)+\beta V^{a}_{-1}(k)\,.\qquad$ (28)
We also introduce one parameter $\Omega(k)=\alpha\Omega_{+}(k)+\beta\Omega_{-}(k)$ instead of two components $\Omega_{+}(k)$ and $\Omega_{-}(k)$ related by space inversion.
As we have already seen representations $V_{+1}^{a}(k)$ and $V_{-1}^{a}(k)$ are one dimensional which means that each of them has one degree of freedom. So, field $A^{a}(k)$ has two degrees of freedom and as we will see it describes the photon.
Now, we are ready to introduce action as gauge invariant expressions of $A^{a}(k)$. First we perform transition from standard momentum $k^{a}$ to arbitrary one $p^{a}$ and then to coordinate dependent fields using $p^{a}=i\partial^{a}$. For real field $A^{a}(x)$, according to (14) and (28) we have
$A^{a}(x)=\int d^{4}p\Big{(}A^{a}(p)+A^{a}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}V^{a}_{+1}(p)+V^{a}_{+1}(-p)\Big{)}+\beta\Big{(}V^{a}_{-1}(p)+V^{a}_{-1}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (29)
With the help of second line in (4.4) we obtain
$\delta A^{a}(x)=\int d^{4}p\,p^{a}\Big{[}\alpha\Big{(}\Omega_{+1}(p)-\Omega_{+1}(-p)\Big{)}+\beta\Big{(}\Omega_{-1}(p)-\Omega_{-1}(-p)\Big{)}\Big{]}e^{-ipx}\,,$ (30)
and using (17) and (19)
$\delta A^{a}(x)=\partial^{a}\Big{(}\alpha\Omega_{+1}(x)+\beta\Omega_{-1}(x)\Big{)}=\partial^{a}\Omega(x)\,.$ (31)
Action should depend only on the gauge invariant combination of vector fields, in fact of field strength $F_{ab}=\partial_{a}A_{b}-\partial_{b}A_{a}$. Since it must be scalar we can take
$I_{0}=-\frac{1}{4}\int d^{4}xF_{ab}F^{ab}\,.$ (32)
It is not obvious that field equation, $\partial^{2}A^{a}-\partial^{a}\partial_{b}A^{b}=0$, which follows from (32) has two degrees of freedom. The gauge transformation (31) has one parameter. We can use Lorentz gauge fixing $\partial_{a}A^{a}=0$ which kills one component of vector field and after that equation of motion takes the form $\partial^{2}A^{a}=0$. But now both gauge fixing and field equation are invariant under additional special gauge transformation $\delta A^{a}=\partial^{a}\Omega_{0}$, such that $\partial^{2}\Omega_{0}=0$. Corresponding gauge fixing kills additional degree of freedom and consequently electromagnetic field $A^{a}(x)$ has two degrees of freedom.
Interaction with other fields can be described by the action
$I_{int}\Big{(}A^{a}\Big{)}=\int d^{4}xA^{a}J_{a}\,,$ (33)
with requirement that it is gauge invariant $I_{int}\Big{(}A^{a}+\delta A^{a}\Big{)}=I_{int}\Big{(}A^{a}\Big{)}$. After partial integration we obtain $\int d^{4}x\Omega\,\partial^{a}J_{a}=0$ and for arbitrary $\Omega$ it produces condition on current to be conserved $\partial^{a}J_{a}=0$.
The complete action is
$I=I_{0}+I_{int}=\int d^{4}x\Big{(}-\frac{1}{4}F_{ab}F^{ab}+A^{a}J_{a}\Big{)}\,.$ (34)
Variation with respect to $A^{a}$ produces Maxwell equations.
5 Second rank tensor
In this section we will consider general features of second rank tensor.
5.1 Projection operators
In the case of second rank tensor we have $A\to(ab)$ and $\Psi^{A}\to T^{ab}$. We will keep notations $(P_{i})^{ab}{}_{cd}$ for projectors and $(S_{12})^{ab}{}_{cd}$ for spin operator of second ranktensor. Since we will use abbreviated notation we introduce new notations for vector operators, $(P_{0})^{a}{}_{b}=(\pi_{0})^{a}{}_{b}$ and $(P_{\pm 1})^{a}{}_{b}=(\pi_{\pm})^{a}{}_{b}$ for projectors and $(S_{12})^{a}{}_{b}=\rho^{a}{}_{b}$ for spin operator.
According to (2) in abbreviated notation representation of spin operator for second rank tensor takes a form
$(S_{12})^{ab}{}_{cd}=\rho^{a}{}_{c}\delta^{b}_{d}+\delta^{a}_{c}\rho^{b}{}_{d}=\Big{(}\rho\delta+\delta\rho\Big{)}^{ab}{}_{cd}\,,\qquad$ (1)
where $\rho^{a}{}_{b}$ has been defined in (1). Note that multiplication in abbreviated notation means that we multiply first term with first term and second term with second term $(A\ast B)(C\ast D)=(AC\ast BD)$.
Using relations $\rho=\pi_{+}-\pi_{-}$ and $\delta=\pi_{0}+\pi_{+}+\pi_{-}$ we obtain (for simplicity we will omit indices)
$S_{12}=P_{1}+2P_{2}-P_{-1}-2P_{-2}\,,\qquad\delta\delta=P_{1}+P_{2}+P_{-1}+P_{-2}+P_{0}\,,\qquad$ (2)
where
$P_{0}=\pi_{0}\pi_{0}+\pi_{+}\pi_{-}+\pi_{-}\pi_{+}\,,\qquad P_{\pm 1}=\pi_{0}\pi_{\pm}+\pi_{\pm}\pi_{0}\,,\qquad P_{\pm 2}=\pi_{\pm}\pi_{\pm}\,.\qquad$ (3)
Using the fact that $\pi_{0},\pi_{+}$ and $\pi_{-}$ are projectors we can conclude that $(P_{i})^{ab}{}_{cd}$ $(i=0,\pm 1,\pm 2)$ are projectors also.
5.2 Solutions of $S_{12}$ equation
For second rank tensor characteristic matrix takes the form
$S_{12}-\lambda\sigma\delta\delta=-\lambda\sigma P_{0}+(1-\lambda\sigma)P_{1}+(2-\lambda\sigma)P_{2}-(1+\lambda\sigma)P_{-1}-(2+\lambda\sigma)P_{-2}\,.\qquad$ (4)
Since $(P_{i})^{ab}{}_{cd}$ are projectors, characteristic polynomial is already factored and we have
$\det\Big{(}S_{12}-\lambda\sigma\delta\delta\Big{)}=\Big{(}-\lambda\sigma\Big{)}^{d_{0}}\Big{(}1-\lambda\sigma\Big{)}^{d_{1}}\Big{(}1+\lambda\sigma\Big{)}^{d_{-1}}\Big{(}2-\lambda\sigma\Big{)}^{d_{2}}\Big{(}2+\lambda\sigma\Big{)}^{d_{-2}}\prod_{i=-2}^{2}\det P_{i}=0\,,\qquad\qquad$ (5)
where $d_{i}$ are dimensions of eigenspaces $\Xi_{i}$. Dimension of subspace projected by $\pi_{0}$ is $2$ while dimensions of subspaces projected by $\pi_{+}$ and $\pi_{-}$ are $1$, so that in our case $d_{0}=6\,,\,d_{\pm 1}=4\,,\,d_{\pm 2}=1$. It is in agreement with general relation since $6+2(4+1)=4^{2}$.
Then with the help of relation $\sigma^{2}=1$ we obtain
$\det\Big{(}S_{12}-\lambda\sigma\delta\delta\Big{)}=\Big{(}-\lambda\Big{)}^{6}\Big{(}1-\lambda^{2}\Big{)}^{4}\Big{(}4-\lambda^{2}\Big{)}\prod_{i=-2}^{2}\det P_{i}=0\,.\qquad$ (6)
Using the fact that $\det P_{i}\neq 0$ we can find that spectrum of helicities is
$\lambda_{0}=0\,,\qquad\lambda_{\pm 1}=\pm 1\,,\qquad\lambda_{\pm 2}=\pm 2\,.\qquad$ (7)
Consequently, the eigenfunctions of operator $S_{12}$ in full notation have the form
$T_{i}^{ab}(k)=(P_{i})^{ab}{}_{cd}T^{cd}(k)\,,\qquad(i=0,\pm 1,\pm 2)$ (8)
where projection operators $(P_{i})^{ab}{}_{cd}$ are defined in (3).
6 Second rank tensor with helicities $\lambda=\pm 2$ and general relativity in weak field approximation
Second rank tensor with helicities $\lambda=\pm 2$ is of particular importance since its symmetric part (metric) contribute to general relativity.
6.1 Gauge transformation for highest helicities $\lambda=\pm 2$ of second rank tensors
Each component of the second rank tensor with highest helicities $\lambda=\pm 2\,$
$T_{\pm 2}^{ab}(k)=(P_{\pm 2})^{ab}{}_{cd}T^{cd}(k)=(\pi_{\pm 1})^{a}{}_{c}(\pi_{\pm 1})^{b}{}_{d}T^{cd}(k)=\frac{1}{4}(\breve{p}_{\mp})_{c}(\breve{p}_{\mp})_{d}T^{cd}(k)\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}\,,\qquad$ (1)
is one dimensional representation and consequently it has one degree of freedom. Note that by definition it is symmetric tensor and relation (8) is satisfy.
With the help of (2) we can express little group generators for massless second rank tensors $(\Pi_{1})^{ab}{}_{cd}$ and $(\Pi_{2})^{ab}{}_{cd}$ in terms of corresponding generators for vectors $(\Pi_{1})^{a}{}_{b}$ and $(\Pi_{2})^{a}{}_{b}$
$(\Pi_{1})^{ab}{}_{cd}=(\Pi_{1})^{a}{}_{c}\delta^{b}_{d}+\delta^{a}_{c}(\Pi_{1})^{b}{}_{d}\,,\qquad(\Pi_{2})^{ab}{}_{cd}=(\Pi_{2})^{a}{}_{c}\delta^{b}_{d}+\delta^{a}_{c}(\Pi_{2})^{b}{}_{d}\,.$ (2)
In that case instead of general expression for gauge transformation (27) we have
$\delta T_{i}^{ab}(\varepsilon_{1},\varepsilon_{2})(k)=i\Big{(}\varepsilon_{1}\Pi_{1}+\varepsilon_{2}\Pi_{2}\Big{)}^{ab}{}_{cd}\,T_{i}^{cd}(k)\,.\qquad$ (3)
Using expressions for highes helicities (1) and gauge transformation for basic vector $\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}$ we obtain
$\delta T_{\pm 2}^{ab}(k)=k^{a}\Omega_{\pm}^{b}+k^{b}\Omega_{\pm}^{a}\,,\qquad\Omega_{\pm}^{a}(k)=\frac{1}{4}(\breve{p}_{\mp})_{c}(\breve{p}_{\mp})_{d}T^{cd}\breve{p}_{\pm}^{a}\varepsilon_{\pm}\,.\qquad$ (4)
Note that again ${\cal P}T_{\pm 2}^{ab}(k)=T_{\mp 2}^{ab}(k)$ and ${\cal P}\Omega_{\pm}^{a}=\Omega_{\mp}^{a}$.
Here we came to a similar conclusion for highest helicities of second rank tensor as in the case of vector field: it is not physical since it is not Lorentz invariant.
6.2 General relativity in weak field approximation
As in the case of vector fields, components of massless second rank tensors of opposite helicities $\lambda=2$ and $\lambda=-2$ are related by transformation of space inversion. Since gravitation interaction is symmetric under space inversion it is common to treat both components of incomplete irreducible representation of proper Poincare group as a single particle called graviton. So, instead of two polarizations $T_{+2}^{ab}(k)$ and $T_{-2}^{ab}(k)$, both with one degree of freedom, we will introduce one symmetric tensor $h^{ba}(k)=h^{ab}(k)$ with two degrees of freedom, which is incomplete irreducible representation of Poincare group with space inversion. Using (11) we obtain
$h^{ab}(k)=\alpha T^{ab}_{+2}(k)+\beta T^{ab}_{-2}(k)\,.\qquad$ (5)
We also introduce one parameter $\Omega^{a}(k)=\alpha\Omega_{+}^{a}(k)+\beta\Omega_{-}^{a}(k)$ instead of two components $\Omega_{+}^{a}(k)$ and $\Omega_{-}^{a}(k)$ related by space inversion.
If we go from standard momentum frame to arbitrary frame and then to coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, we can obtain gauge transformation in coordinate representation. For real field, according to (14) and (5) we have
$h^{ab}(x)=\int d^{4}p\Big{(}h^{ab}(p)+h^{ab}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}T^{ab}_{+2}(p)+T^{ab}_{+2}(-p)\Big{)}+\beta\Big{(}T^{ab}_{-2}(p)+T^{ab}_{-2}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (6)
With the help of (4) we obtain gauge transformation
$\delta h^{ab}(x)=\int d^{4}p\,\Big{\{}p^{a}\Big{[}\alpha\Big{(}\Omega_{+}^{b}(p)-\Omega_{+}^{b}(-p)\Big{)}+\beta\Big{(}\Omega_{-}^{b}(p)-\Omega_{-}^{b}(-p)\Big{)}\Big{]}$
$+p^{b}\Big{[}\alpha\Big{(}\Omega_{+}^{a}(p)-\Omega_{+}^{a}(-p)\Big{)}+\beta\Big{(}\Omega_{-}^{a}(p)-\Omega_{-}^{a}(-p)\Big{)}\Big{]}\Big{\}}e^{-ipx}\,,$ (7)
and using (17) and (19)
$\delta h^{ab}(x)=\partial^{a}\Big{(}\alpha\Omega_{+}^{b}(x)+\beta\Omega_{-}^{b}(x)\Big{)}+\partial^{b}\Big{(}\alpha\Omega_{+}^{a}(x)+\beta\Omega_{-}^{a}(x)\Big{)}=\partial^{a}\Omega^{b}(x)+\partial^{b}\Omega^{a}(x)\,.\qquad$ (8)
In order to get field equation we need gauge invariant function of tensor $h^{ab}(x)$. It is linear combination of massless Klein-Gordon equation and supplementary conditions $G^{b}$
$\partial^{2}h_{ab}=0\,,\qquad G^{b}=\partial_{a}h^{ab}-\frac{1}{2}\partial^{b}h^{a}{}_{a}=0\,.$ (9)
The supplementary conditions are in fact gauge fixing which kill all unwanted degrees of freedom.
These equations can be combined into one equation as linear combination of Klein-Gordon equation and supplementary conditions
$\partial^{2}h_{ab}+\alpha\Big{(}\partial_{a}G_{b}+\partial_{b}G_{a}\Big{)}=0\,.$ (10)
It must be gauge invariant and using gauge transformation of gauge fixing condition $\delta G_{a}=\partial^{2}\Omega_{a}$ we obtain $\alpha=-1$.
Consequently, equation of motion takes a form of Ricci curvature
$R_{ab}=\partial^{2}h_{ab}-\Big{(}\partial_{a}G_{b}+\partial_{b}G_{a}\Big{)}=\partial^{2}h_{ab}-\partial_{a}\partial^{c}h_{cb}-\partial_{b}\partial^{c}h_{ca}+\partial_{a}\partial_{b}h^{c}{}_{c}=0\,.$ (11)
Then scalar curvature is
$R=\eta^{a}R_{ab}=\partial^{2}h-2\partial^{a}G_{a}=2\Big{(}\partial^{2}h-\partial^{a}\partial^{b}h_{ab}\Big{)}\,,$ (12)
and Einstein tensor for free field equation
$G^{ab}(x)=R^{ab}-\frac{1}{2}\eta^{ab}R=\partial^{2}\Big{(}h_{ab}-\frac{1}{2}\eta^{ab}h\Big{)}-\Big{(}\partial_{a}G_{b}+\partial_{b}G_{a}\Big{)}+\eta^{ab}\partial^{c}G_{c}$
$=\partial_{c}\partial^{b}h^{ac}-\partial^{2}h^{ab}-\partial^{a}\partial^{b}h+\partial^{a}\partial_{c}h^{cb}-\eta^{ab}\Big{(}\partial_{c}\partial_{d}h^{cd}-\partial^{2}h\Big{)}=0\,.\qquad$ (13)
All these tensors can be obtained as small perturbation around flat metric using relation $g_{ab}=\eta_{ab}+h_{ab}$, see Refs.[11, 12].
Field equation (6.2) can be obtained from the action
$I_{0}(h^{ab})=\int d^{4}x\Big{(}\frac{1}{2}\partial_{c}h_{ab}\partial^{c}h^{ab}-\partial_{c}h_{ab}\partial^{b}h^{ac}+\partial_{c}h^{ac}\partial_{a}h-\frac{1}{2}\partial_{a}h\partial^{a}h\Big{)}\,.$ (14)
Similarly as in the case of vector field it can be shown (see for example Ref.[12]) that field $h^{ab}$ which satisfy this equation has two degrees of freedom. The symmetric tensor $h^{ab}$ has $10$ components. To fix gauge invariance (8) we can use $4$ gauge fixings conditions $G^{b}=\partial_{a}h^{ab}-\frac{1}{2}\partial^{b}h^{a}{}_{a}=0$ which kills $4$ component of tensor field and after that equation of motion takes the form $\partial^{2}\Big{(}h^{ab}-\frac{1}{2}\eta^{ab}h\Big{)}=0$. But now both gauge fixing and field equation are invariant under additional special gauge transformation $\delta h^{ab}(x)=\partial^{a}\Omega_{0}^{b}(x)+\partial^{b}\Omega_{0}^{a}(x)$, such that $\partial^{2}\Omega_{0}^{a}=0$. Corresponding gauge fixing kills additional $4$ degree of freedom and consequently gravitation field $h^{ab}(x)$ has two degrees of freedom.
Interaction with matter field $\Phi^{A}$ we can describe with the action quadratic in $h^{ab}$
$I_{int}(h^{ab})=\int d^{4}x{\cal L}(h^{ab},\Phi^{A})\,.$ (15)
Note that the integration measure is the flat Minkowski measure $d^{4}x$, as Lagrangian is already quadratic in $h^{ab}$.
From requirement that $I_{int}(h^{ab})$ is gauge invariant $I_{int}(h^{ab}+\delta h^{ab})=I_{int}(h^{ab})$ we have
$\int d^{4}x\delta h^{ab}\frac{\partial{\cal L}}{\partial h^{ab}}=\int d^{4}x\Big{(}\partial^{a}\Omega^{b}+\partial^{b}\Omega^{a}\Big{)}\Theta_{ab}=0\,.\qquad\Theta_{ab}=\frac{\partial{\cal L}}{\partial h^{ab}}$ (16)
Using partial integration and fact that $\Theta_{ab}$ is symmetric tensor we obtain that $\Theta_{ab}$ is the conserved energy-momentum tensor, $\partial_{a}\Theta^{ab}=0$.
The complete action is $I(h^{ab})=I_{0}(h^{ab})+I_{int}(h^{ab})$. Its variation with respect to $h^{ab}$ produces complete Einstein equations in weak field approximation
$G_{ab}(h^{ab})=\Theta_{ab}\,.$ (17)
7 Second rank tensor with helicities $\lambda=\pm 1$ produces Maxwell equations
Beside highes helicities $\lambda=\pm 2$, second rank tensor has also components with helicities $\lambda=\pm 1$. We expect that these components describe Maxwell equations in terms of field strength.
7.1 Second rank tensor with helicities $\lambda=\pm 1$
The components of second rank tensor with helicities $\lambda=\pm 1$ have a form
$T_{\pm 1}^{ab}(k)=(P_{\pm 1})^{ab}{}_{cd}T^{cd}(k)=\Big{(}(\pi_{0})^{a}{}_{c}(\pi_{\pm})^{b}{}_{d}+(\pi_{\pm})^{a}{}_{c}(\pi_{0})^{b}{}_{d}\Big{)}T^{cd}(k)\,.$ (1)
Since $\pi_{0}$ project to two dimensional space and $\pi_{+}$ and $\pi_{-}$ to one dimensional space, both components $T_{\pm 1}^{ab}(k)$ are four dimensional representations and they have four degrees of freedom.
To avoid influence of nontrivial gauge transformations of $(\pi_{0+})^{a}{}_{b}$ we will save only $(\pi_{0-})^{a}{}_{b}$ part (gauge invariant part of $(\pi_{0})^{a}{}_{b}$) which means that we will take $(\pi_{0})^{a}{}_{b}\to(\pi_{0-})^{a}{}_{b}$. With that choice we obtain
$T_{\pm 1(-)}^{ab}(k)=(P_{\pm 1(-)})^{ab}{}_{cd}T^{cd}(k)=\Big{(}(\pi_{0-})^{a}{}_{c}(\pi_{\pm})^{b}{}_{d}+(\pi_{\pm})^{a}{}_{c}(\pi_{0-})^{b}{}_{d}\Big{)}T^{cd}(k)\,.$ (2)
Note that now each component $T_{+1(-)}^{ab}(k)$ and $T_{-1(-)}^{ab}(k)$ has two degrees of freedom.
Using expressions for projectors we get
$T_{\pm 1(-)}^{ab}(k)=-\frac{1}{4}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{cd}(k)\Big{)}k^{a}\breve{p}_{\pm}^{b}-\frac{1}{4}\Big{(}(\breve{p}_{\mp})_{c}q_{d}T^{cd}(k)\Big{)}\breve{p}_{\pm}^{a}k^{b}\,.$ (3)
Since $k_{a}k^{a}=0$ and $k_{a}\breve{p}_{\pm}^{a}=0$ there are constraints $k_{a}T_{\pm 1(-)}^{ab}(k)=0$. The fields $T_{\pm 1(-)}^{ab}(k)$ with helicities $\lambda=\pm 1$ contain one basic vector $\breve{p}_{\pm}^{a}$ and consequently relation (8) is satisfy.
For symmetric tensor we obtain
$(T^{S}_{\pm 1(-)})^{(ab)}(k)=-\frac{1}{4}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{(cd)}(k)\Big{)}\Big{(}k^{a}\breve{p}_{\pm}^{b}+k^{b}\breve{p}_{\pm}^{a}\Big{)}\equiv\mathbf{T}_{\mp 2}^{S}\Big{(}k^{a}\breve{p}_{\pm}^{b}+k^{b}\breve{p}_{\pm}^{a}\Big{)}\,,\qquad$ (4)
or
$(T^{S}_{\pm 1(-)})^{(ab)}(k)=k^{a}\varphi_{\pm}^{b}(k)+k^{b}\varphi_{\pm}^{a}(k)\,,\qquad\varphi_{\pm}^{a}=\mathbf{T}_{\mp 2}^{S}\breve{p}_{\pm}^{a}\,,$ (5)
and for antisymmetric one
$(T^{A}_{\pm 1(-)})^{[ab]}(k)=-\frac{1}{4}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{[cd]}(k)\Big{)}\Big{(}k^{a}\breve{p}_{\pm}^{b}-k^{b}\breve{p}_{\pm}^{a}\Big{)}\equiv\mathbf{T}_{\mp 2}^{A}\Big{(}k^{a}\breve{p}_{\pm}^{b}-k^{b}\breve{p}_{\pm}^{a}\Big{)}\,,\qquad$ (6)
or
$(T^{A}_{\pm 1(-)})^{[ab]}(k)=k^{a}A_{\pm}^{b}(k)-k^{b}A_{\pm}^{a}(k)\,,\qquad A_{\pm}^{a}=\mathbf{T}_{\mp 2}^{A}\breve{p}_{\pm}^{a}\,.$ (7)
Here $\varphi_{\pm}^{a}$ and $A_{\pm}^{a}$ are new fields defined above. Note that both $\varphi_{\pm}^{a}$ and $A_{\pm}^{a}$ are proportional to $\breve{p}_{\pm}^{a}$, as well as fields $V_{\pm}^{a}$ in Eq.(4.3). According to scalar product of basic vectors $k_{a}\breve{p}_{\pm}^{a}=0$ we can conclude that $k_{a}\varphi_{\pm}^{a}=0$ and $k_{a}A_{\pm}^{a}=0$, which means that these solutions are in Lorentz gauge. We will see that these fields describe electromagnetic interaction.
Finally, each component $(T^{S})_{+1(-)}^{ab}(k)$, $(T^{S})_{-1(-)}^{ab}(k)$, $(T^{A})_{+1(-)}^{ab}(k)$ and $(T^{A})_{-1(-)}^{ab}(k)$ have one degree of freedom.
7.2 Gauge transformation for second rank tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Using gauge transformations for basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=k^{a}\varepsilon_{\pm}$ we can obtain gauge transformations for symmetric and antisymmetric parts of second rank tensor with helicities $\lambda=\pm 1$
$\delta(T^{S}_{\pm 1(-)})^{(ab)}(k)=k^{a}k^{b}\omega_{\pm}\,,\qquad\omega_{\pm}(k)=-\frac{1}{2}\Big{(}q_{c}(\breve{p}_{\mp})_{d}T^{(cd)}(k)\Big{)}\varepsilon_{\pm}\,,\qquad$
$\delta(T^{A}_{\pm 1(-)})^{[ab]}(k)=0\,.\qquad\qquad\qquad\qquad$ (8)
So, symmetric tensor is gauge dependent and antisymmetric one is gauge invariant.
7.2.1 Symmetric tensor with helicities $\lambda=\pm 1$
Note that under space inversion ${\cal P}(T^{S}_{\pm 1(-)})^{(ab)}(k)=(T^{S}_{\mp 1(-)})^{(ab)}(k)$, ${\cal P}\varphi^{a}_{\pm}(k)=\varphi^{a}_{\mp}(k)$ and ${\cal P}\omega_{\pm}(k)=\omega_{\mp}(k)$. So, according to (11) instead of two components $(T^{S}_{+1(-)})^{(ab)}(k)$ and $(T^{S}_{-1(-)})^{(ab)}(k)$ we will introduce one symmetric tensor
$\varphi^{ab}(k)=\alpha(T^{S}_{+1(-)})^{(ab)}(k)+\beta(T^{S}_{-1(-)})^{(ab)}(k)\,,\qquad$ (9)
and instead of two components $\varphi^{a}_{\pm}$ we will introduce one vector field
$\varphi^{a}(k)=\alpha\varphi^{a}_{+}(k)+\beta\varphi^{a}_{-}(k)\,,\qquad$ (10)
which are incomplete irreducible representation of Poincare group with space inversion. The representation of the field $\varphi^{ab}(k)$ as well as of $\varphi^{a}(k)$ is two dimensional and consequently they have two degrees of freedom. Since the parameters $\omega_{\pm}(k)$ are also connected by parity transformation we can introduce one parameter $\omega(k)=\alpha\omega_{+}(k)+\beta\omega_{-}(k)$.
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, we obtain expression for symmetric tensor $\varphi^{ab}(x)$ in terms of vector field $\varphi^{a}(x)$ and theirs gauge transformations. For real field, according to (14) and (9) we have
$\varphi^{ab}(x)=\int d^{4}p\Big{(}\varphi^{ab}(p)+\varphi^{ab}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}(T^{S}_{+1(-)})^{(ab)}(p)+(T^{S}_{+1(-)})^{(ab)}(-p)\Big{)}+\beta\Big{(}(T^{S}_{-1(-)})^{(ab)}(p)+(T^{S}_{-1(-)})^{(ab)}(-p)\Big{)}\Big{]}e^{-ipx}\,.\qquad$ (11)
Then with the help of (5) we obtain
$\varphi^{ab}(x)=\int d^{4}p\,\Big{\{}p^{a}\Big{[}\alpha\Big{(}\varphi_{+}^{b}(p)-\varphi_{+}^{b}(-p)\Big{)}+\beta\Big{(}\varphi_{-}^{b}(p)-\varphi_{-}^{b}(-p)\Big{)}\Big{]}$
$+p^{b}\Big{[}\alpha\Big{(}\varphi_{+}^{a}(p)-\varphi_{+}^{a}(-p)\Big{)}+\beta\Big{(}\varphi_{-}^{a}(p)-\varphi_{-}^{a}(-p)\Big{)}\Big{]}\Big{\}}e^{-ipx}\,,$ (12)
and using (17) and (19)
$\varphi^{ab}(x)=\partial^{a}\Big{(}\alpha\varphi_{+}^{b}(x)+\beta\varphi_{-}^{b}(x)\Big{)}+\partial^{b}\Big{(}\alpha\varphi_{+}^{a}(x)+\beta\varphi_{-}^{a}(x)\Big{)}=\partial^{a}\varphi^{b}(x)+\partial^{b}\varphi^{a}(x)\,.\qquad$ (13)
To obtain gauge transformation we use (7.2) and (7.2.1)
$\delta\varphi^{ab}(x)(x)=\int d^{4}p\,p^{a}p^{b}\Big{[}\alpha\Big{(}\omega_{+}(p)-\omega_{+}(-p)\Big{)}+\beta\Big{(}\omega_{-}(p)-\omega_{-}(-p)\Big{)}\Big{]}e^{-ipx}\,,$ (14)
and with the help of (18) and (19)
$\delta\varphi^{ab}(x)=\partial^{a}\partial^{b}\Big{(}\alpha\omega_{+}(x)+\beta\omega_{-}(x)\Big{)}=\partial^{a}\partial^{b}\omega(x)\,.\qquad$ (15)
So, gauge transformation of field $\varphi^{a}(x)$ takes the form
$\delta\varphi^{a}(x)=\partial^{a}\omega_{1}(x)\,,\qquad\omega_{1}(x)=\frac{1}{2}\omega(x)\,.$ (16)
The field $\varphi^{a}(x)$ behaves like electromagnetic field. It is vector field, it has the same helicity, the same gauge transformation and the same number of degrees of freedom. So, this field should not be considered as new field.
7.2.2 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Antisymmetric part is gauge invariant and consequently Lorentz invariant and we can use it as physical field. Instead of two components $(T^{A}_{+1(-)})^{[ab]}(k)$ and $(T^{A}_{-1(-)})^{[ab]}(k)$, according to (11) we will introduce one antisymmetric tensor
$F^{ab}(k)=\alpha(T^{A}_{+1(-)})^{[ab]}(k)+\beta(T^{A}_{-1(-)})^{[ab]}(k)\,,\qquad$ (17)
which is irreducible representation of Poincare group with space inversion. Similarly, from (7) we have
$F^{ab}(k)=k^{a}A^{b}(k)-k^{b}A^{a}(k)\,,\qquad A^{a}(k)=\alpha A_{+}^{a}(k)+\beta A_{-}^{a}(k)\,.$ (18)
Using (6) and scalar product of basic vectors we can conclude that there are two constraints on field $F^{ab}(k)$
$k_{a}F^{ab}(k)=0\,,\qquad\varepsilon_{abcd}k^{b}F^{cd}(k)=0\,.\qquad$ (19)
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, in the same way as before we get
$\partial_{a}F^{ab}(x)=0\,,\qquad\varepsilon_{abcd}\partial^{b}F^{cd}(x)=0\,,\qquad$ (20)
which are just vacuum Maxwell equations. As well as in the previous case the representation of the field $F^{ab}(x)$ is two dimensional and consequently it has two degrees of freedom. We could have expected that, since it describes Maxwell equations.
To express coupling of electromagnetic fields with matter (to a conserved current) we must solve second equation which produces
$F_{ab}(x)=\partial_{a}A_{b}(x)-\partial_{b}A_{a}(x)\,,\qquad$ (21)
which also follows from (18).
Note that $A_{a}(x)$ is gauge field since $F_{ab}(x)$ is invariant under transformation $\delta A_{a}(x)=\partial_{a}\Omega$. Then interaction with a matter can be express as in (33).
8 Fourth rank tensor with helicities $\lambda=\pm 2$ produces general relativity in weak field approximation
Since a symmetric second-rank tensor field with helicities $\lambda=\pm 1$ describes electromagnetic interactions through the field strength tensor, we naturally expect that a fourth-rank tensor field with helicities $\lambda=\pm 2$ and appropriate symmetry properties should describe gravitational interactions through the Riemann curvature tensor.
For the fourth-rank tensor case, the field index $A$ corresponds to a quadruple index $A\to(abcd)$ with $\Psi^{A}\to T^{abcd}$. In this section, we will maintain the compact notation using index $A$ for simplicity, reverting to the explicit index notation $abcd$ only when discussing specific symmetry properties.
8.1 Fourth rank tensor
We are going to solve equation
$\Big{(}S^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}\Big{)}T^{B}=0\,,$ (1)
where $T^{A}$ is fourth rank tensor, $T^{A}=T^{abcd}$. The problem comes down to that of vector fields. According to (2) in abbreviated notation spin operator takes the form
$(S_{12})^{A}{}_{B}=\Big{(}\rho\delta^{3}+\delta\rho\delta^{2}+\delta^{2}\rho\delta+\delta^{3}\rho\Big{)}^{A}{}_{B}\,.\qquad$ (2)
Using expressions $\rho=\pi_{+}-\pi_{-}$ and $\delta=\pi_{0}+\pi_{+}+\pi_{-}$ after multiplications we obtain
$(S_{12})^{A}{}_{B}=\sum_{i=-4}^{4}i\Big{(}P_{i}\Big{)}^{A}{}_{B}\,,\qquad$ (3)
and
$\delta^{A}_{B}=(\delta^{4})^{A}{}_{B}=\Big{(}(\pi_{0}+\pi_{+}+\pi_{-})^{4}\Big{)}^{A}{}_{B}=\sum_{i=-4}^{4}\Big{(}P_{i}\Big{)}^{A}{}_{B}\,,$ (4)
where the multiplication factor $i$ in (3) is consequence of combinatorics.
Here $(P_{i})^{A}{}_{B}$ is defined as the sum of all possible terms $\pi_{i_{1}}\pi_{i_{2}}\pi_{i_{3}}\pi_{i_{4}}$ such that $\sum_{m=1}^{4}i_{m}=i$ where we assign $\pi_{+}\equiv\pi_{1}$ and $\pi_{-}\equiv\pi_{-1}$. The sequence of multipliers is significant. For example $\pi_{+}\pi_{0}\pi_{+}\pi_{0}$ is different from $\pi_{0}\pi_{+}\pi_{+}\pi_{0}$. Since $\pi_{0},\pi_{+}$ and $\pi_{-}$ are projectors, it follows that all $P_{i}$ for $i=0,\pm 1,\pm 2,\pm 3,\pm 4$ are also projectors, satisfying the orthogonality condition $P_{i}P_{j}=\delta_{ij}P_{i}$.
To obtain spectrum of helicities we are going to solve consistency condition that characteristic polynomial vanishes, $\det(S_{12}-\lambda\sigma\delta)^{A}{}_{B}=0$. Using (3) and (4) we obtain expression for characteristic polynomial
$(S_{12})^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}=\sum_{i=-4}^{4}(i-\lambda\sigma)(P_{i})^{A}{}_{B}\,.\qquad$ (5)
Since $(P_{i})^{A}{}_{B}$ are projection operators, expression $S_{12}-\lambda\sigma\delta$ is already factored and we have
$\det(S_{12}-\lambda\sigma\delta)=\prod_{i=-4}^{4}(i-\lambda\sigma)^{d_{i}}\,\det P_{i}=0\,,\qquad$ (6)
where $d_{i}$ are dimensions of eigenspaces $\Xi_{i}$. In our case
$d_{4}=1\,,\qquad d_{3}=8\,,\qquad d_{2}=28\,,\qquad d_{1}=56\,,\qquad d_{0}=70\,,$ (7)
with general relation $d_{i}=d_{-i}$. Note that $2(1+8+28+56)+70=4^{4}$.
Then with the help of relation $\sigma^{2}=1$ we obtain
$\det(S_{12}-\lambda\sigma\delta)=(-\lambda)^{d_{0}}\prod_{i=1}^{4}(i^{2}-\lambda^{2})^{d_{i}}\,\prod_{i=-4}^{4}\det P_{i}=0\,.\qquad$ (8)
Using the fact that $\det P_{i}\neq 0$ we can conclude that
$(-\lambda)^{d_{0}}\prod_{i=1}^{4}(i^{2}-\lambda^{2})^{d_{i}}=0\,,\qquad$ (9)
which produces spectrum of helicities for four rank tensor
$\lambda_{0}=0\,,\qquad\lambda_{\pm i}=\pm i\,.\qquad i=(1,2,3,4)$ (10)
The eigenfunctions, which are our candidates for irreducible representations, for four rank massless tensors are
$T_{i}^{A}(k)=(P_{i})^{A}{}_{B}T^{B}(k)\,.\qquad(i=0,\pm 1,\pm 2,\pm 3,\pm 4)$ (11)
8.2 General relativity in terms of Riemann tensor
In this subsection we will find irreducible representations of fourth rank tensor with helicity $\lambda=\pm 2$. First we obtain irreducible representations of proper Poincare group, $T_{\pm 2(-)}^{abcd}(k)$ and then of wider group which combined Poincare group and parity, $R^{abcd}(k)$. The last one is linearized Riemann tensor. Finally, we obtain Einstein equations in vacuum.
8.2.1 Irreducible representations of proper Poincare group
The fourth rank tensor with helicities $\lambda=\pm 2$ has a form
$T_{\pm 2}^{abcd}(k)=(P_{\pm 2})^{abcd}{}_{efgh}T^{efgh}(k)$
$=\Big{[}\sum_{\Sigma_{1}}(\pi_{+})^{a}{}_{e}(\pi_{+})^{b}{}_{f}(\pi_{0})^{c}{}_{g}(\pi_{0})^{d}{}_{h}+\sum_{\Sigma_{2}}(\pi_{+})^{a}{}_{e}(\pi_{+})^{b}{}_{f}(\pi_{+})^{c}{}_{g}(\pi_{-})^{d}{}_{h}\Big{]}T^{efgh}(k)\,,\qquad$ (12)
where $\Sigma_{1}$ and $\Sigma_{2}$ means sum over all possible permutations of projectors.
As well as in the case of second rank tensor with helicities $\lambda=\pm 1$, to describe particles with helicities $\lambda=\pm 2$ with the help of fourth rank tensor, in every place we will substitute $\pi_{0}$ with its gauge invariant part $\pi_{0-}$. Then fourth rank tensor with helicities $\lambda=\pm 2$ has two terms: one with $2$ vectors $\breve{p}_{\pm}^{a}$ and $2$ vectors $k^{a}$ and the other with $3$ vectors $\breve{p}_{\pm}^{a}$ and one vector $\breve{p}_{\mp}^{a}$
$T_{\pm 2(-)}^{abcd}(k)=\sum_{\Sigma_{1}}\mathbf{T}_{\mp 4}^{\Sigma_{1}}\,\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}k^{c}k^{d}+\sum_{\Sigma_{2}}\mathbf{T}_{\mp 4}^{\Sigma_{2}}\,\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}\breve{p}_{\pm}^{c}\breve{p}_{\mp}^{d}\,,\qquad$ (13)
where $\Sigma_{1}$ and $\Sigma_{2}$ means sum over all possible permutations of basic vectors. The coefficients have a form
$\mathbf{T}_{\mp 4}^{\Sigma_{1}}=\frac{1}{2^{4}}(\breve{p}_{\mp})_{e}\,(\breve{p}_{\mp})_{f}\,q_{g}q_{h}\,T^{efgh}\,.\qquad\mathbf{T}_{\mp 4}^{\Sigma_{2}}=\frac{1}{2^{4}}(\breve{p}_{\mp})_{e}\,(\breve{p}_{\mp})_{f}\,(\breve{p}_{\mp})_{g}(\breve{p}_{\pm})_{h}\,T^{efgh}\,.\qquad$ (14)
If we count numbers of basic vectors $\breve{p}_{\pm}^{a}$ we can conclude that relation (8) $n_{\pm}+n_{\mp}=\lambda$ valid since ±$\pm 2+0=\pm 2$ and ±$\pm 3+(\mp 1)=\pm 2$.
We are interested in tensor antisymmetric in the first pair ($ab$) as well as in the second pair of indices ($cd$). Then only terms of the form
$\mathbf{T}_{\mp 4}^{\Sigma_{1}}\to\mathbf{T}_{\mp 4}^{R}=\frac{1}{2^{4}}\Big{(}q_{e}(\breve{p}_{\mp})_{f}-q_{f}(\breve{p}_{\mp})_{e}\Big{)}\,\Big{(}q_{g}(\breve{p}_{\mp})_{h}-q_{h}(\breve{p}_{\mp})_{g}\Big{)}\,\,T^{efgh}\,,\qquad$ (15)
survive and we obtain
$T_{\pm 2(-)}^{[ab][cd]}(k)=\mathbf{T}_{\mp 4}^{R}\,\Big{(}k^{a}\breve{p}_{\pm}^{b}-k^{b}\breve{p}_{\pm}^{a}\Big{)}\Big{(}k^{c}\breve{p}_{\pm}^{d}-k^{d}\breve{p}_{\pm}^{c}\Big{)}\,.\qquad$ (16)
This tensor is symmetric under exchange of the two pairs of indices ($ab$) and ($cd$). We can rewrite it in terms of second rank symmetric tensor $h_{\pm}^{ab}=\mathbf{T}_{\mp 4}^{R}\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}$ as
$T_{\pm 2(-)}^{[ab][cd]}(k)=k^{a}k^{c}h_{\pm}^{bd}-k^{a}k^{d}h_{\pm}^{bc}-k^{b}k^{c}h_{\pm}^{ad}+k^{b}k^{d}h_{\pm}^{ac}\,.\qquad$ (17)
It is easy to check symmetry properties of tensor $T_{\pm 2(-)}^{[ab][cd]}(k)$: cyclic permutation symmetry and Bianchi identity
$T_{\pm 2(-)}^{a[bcd]}=0\,,\qquad T_{\pm 2(-)}^{ab[cd}k^{e]}=0\,.$ (18)
With the help of expressions for scalar product of the basic vectors and (16) we obtain
$T_{\pm 2(-)}^{ac}(k)=\eta_{bd}T_{\pm 2(-)}^{abcd}(k)=0\,,\qquad T_{\pm 2(-)}(k)=\eta_{ab}T_{\pm 2(-)}^{ab}(k)=0\,.$ (19)
Using the gauge transformations of the basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}$ we have
$\delta T_{\pm 2(-)}^{[ab][cd]}(k)=0\,,$ (20)
which means that fields $T_{\pm 2(-)}^{[ab][cd]}(k)$ are gauge invariant and we can use them as a physical fields. They are irreducible representations of proper Poincare group.
8.2.2 Curvatures as irreducible representations of Poincare group with parity
As explained earlier, instead of two components $T_{+2(-)}^{[ab][cd]}(k)$ and $T_{-2(-)}^{[ab][cd]}(k)$ we will introduce tensor $R^{abcd}(k)$ which is irreducible representation of Poincare group with space inversion. With the help of (11) we have
$R^{abcd}(k)=\alpha T_{+2(-)}^{[ab][cd]}(k)+\beta T_{+2(-)}^{[ab][cd]}(k)=k^{a}k^{c}h^{bd}(k)-k^{a}k^{d}h^{bc}(k)-k^{b}k^{c}h^{ad}(k)+k^{b}k^{d}h^{ac}(k)\,,\qquad$
$h^{ab}(k)=\alpha h_{+}^{ab}(k)+\beta h_{-}^{ab}(k)=\alpha\mathbf{T}_{-4}^{R}(k)\breve{p}_{+}^{a}\breve{p}_{+}^{b}+\beta\mathbf{T}_{+4}^{R}(k)\breve{p}_{-}^{a}\breve{p}_{-}^{b}\,.\qquad\qquad\qquad$ (21)
The coordinate representation for real fields is
$R^{abcd}(x)=\int d^{4}p\Big{(}R^{abcd}(p)+R^{abcd}(-p)\Big{)}e^{-ipx}\,,\qquad$ (22)
and with the help of (18)
$R^{abcd}(x)=\partial^{a}\partial^{c}h^{bd}(x)-\partial^{a}\partial^{d}h^{bc}(x)-\partial^{b}\partial^{c}h^{ad}(x)+\partial^{b}\partial^{d}h^{ac}(x)\,.\qquad$ (23)
where according to (19)
$h^{ab}(x)=-\int d^{4}p\Big{[}h^{ab}(p)+h^{ab}(-p)\Big{]}e^{-ipx}\,.\qquad$ (24)
In the expression (23) we can recognize Riemann tensor see Ref.[11], and in the expression (24) metric tensor in weak field approximation.
Then we can calculate linearized Ricci tensor
$R^{bd}(x)=\eta_{ac}R^{abcd}(x)=\partial^{2}h^{bd}(x)-\partial_{a}\partial^{d}h^{ba}(x)-\partial^{b}\partial_{a}h^{ad}(x)+\partial^{b}\partial^{d}h(x)\,,\qquad$ (25)
linearized scalar curvature
$R(x)=\eta_{bd}R^{bd}=\partial^{2}h-\partial_{a}\partial_{b}h^{ba}-\partial_{a}\partial_{b}h^{ab}+\partial^{2}h=2\Big{(}\partial^{2}h-\partial_{a}\partial_{b}h^{ba}\Big{)}\,,\qquad$ (26)
and linearized Einstein tensor
$G^{bd}(x)=R^{bd}-\frac{1}{2}\eta^{bd}R=\partial^{2}h^{bd}-\partial_{a}\partial^{d}h^{ba}-\partial^{b}\partial_{a}h^{ad}+\partial^{b}\partial^{d}h-\eta^{bd}\Big{(}\partial^{2}h-\partial_{a}\partial_{b}h^{ba}\Big{)}\,.\qquad$ (27)
8.2.3 Einstein equations in vacuum
Using expressions (18), (20) and (8.2.2), and substituting $k^{a}\to p^{a}\to i\partial^{a}$ we can conclude that tensor $R^{abcd}(x)$ has standard symmetry properties of Riemann tensor: cyclic permutation symmetry and Bianchi identity
$R^{a[bcd]}(x)=0\,,\qquad R^{ab[cd}k^{e]}(x)=0\,,$ (28)
and that it is gauge invariant
$\delta R^{abcd}(x)=0\,.\qquad$ (29)
So, it can be used as physical field.
From (19) we obtain that Ricci and scalar curvatures vanish
$R^{ac}(x)=\eta_{bd}R^{abcd}(x)=0\,,\qquad R(x)=\eta_{ab}R^{ab}(x)=0\,,$ (30)
which produces Einstein equations in vacuum
$G^{ac}(x)=R^{ac}(x)-\frac{1}{2}\eta^{ac}R(x)=0\,.$ (31)
To express coupling of gravitational fields with matter (to a conserved energy-momentum tensor) we must solve Bianchi identity in terms of symmetric tensor $h^{ab}(x)$. The solution is Eq. (23) and then Einstein tensor takes the form (27). Consequently, interaction with matter can be express as in (15).
8.2.4 Dimensions of the representations and degrees of freedom
Let us count dimensions of these representations. As we have already shown dimension of each of representation $T_{\pm 2}^{abcd}(k)$ defined in Eq.(8.2.1) is $d_{2}=28$, where we used the fact that subspace corresponding to $\pi_{0}$ is two dimensional. The first term is $6\cdot 2\cdot 2=24$ dimensional where $6$ is number of combination of $4$ things taken $2$ at a time and $2\cdot 2$ is dimension of subspace corresponding to two operators $\pi_{0}$. The second term is $4$ dimensional so that $24+4=28$. After substitution two dimensional projector $\pi_{0}$ with one dimensional $\pi_{0-}$ we obtain that $T_{\pm 2(-)}^{abcd}(k)$, defined in Eq.(13) is $6\cdot 1\cdot 1+4=10$ dimensional representation, which could be expected for symmetric second ranktensor.
Using symmetry properties the second term in (13) disappears and we are left with $6$ dimensional representation, describing with Eq.(16). Since in that equation we are missing terms with $k^{a}k^{b}$ and $k^{c}k^{d}$, we are left with $4$ dimensional representation. In Eq.(16) both multipliers are antisymmetric. The other three possibilities (that both multipliers are symmetric and that one is symmetric and the other is antisymmetric) complements dimension of such kind of representations to $4$. It means that in our particular case each of representation $T_{+2(-)}^{[ab][cd]}(k)$ and $T_{-2(-)}^{[ab][cd]}(k)$ is one dimensional. Consequently, they are irreducible representations of proper Poincare group with one degree of freedom.
According to discussion above, Riemann tensor $R^{abcd}(k)$ which is irreducible representation of Poincare group with space inversion, has two degrees of freedom.
9 Massless tensors with arbitrary rank
We are going to solve most general equation for integer helicity
$(S^{A}{}_{B}-\lambda\sigma\delta^{A}_{B})\Psi^{B}=0\,,$ (1)
where $\Psi^{A}$ is tensor with rank $n$ so that $\Psi^{A}=T^{a_{1}a_{2}\cdots a_{n}}$. The problem comes down to that of vector fields.
9.1 Spin operator, delta function and projectors for arbitrary rank tensors in terms of elementary projectors
First we are going to express spin operator $S^{A}{}_{B}$ and delta function $\delta^{A}_{B}$ for fields with arbitrary helicity in terms of corresponding expressions for vector fields.
We will use abbreviated notation (we will omit the indices ). It means that formally we multiply first term with first term and second term with second term
$(A\ast B)(C\ast D)=(AC\ast BD)\,.$ (2)
In abbreviated notation we can write
$(S_{12})^{A}{}_{B}=\Big{(}\rho\delta^{n-1}+\delta\rho\delta^{n-2}+\cdots\delta^{n-1}\rho\Big{)}^{A}{}_{B}=\sum_{k=0}^{n-1}\Big{(}\delta^{k}\rho\delta^{n-k-1}\Big{)}^{A}{}_{B}\,.\qquad$ (3)
Using relations $\rho=\pi_{+}-\pi_{-}$ and $\delta=\pi_{0}+\pi_{+}+\pi_{-}$ we obtain
$(S_{12})^{A}{}_{B}=\sum_{k=0}^{n-1}\Big{(}(\pi_{0}+\pi_{+}+\pi_{-})^{k}\,(\pi_{+}-\pi_{-})\,(\pi_{0}+\pi_{+}+\pi_{-})^{n-k-1}\Big{)}^{A}{}_{B}\,,\qquad$ (4)
and after multiplications
$(S_{12})^{A}{}_{B}=\sum_{k=1}^{n}k\Big{(}P_{k}-P_{-k}\Big{)}^{A}{}_{B}\,.\qquad$ (5)
Delta function for tensors with arbitrary rank has the form
$\delta^{A}_{B}=\delta^{n}=(\pi_{0}+\pi_{+}+\pi_{-})^{n}=\sum_{k=1}^{n}\Big{(}P_{k}+P_{-k}\Big{)}+P_{0}\,.$ (6)
Here $(P_{k})^{A}{}_{B}$ is defined as the sum of all possible terms $\pi_{i_{1}}\pi_{i_{2}}\cdots\pi_{i_{n}}$ such that $\sum_{m=1}^{n}i_{m}=k$ where we assign $\pi_{+}\equiv\pi_{1}$ and $\pi_{-}\equiv\pi_{-1}$. Note that the sequence of multipliers is significant. Since $\pi_{0}$, $\pi_{+}$ and $\pi_{-}$ are projectors it follows that all expressions $P_{k}$ are also projectors, satisfying the orthogonality condition $P_{k}P_{q}=\delta_{kq}P_{k}$ where $(k,q=0,\pm 1,\pm 2,\cdots,\pm n)$. The multiplication factor $k$ in (5) is consequence of combinatorics.
We already had explicit expressions for projection operators in the simplest cases: for vectors $(P_{0})^{a}{}_{b}=(\pi_{0})^{a}{}_{b}\,,(P_{\pm 1})^{a}{}_{b}=(\pi_{\pm})^{a}{}_{b}$ and for second rank tensors in Eq. (3), which are particular cases of projectors for arbitrary rank tensors.
9.2 Solution of spin equation for arbitrary rank
To obtain spectrum of helicities we are going to solve consistency condition
$\det(S_{12}-\lambda\sigma)^{A}{}_{B}=0\,.\qquad$ (7)
Using (5) and (6) we obtain
$(S_{12})^{A}{}_{B}-\lambda\sigma\delta^{A}_{B}=\sum_{k=-n}^{n}(k-\lambda\sigma)(P_{k})^{A}{}_{B}\,.\qquad$ (8)
Since $(P_{k})^{A}{}_{B}$ are projection operators $S_{12}-\lambda\sigma$ is diagonalized and we have
$\det(S_{12}-\lambda\sigma)=\prod_{k=-n}^{n}(k-\lambda\sigma)^{d_{k}}\det P_{k}=0\,,\qquad$ (9)
where $d_{k}$ is dimension of subspace after projection with $P_{k}$.
With the help of relations $\sigma^{2}=1$ and $d_{k}=d_{-k}$ we obtain
$\det(S_{12}-\lambda\sigma\delta)=(-\lambda\sigma)^{d_{0}}\prod_{k=1}^{n}(k^{2}-\lambda^{2})^{d_{k}}\,\prod_{k=-n}^{n}\det P_{k}=0\,.\qquad$ (10)
Using the fact that $\det P_{k}\neq 0$ and $\sigma\neq 0$ we can conclude that
$(-\lambda)^{d_{0}}\prod_{k=1}^{n}(k^{2}-\lambda^{2})^{d_{k}}=0\,,\qquad$ (11)
which produces eigenvalues of operator $(S_{12})^{A}{}_{B}$ and consequently all possible helicities for $n$ rank tensor
$\lambda_{0}=0\,,\qquad\lambda_{k}=k\,,\qquad\lambda_{-k}=-k\,,\qquad k=(1,2,\cdots,n)\,.$ (12)
The eigenfunctions, which are our candidates for irreducible representations, for massless tensors with arbitrary rank are
$\Psi_{m}^{A}=(P_{m})^{A}{}_{B}\Psi^{B}\,.\qquad(m=0,\pm 1,\pm 2,\cdots,\pm n)$ (13)
9.3 Dimension of subspaces $d_{k}$
The subspace projected by $\pi_{0}$ has dimension $2$ while those projected by $\pi_{+}$ and $\pi_{-}$ each have dimension $1$. Since $d_{k}=d_{-k}$ we obtain following expressions for the dimensions $d_{k}$.
If $P_{k}$ contain $m$ operators $\pi_{0}$ then for even $m$ and $r=\frac{1}{2}(n+k)$, $s=\frac{m}{2}$
$d_{k}=\sum_{s=0}^{r-k}\frac{n!}{(r-s)!(r-k-s)!(2s)!}2^{2s}\,,$ (14)
and for odd $m$ and $r=\frac{1}{2}(n+k-1)$, $s=\frac{m-1}{2}$
$d_{k}=\sum_{s=0}^{r-k}\frac{n!}{(r-s)!(r-k-s)!(2s+1)!}2^{2s+1}\,.$ (15)
In particular we have
$d_{n}=1\,,\qquad d_{n-1}=2n\,,\qquad$
$d_{n-2}=n(2n-1)\,,\qquad d_{n-3}=n(n-1)\frac{4n-2}{3}\,.\qquad$ (16)
In next two sections we will consider the highest helicity which is always one dimensional.
10 Components with highest helicity
In previous sections, we demonstrated that our principle field equations successfully reproduce all known cases with lower helicities $\lambda=0,1,2$. We now turn to a comparison of our framework with established results for higher helicity fields $\lambda\geq 3$.
The general theory of free fields with arbitrary helicity $n$ was first developed by Fronsdal [7]. We will show that Fronsdal’s formulation emerges as a special case of our approach when restricted to the highest helicity component.
10.1 Gauge transformation for massless tensors of arbitrary rank
To find little group transformation for massless tensors of arbitrary rank we will use little group transformation for massless vector fields. With the help of (2) in abbreviated notation we have
$(\Pi_{1})^{A}{}_{B}=\sum_{k=0}^{n-1}\Big{(}\delta^{k}\hat{\Pi}_{1}\delta^{n-k-1}\Big{)}^{A}{}_{B}\,,\qquad(\Pi_{2})^{A}{}_{B}=\sum_{k=0}^{n-1}\Big{(}\delta^{k}\hat{\Pi}_{2}\delta^{n-k-1}\Big{)}^{A}{}_{B}\,,$ (1)
where $(\hat{\Pi}_{1})^{a}{}_{b}$ and $(\hat{\Pi}_{2})^{a}{}_{b}$ in the sum are expressions for vector fields and $\delta^{k}$ is product of $k$ corresponding $\delta$ functions.
Tensor fields with helicity $\lambda_{i}$ has a form
$T_{\pm i}^{a_{1}a_{2}\cdots a_{n}}(k)=(P_{\pm i})^{a_{1}a_{2}\cdots a_{n}}{}_{b_{1}b_{2}\cdots b_{n}}T^{b_{1}b_{2}\cdots b_{n}}(k)\,.\qquad(i=0,1,2,\cdots,n)$ (2)
With help of (27) we can obtain its gauge transformations
$\delta T_{\pm i}^{a_{1}a_{2}\cdots a_{n}}(\varepsilon_{1},\varepsilon_{2},k)=\sum_{k=0}^{n-1}\Big{(}\delta^{k}\hat{\Delta}\delta^{n-k-1}\Big{)}^{a_{1}a_{2}\cdots a_{n}}{}_{b_{1}b_{2}\cdots b_{n}}T_{\pm i}^{b_{1}b_{2}\cdots b_{n}}(k)\,,\qquad$ (3)
where $\hat{\Delta}^{a}{}_{b}=i\Big{(}\varepsilon_{1}\hat{\Pi}_{1}+\varepsilon_{2}\hat{\Pi}_{2}\Big{)}^{a}{}_{b}$ is expression for vector fields.
Using expressions for eigenfunctions (2) and for projection operators as well as for transformation of basic vector fields (26) and the fact that $\hat{\Delta}^{a}{}_{b}$ acts on vectors and not on scalars we can obtain gauge transformation of arbitrary tensor.
The general expression for highest helicity tensor (for example tensor of rank $n$ with helicities $\lambda_{\pm n}=\pm n$) has a form
$T_{\pm n}^{a_{1}a_{2}\cdots a_{n}}=(P_{\pm n})^{a_{1}a_{2}\cdots a_{n}}{}_{b_{1}b_{2}\cdots b_{n}}T^{b_{1}b_{2}\cdots b_{n}}=\mathbf{T}_{\mp n}\,\breve{p}_{\pm}^{a_{1}}\breve{p}_{\pm}^{a_{2}}\cdots\breve{p}_{\pm}^{a_{n}}\,,\qquad$ (4)
where
$\mathbf{T}_{\mp n}=\frac{(-1)^{n}}{2^{n}}(\breve{p}_{\mp})_{b_{1}}\,(\breve{p}_{\mp})_{b_{2}}\,\cdots\,(\breve{p}_{\mp})_{b_{n}}\,T^{b_{1}b_{2}\cdots b_{n}}\,.\qquad$ (5)
If $T^{a_{1}a_{2}\cdots a_{n}}$ is antisymmetric in any two indices then $\mathbf{T}_{\mp n}$ and consequently $T_{\pm n}^{a_{1}a_{2}\cdots a_{n}}$ go to zero. Therefore representation with highest helicity is totally symmetric.
We can understand this fact with the help of basic vectors. The basic vectors $\breve{p}_{\pm}^{a}$ carry helicities ±$\pm 1$ and both basic vectors $k^{a}$ and $q^{a}$ carry helicities 00. If $n_{\pm}$ are numbers of vectors $\breve{p}_{\pm}^{a}$ in the representation of field $\Psi^{A}$, then its helicity is
$\lambda=n_{+}-n_{-}\,.$ (6)
So, representation with highest helicity contains all basic vectors $\breve{p}_{+}^{a}$ or all basic vectors $\breve{p}_{-}^{a}$ and obviously is totally symmetric.
Its gauge transformation is
$\delta T_{\pm n}^{a_{1}a_{2}\cdots a_{n}}=\sum_{i=1}^{n}k^{a_{i}}\Omega_{\pm n}^{a_{1}a_{2}\cdots a_{i-1}a_{i+1}a\cdots_{n}}\,,\qquad$ (7)
where gauge parameter has the form
$\Omega_{\pm n}^{a_{1}a_{2}\cdots a_{i-1}a_{i+1}a\cdots_{n}}=\mathbf{T}_{\mp n}\,\varepsilon_{\pm}\breve{p}_{\pm}^{a_{1}}\breve{p}_{\pm}^{a_{2}}\cdots\breve{p}_{\pm}^{a_{i-1}}\breve{p}_{\pm}^{a_{i+1}}\cdots\breve{p}_{\pm}^{a_{n}}\,,\qquad$ (8)
10.2 Highest helicity tensor $h_{A}(k)$ symmetric under space inversion in the frame of standard momentum
From now on we will use the simplified notation. With capital $A$ we will denote set of all indices $A=\{a_{1}a_{2}\cdots a_{n}\}$. The set of all indices except $a_{i}$ we will denote as $A_{i}=\{a_{1}a_{2}\cdots a_{i-1}a_{i+1}\cdots a_{n}\}$ and similarly set of all indices except $a_{i}$ and $a_{j}$ we will denote as $A_{ij}$.
If the tensor is symmetric under space inversion then instead of two components $T^{A}_{+n}(k)$ and $T^{A}_{-n}(k)$ which are incomplete irreducible representation of proper Poincare group we will introduce one vector field $h^{A}(k)$ which is incomplete irreducible representation of wider group which includes space inversion. Then we obtain
$h^{A}(k)=\alpha T_{+n}^{A}(k)+\beta T_{-n}^{A}(k)\,,$ (9)
We also introduce one parameter $\Omega^{A_{i}}(k)=\alpha\Omega_{+n}^{A_{i}}(k)+\beta\Omega_{-n}^{A_{i}}(k)$ instead of two components $\Omega_{+n}^{A_{i}}(k)$ and $\Omega_{-n}^{A_{i}}(k)$ related by space inversion.
The field $h^{A}(k)$ is symmetric in all indices
$h_{a_{a}a_{2}\cdots a_{n}}=h_{(a_{a}a_{2}\cdots a_{n})}\,,$ (10)
and using $\eta^{a_{1}a_{2}}\breve{p}_{\pm}^{a_{1}}\breve{p}_{\pm}^{a_{2}}=0$ and $k_{a}\breve{p}_{\pm}^{a}=0$ we conclude that
$\eta^{ab}h_{abA_{ij}}(k)=0\,,\qquad k^{a}h_{aA_{i}}(k)=0\,.$ (11)
The gauge transformation of field $h_{A}(k)$ has the form
$\delta h_{A}(k)=\sum_{i=1}^{n}k_{a_{i}}\Omega_{A_{i}}(k)\,.\qquad$ (12)
With the help of expression $\eta^{a_{1}a_{2}}\breve{p}_{\pm}^{a_{1}}\breve{p}_{\pm}^{a_{2}}=0$ and using (8) we also have condition on parameter $\Omega$
$\eta^{ab}\Omega_{abA_{ijk}}(k)=0\,.\qquad(n\geq 3)$ (13)
11 Relations with previous work
In this section, we employ our framework to systematically recover fundamental results from established studies of highest-helicity fields.
11.1 The highest-helicity fields in coordinate representation necessarily satisfy the Fierz-Pauli constraints
Let us go from standard momentum frame to arbitrary frame and then to coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$. Then we will consider field $h^{A}(x)$
$h^{A}(x)=\int d^{4}p\Big{(}h^{A}(p)+h^{A}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}T^{A}_{+n}(p)+T^{A}_{+n}(-p)\Big{)}+\beta\Big{(}T^{A}_{-n}(p)+T^{A}_{-n}(-p)\Big{)}\Big{]}e^{-ipx}\,,$ (1)
and gauge parameters $\Omega^{A_{i}}(x)$ in coordinate representation
$\Omega^{A_{i}}(x)=\int d^{4}p\Big{(}\Omega^{A_{i}}(p)+\Omega_{A_{i}}(-p)\Big{)}e^{-ipx}$
$=\int d^{4}p\Big{[}\alpha\Big{(}\Omega^{A_{i}}_{+n}(p)+\Omega^{A_{i}}_{+n}(-p)\Big{)}+\beta\Big{(}\Omega^{A_{i}}_{-n}(p)+\Omega^{A_{i}}_{-n}(-p)\Big{)}\Big{]}e^{-ipx}\,.$ (2)
The fields satisfy massless Klein-Gordon equation and has following gauge transformation
$\partial^{2}h_{A}(x)=0\,,\qquad\delta h_{A}(x)=\sum_{i=1}^{n}\partial_{a_{i}}\Omega_{A_{i}}(x)\,,\qquad$ (3)
when gauge parameters according to (13) satisfy Fronsdal condition [7]
$\eta^{ab}\Omega_{abA_{ijk}}(x)=0\,.\qquad(n\geq 3)\,.$ (4)
Later we will see that this is necessary condition that equation of motion is gauge invariant.
As well as fields $T^{A}_{\pm n}$ the field $h_{A}$ is totally symmetric and satisfy the same supplementary conditions
$h_{a_{a}a_{2}\cdots a_{n}}(x)=h_{(a_{a}a_{2}\cdots a_{n})}(x)\,,\qquad\eta^{ab}h_{abA_{ij}}(x)=0\,,\qquad\partial^{a}h_{aA_{i}}(x)=0\,.\qquad$ (5)
These are exactly Fierz-Pauli constraints [5], which eliminate the unwanted components ($\lambda<n$) of the tensor field $h_{A}(x)$.
11.2 Fronsdal approach
In this subsection we will derive Fronsdal equations and Fronsdal action [7]. The corresponding field consist of two components with highest helicity, which are one dimensional. Consequently the field has two degrees of freedom. To satisfy that requirement we must impose a condition on the field to be double-traceless.
11.2.1 Derivation of Fronsdal equations
Consequently, we obtain equations for tensor field $h_{A}$ with helicity $\lambda=n$ in arbitrary frame
$\partial^{2}h_{A}(x)=0\,,\qquad\partial^{a}h_{aA_{i}}(x)=0\,,\qquad\eta^{ab}h_{abA_{ij}}(x)=0\,.\qquad$ (6)
This shows that field equations consists of massless Klein-Gordon equation plus supplementary conditions (gauge fixing). Supplementary conditions define the minimal set of auxiliary fields suitable for the description of the action found in Refs.[6] and [7].
Let us combine two conditions from (6) into one condition known as de Donder gauge
$G_{A_{i}}=\partial^{b}h_{bA_{i}}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}h^{b}{}_{bA_{ij}}=0\,.\qquad$ (7)
Using expression (12) we can find particular gauge transformations
$\delta h_{A}(k)=\sum_{i=1}^{n}\partial_{a_{i}}\Omega_{A_{i}}(k)\,,\qquad\delta h_{bA_{i}}=\partial_{b}\Omega_{A_{i}}+\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}\Omega_{bA_{ij}}\,,\qquad$
$\delta h^{b}{}_{bA_{ij}}=2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}=2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\,.\qquad$ (8)
Then gauge variation of the de Donder tensor is
$\delta G_{A_{i}}=\partial^{b}\Big{(}\partial_{b}\Omega_{A_{i}}+\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}\Omega_{bA_{ij}}\Big{)}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\partial_{a_{j}}\Big{(}2\partial^{b}\Omega_{bA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\Big{)}$
$=\partial^{2}\Omega_{A_{i}}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{j}}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\,.\qquad$ (9)
Equations (6) can be combined into one equation as linear combination of massless Klein-Gordon equation and supplementary conditions in terms of de Donder gauge
$\partial^{2}h_{A}+\sum_{i=1}^{n}a_{i}\partial_{a_{i}}G_{A_{i}}=0\,.\qquad$ (10)
We can find coefficients $a_{i}$, from requirements that this equation is gauge invariant which means that
$\partial^{2}\sum_{i=1}^{n}\partial_{a_{i}}\Omega_{A_{i}}+\sum_{i=1}^{n}a_{i}\partial_{a_{i}}\Big{(}\partial^{2}\Omega_{A_{i}}-\frac{1}{2}\sum_{j=1\,j\neq i}^{n}\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{j}}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}\Big{)}=0\,.\qquad$ (11)
This is satisfied for $a_{i}=-1$ and
$\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{i}}\partial_{a_{j}}\partial_{a_{k}}\Omega^{b}{}_{bA_{ijk}}=0\,.$ (12)
Consequently, in order to ensure that the Fronsdal equation is gauge invariant, we must require that the gauge parameter is traceless
$\Omega^{b}{}_{bA_{ijk}}=0\,.$ (13)
In our approach, this condition is satisfied, as we have shown in (4).
So, equation of motion takes the form of Ricci like tensor
$R_{A}(h)\equiv\partial^{2}h_{A}-\sum_{i=1}^{n}\partial_{a_{i}}G_{A_{i}}=\partial^{2}h_{A}-\sum_{i=1}^{n}\partial_{a_{i}}\partial^{b}h_{bA_{i}}+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\partial_{a_{i}}\partial_{a_{j}}h^{b}{}_{bA_{ij}}=0\,.\qquad$ (14)
11.2.2 Fronsdal action
Then scalar like tensor is
$R_{A_{ij}}=\eta^{a_{i}a_{j}}R_{A}=\partial^{2}h^{a}{}_{aA_{ij}}-2\partial^{a}G_{aA_{ij}}-\sum_{i=1}^{n}\partial_{a_{i}}G^{a}{}_{aA_{ijk}}$
$=2\partial^{2}h^{a}{}_{aA_{ij}}-2\partial^{a}\partial^{b}h_{abA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\partial^{b}h^{a}{}_{abA_{ijk}}\,,\qquad\qquad$ (15)
and Einstein like tensor
$G_{A}=R_{A}-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\eta_{a_{i}a_{j}}R_{A_{ij}}=\partial^{2}h_{A}-\sum_{i=1}^{n}\partial_{a_{i}}\partial^{b}h_{bA_{i}}+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\partial_{a_{i}}\partial_{a_{j}}h^{b}{}_{bA_{ij}}$
$-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\eta_{a_{i}a_{j}}\Big{(}2\partial^{2}h^{a}{}_{aA_{ij}}-2\partial^{a}\partial^{b}h_{abA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}\partial_{a_{k}}\partial^{b}h^{a}{}_{abA_{ijk}}\Big{)}\,.\qquad$ (16)
So, we can take the action in the form
$S_{n}(h)=\frac{1}{2}\int d^{4}xh^{A}G_{A}=\frac{1}{2}\int d^{4}x\Big{[}h^{A}\partial^{2}h_{A}-\sum_{i=1}^{n}h^{A}\partial_{a_{i}}\partial^{b}h_{bA_{i}}+\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}h^{A}\partial_{a_{i}}\partial_{a_{j}}h^{b}{}_{bA_{ij}}$
$-\frac{1}{2}\sum_{i=1}^{n}\sum_{j=1\,j\neq i}^{n}\eta_{a_{i}a_{j}}\Big{(}2h^{A}\partial^{2}h^{a}{}_{aA_{ij}}-2h^{A}\partial^{a}\partial^{b}h_{abA_{ij}}+\sum_{k=1\,k\neq i,j}^{n}h^{A}\partial_{a_{k}}\partial^{b}h^{a}{}_{abA_{ijk}}\Big{)}\Big{]}\,.\qquad\qquad$ (17)
Using the notation $h_{A_{ij}}\equiv\eta^{ab}h_{abA_{ij}}=h^{a}{}_{aA_{ij}}$, up to total derivatives we have
$S_{n}(h)=\frac{1}{2}\int d^{4}x\Big{[}-\partial^{a}h^{A}\partial_{a}h_{A}+n\partial_{a}h^{aA_{i}}\partial^{b}h_{bA_{i}}+n(n-1)\partial_{a}\partial_{b}h^{abA_{ij}}h^{c}{}_{cA_{ij}}$
$+\frac{n(n-1)}{2}\partial^{c}h_{b}{}^{bA_{ij}}\partial_{c}h^{a}{}_{aA_{ij}}+\frac{n(n-1)(n-2)}{4}\partial_{c}h_{b}{}^{bcA_{ijk}}\partial^{b}h^{a}{}_{abA_{ijk}}\Big{]}\,.\qquad$ (18)
in complete agreement with Ref.[7].
11.2.3 Propagating degrees of freedom of Fronsdal field
From general consideration we know that highest helicity representation is one dimensional and carry one degree of freedom. Then field in Fronsdal equation consists of two highest helicity components and must carry two degrees of freedom. We will assume that it is $k$-traceless tensor and we will find $k$ so that corresponding tensor carries two degrees of freedom. Note that the condition $n\geq 2k$ must be satisfied.
Completely symmetric $k$-traceless tensor with helicity $n$ has
$h^{k}_{A}={n+3\choose 3}-{n+3-2k\choose 3}\,,\qquad\qquad$ (19)
independent components. After sme calculation we obtain
$h^{k}_{A}=n^{2}k-2nk(k-2)-4k^{2}+\frac{k}{3}\Big{(}4k^{2}+11\Big{)}\,.\qquad\qquad$ (20)
Some of these components are related by the gauge condition $G_{A_{i}}=0$. Since de Donder tensor is traceless, $\eta^{ab}G_{abA_{ijk}}=0$, we should imposes
$\Omega^{\prime}_{A_{i}}={n+2\choose 3}-{n\choose 3}=n^{2}\,,\qquad(n\geq 3)\qquad$ (21)
conditions. Note that parameter $\Omega_{A_{i}}$ has one index less then field $h_{A}$. Consequently comparing with (19) in the first term we will take $n\to n-1$. Because parameter $\Omega_{A_{i}}$ is traceless in the second term we will take $k=1$ and $n\to n-1$.
In de Donder gauge, the Fronsdal equation becomes
$\partial^{2}h_{A}=0\,.\qquad$ (22)
Note that de Donder gauge does not fix the gauge completely. Both Fronsdal equation and de Donder gauge are invariant under residual gauge transformation given by traceless gauge parameters $\omega_{A_{i}}$ which obey $\partial^{2}\omega_{A_{i}}=0$. This residual gauge kills additional $\omega^{\prime}_{A_{i}}=n^{2}$ components.
Therefore, since complete number of degrees of freedom must be equal to two we have
$h^{k}_{A}-\Omega^{\prime}_{A_{i}}-\omega^{\prime}_{A_{i}}=n^{2}k-2nk(k-2)-4k^{2}+\frac{k}{3}\Big{(}4k^{2}+11\Big{)}-2n^{2}=2\,.\qquad\qquad$ (23)
This is cubic equation for $k$ with one solution $k=2$. There is one more solution with integers $k=3$ and $n=3$. But it does not satisfy condition $n\geq 2k$.
This means that double-traceless fields for arbitrary helicity in Fronsdal equation indeed propagates two degrees of freedom.
11.3 Relation with de Wit Freedman approach
Let us shorty comment de Wit and Freedman approach [8], based on a hierarchy of generalized Christoffel symbols. Their final result is Fronsdal’s action which we obtained in previous section. Here we are going to comment their intermediate result: the gauge invariant generalized curvature.
We are interested in tensors of the form
$T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)=\mathbf{T}_{\mp 2n}^{R}\,\prod_{i=1}^{n}\Big{(}k^{a_{i}}\breve{p}_{\pm}^{b_{i}}-k^{b_{i}}\breve{p}_{\pm}^{a_{i}}\Big{)}\,.\qquad$ (24)
Using transformations $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=\varepsilon_{\pm}k^{a}$ we can conclude that every term is gauge invariant $\delta\Big{(}k^{a_{i}}\breve{p}_{\pm}^{b_{i}}-k^{b_{i}}\breve{p}_{\pm}^{a_{i}}\Big{)}=0$. Consequently tensors $T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)$ are gauge invariant and we can use them as a physical fields. They are irreducible representations of proper Poincare group.
The pair exchange property
$T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)=(-1)^{n}T_{\pm 2n(-)}^{(b_{1}b_{2}\cdots b_{n})|(a_{1}a_{2}\cdots a_{n})}(k)\,,\qquad$ (25)
follows from the fact that expression $k^{a_{i}}\breve{p}_{\pm}^{b_{i}}-k^{b_{i}}\breve{p}_{\pm}^{a_{i}}$ change the sign after exchange $a_{i}$ with $b_{i}$.
Using the fact that
$k^{a}\Big{(}k^{b}\breve{p}_{\pm}^{c}-k^{c}\breve{p}_{\pm}^{b}\Big{)}+cyclic(a,b,c)=0\,,\qquad\breve{p}_{\pm}^{a}\Big{(}k^{b}\breve{p}_{\pm}^{c}-k^{c}\breve{p}_{\pm}^{b}\Big{)}+cyclic(a,b,c)=0\,,\qquad$ (26)
we can check symmetry properties: cyclic permutation symmetry
$T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)+cyclic(a_{i},b_{i},b_{j})=0\,,\qquad$ (27)
and Bianchi identity
$k^{a}T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)+cyclic(a,a_{i},b_{i})=0\,.\qquad$ (28)
With the help of expressions for scalar product of the basic vectors and (24) we obtain
$\eta_{a_{i}b_{i}}T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)=0\,.$ (29)
We can rewrite (24) in terms of $n$ rank symmetric tensor
$h_{\pm}^{c_{1}c_{2}\cdots c_{n}}(k)=\mathbf{T}_{\mp 2n}^{R}\,\breve{p}_{\pm}^{c_{1}}\breve{p}_{\pm}^{c_{2}}\cdots\breve{p}_{\pm}^{c_{n}}\,.\qquad$ (30)
as
$T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)=\sum_{m=0}^{n}(-1)^{n-m}k^{a_{1}}\cdots k^{a_{m}}k^{b_{m+1}}\cdots k^{b_{n}}h_{\pm}^{b_{1}\cdots b_{m}a_{m+1}\cdots a_{n}}\,.\qquad$ (31)
11.3.1 Generalized curvatures in coordinate representation
As explained earlier, with the help of (11) instead of two components $T_{+2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)$ and $T_{-2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)$ we will introduce tensor
$R^{a_{1}a_{2}\cdots a_{n},b_{1}b_{2}\cdots b_{n}}(k)=\alpha T_{+2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)+\beta T_{-2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)\,,\qquad\qquad\qquad$ (32)
and instead of two components $h_{+}^{a_{1}a_{2}\cdots a_{n}}(k)$ and $h_{-}^{a_{1}a_{2}\cdots a_{n}}(k)$ we will introduce tensor
$h^{a_{1}a_{2}\cdots a_{n}}(k)=\alpha h_{+}^{a_{1}a_{2}\cdots a_{n}}(k)+\beta h_{-}^{a_{1}a_{2}\cdots a_{n}}(k)\,.\qquad\qquad\qquad$ (33)
Both tensors are irreducible representations of Poincare group with space inversion.
The coordinate representations for real fields are generalized Riemann tensor see Ref.[8]
$R^{a_{1}a_{2}\cdots a_{n},b_{1}b_{2}\cdots b_{n}}(x)=\int d^{4}p\Big{(}R^{a_{1}a_{2}\cdots a_{n},b_{1}b_{2}\cdots b_{n}}(p)+R^{a_{1}a_{2}\cdots a_{n},b_{1}b_{2}\cdots b_{n}}(-p)\Big{)}e^{-ipx}\,.\qquad$ (34)
and generalized metric tensor
$h^{a_{1}a_{2}\cdots a_{n}}(x)=\int d^{4}p\Big{(}h^{a_{1}a_{2}\cdots a_{n}}(p)+h^{a_{1}a_{2}\cdots a_{n}}(-p)\Big{)}e^{-ipx}\,.\qquad$ (35)
Just like a tensors $T_{\pm 2n(-)}^{(a_{1}a_{2}\cdots a_{n})|(b_{1}b_{2}\cdots b_{n})}(k)$ generalized Riemann tensor $R^{a_{1}a_{2}\cdots a_{n},b_{1}b_{2}\cdots b_{n}}(x)$ is gauge invariant and satisfy cyclic permutation symmetry and Bianchi identity.
For $n=1$ we obtain Maxwell field strength
$R^{ab}=\partial^{a}h^{b}-\partial^{b}h^{a}\,,$ (36)
and for $n=2$ Riemann tensor in weak field approximation (23)
$R^{abcd}(x)=\partial^{a}\partial^{c}h^{bd}-\partial^{a}\partial^{d}h^{bc}-\partial^{b}\partial^{c}h^{ad}+\partial^{b}\partial^{d}h^{ac}\,.\qquad$ (37)
Note that tensor $R^{abcd}$, which denoted in Ref.[8] with prime, is symmetric under exchange both indices $ab$ with $cd$ and not under separately exchange $a$ with $c$ or $b$ with $d$. We can construct symmetric tensor in both $ac$ and $bd$ indices as linear combination of antisymmetric Riemann tensors in the same way as in equation (2.10) of Ref.[8]
$R_{S}^{abcd}=\frac{1}{2}\Big{(}R^{acbd}+R^{adbc}\Big{)}\,,\qquad$ (38)
where $R_{S}^{abcd}$ is symmetric tensor of Ref.[8] which in that article generalizes to higher spin.
The generalized Riemann tensor of Ref.[8] can be obtained after symmetrization of the tensor (34) with respect of all indices $a_{i}$ as well as with respect of all indices $b_{i}$.
12 Arbitrary rank tensor with helicity $\lambda=1$ produces Maxwell equations and with $\lambda=2$ Einstein equations
Arbitrary rank tensor with $n\geq 2$ components contains also helicities $\lambda=\pm 1$ and $\lambda=\pm 2$. We will show that these tensors describe Maxwell equations and Einstein equations in weak field approximation.
12.1 Arbitrary rank tensor with helicities $\lambda=\pm 1$
The components of $n$ rank tensor with helicities $\lambda=\pm 1$ have a form
$T_{\pm 1}^{A}(k)=(P_{\pm 1})^{A}{}_{B}T^{B}(k)=\sum_{i=1}^{n}(\pi_{\pm})^{a_{i}}{}_{b_{i}}(\pi_{0})^{A_{i}}{}_{B_{i}}T^{b_{i}B_{i}}(k)\,.$ (1)
Since $\pi_{0}$ project to two dimensional space and $\pi_{+}$ and $\pi_{-}$ to one dimensional space, both components $T_{\pm 1}^{A}(k)$ are $2^{n-1}n$ dimensional representations and they have $2^{n-1}n$ degrees of freedom.
To avoid influence of nontrivial gauge transformations of $(\pi_{0+})^{a}{}_{b}$ we will save only $(\pi_{0-})^{a}{}_{b}$ part (gauge invariant part of $(\pi_{0})^{a}{}_{b}$) which means that we will take $(\pi_{0})^{a}{}_{b}\to(\pi_{0-})^{a}{}_{b}$. With that choice we obtain
$T_{\pm 1(-)}^{A}(k)=(P_{\pm 1(-)})^{A}{}_{B}T^{B}(k)=\sum_{i=1}^{n}(\pi_{\pm})^{a_{i}}{}_{b_{i}}(\pi_{0-})^{A_{i}}{}_{B_{i}}T^{b_{i}B_{i}}(k)\,.$ (2)
Note that now each component $T_{+1(-)}^{A}(k)$ and $T_{-1(-)}^{A}(k)$ has $n$ degrees of freedom.
We will consider tensors symmetric in all indices. The other $n-1$ possibilities are antisymmetric expression in $a_{i}$ and one of $n-1$ indices from $A_{i}$. It means that each component $(T)_{+1(-)}^{A}(k)$ and $(T)_{-1(-)}^{A}(k)$ has one degree of freedom. From now on, we will assume that all tensors are totaly symmetric in all indices, but we will not denote symmetrization explicitly. For example $T_{\pm 1(-)}^{a_{i}A_{i}}$ means totaly symmetric in all indices $A$.
Using expressions for projectors we get
$T_{\pm 1(-)}^{a_{i}A_{i}}(k)=-\frac{1}{2^{n}}\Big{(}(\breve{p}_{\mp})_{b_{j}}q_{B_{j}}T^{b_{i}B_{i}}(k)\Big{)}\sum_{i=1}^{n}k^{A_{i}}\breve{p}_{\pm}^{a_{i}}\equiv\mathbf{T}_{\mp n}\sum_{i=1}^{n}k^{A_{i}}\breve{p}_{\pm}^{a_{i}}\,.$ (3)
The fields $T_{\pm 1(-)}^{A}(k)$ with helicities $\lambda=\pm 1$ contain one basic vector $\breve{p}_{\pm}^{a}$ and consequently relation (8) is satisfy.
Therefore, we can write totally symmetric tensor in the form
$(T_{\pm 1(-)})^{A}(k)=\sum_{i=1}^{n}k^{A_{i}}A_{\pm}^{a_{i}}(k)\,,\qquad A_{\pm}^{a}=\mathbf{T}_{\mp n}\breve{p}_{\pm}^{a}\,,$ (4)
Here $A_{\pm}^{a}$ is new field. Note that it is proportional to $\breve{p}_{\pm}^{a}$, as well as fields $V_{\pm}^{a}$ in Eq.(4.3). According to scalar product of basic vectors $k_{a}\breve{p}_{\pm}^{a}=0$ we can conclude that $k_{a}A_{\pm}^{a}=0$, which means that these solutions are in Lorentz gauge. We will see that these fields describe electromagnetic interaction.
12.2 Gauge transformation for arbitrary rank tensor with helicities $\lambda=\pm 1$ and Maxwell equations
Using gauge transformations for basic vectors $\delta k^{a}=0$ and $\delta\breve{p}_{\pm}^{a}=k^{a}\varepsilon_{\pm}$ we can obtain gauge transformations for totally symmetric arbitrary rank tensor with helicities $\lambda=\pm 1$
$\delta(T_{\pm 1(-)})^{A}(k)=k^{A}\omega_{\pm}\,,\qquad\omega_{\pm}(k)=n\mathbf{T}_{\mp n}\varepsilon_{\pm}\,.\qquad$ (5)
and for field $A_{\pm}^{a}$
$\delta A_{\pm}^{a}(k)=k^{a}\tilde{\omega}_{\pm}\,,\qquad\tilde{\omega}_{\pm}(k)=\mathbf{T}_{\mp n}\varepsilon_{\pm}\,.\qquad$ (6)
Since $k^{a}$ and $q^{a}$ are invariant under space inversion and ${\cal P}\breve{p}_{\pm}^{a}=-\breve{p}_{\mp}^{a}$ and ${\cal P}\varepsilon_{\pm}=-\varepsilon_{\mp}$ we have ${\cal P}(T_{\pm 1(-)})^{A}(k)=(T_{\mp 1(-)})^{A}(k)$, ${\cal P}A^{a}_{\pm}(k)=A^{a}_{\mp}(k)$ and ${\cal P}\omega_{\pm}(k)=\omega_{\mp}(k)$. So, according to (11) instead of two components $(T_{+1(-)})^{A}(k)$ and $(T_{-1(-)})^{A}(k)$ we will introduce one symmetric tensor
$T^{A}(k)=\alpha(T_{+1(-)})^{A}(k)+\beta(T_{-1(-)})^{A}(k)\,,\qquad$ (7)
and instead of two components $A^{a}_{\pm}$ we will introduce one vector field
$A^{a}(k)=\alpha A^{a}_{+}(k)+\beta A^{a}_{-}(k)\,,\qquad$ (8)
which are incomplete irreducible representation of Poincare group with space inversion. The representation of the field $T^{A}(k)$ as well as of $A^{a}(k)$ is two dimensional and consequently both fields have two degrees of freedom. Since the parameters $\omega_{\pm}(k)$ and ${\tilde{\omega}}_{\pm}(k)$ are also connected by parity transformation we can introduce one parameter $\omega(k)=\alpha\omega_{+}(k)+\beta\omega_{-}(k)$ and one parameter ${\tilde{\omega}}(k)=\alpha{\tilde{\omega}}_{+}(k)+\beta{\tilde{\omega}}_{-}(k)$
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, we obtain expression for totally symmetric tensor $T^{A}(x)$ in terms of vector field $A^{a}(x)$ and theirs gauge transformations. For real field, according to (7) and (8)
$T^{A}(x)=\sum_{i=1}^{n}\partial^{A_{i}}A^{a_{i}}(x)\,.\qquad$ (9)
To obtain gauge transformation we use (7.2) and (6) and obtain
$\delta T^{A}(x)=\partial^{A}\omega(x)\,.\qquad\delta A^{a}(x)=\partial^{a}\omega_{1}(x)\,,\qquad\omega_{1}(x)=\frac{1}{n}\omega(x)\,.$ (10)
The field $A^{a}(x)$ behaves like electromagnetic field. It is vector field, it has the same helicity, the same gauge transformation and the same number of degrees of freedom. So, this describes electromagnetic interaction.
12.2.1 Antisymmetric tensor with helicities $\lambda=\pm 1$ and Maxwell equations
We already mentioned that there are $n-1$ possibe antisymmetric expressions in $a_{i}$ and one of $n-1$ indices from $A_{i}$. It means that each component $(T^{An})_{+1(-)}^{A}(k)$ and $(T^{An})_{-1(-)}^{A}(k)$ has one degree of freedom. Using expressions for projectors we get
$(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)=-\frac{1}{2^{n}}\Big{(}(\breve{p}_{\mp})_{b_{k}}k_{b_{q}}q_{B_{kq}}T^{b_{k}b_{q}B_{kq}}(k)\Big{)}\sum_{i,j=1}^{n}k^{A_{ij}}\Big{(}k^{a_{i}}\breve{p}_{\pm}^{a_{j}}-k^{a_{j}}\breve{p}_{\pm}^{a_{i}}\Big{)}$
$\equiv\mathbf{T}_{\mp n}\sum_{i,j=1}^{n}k^{A_{ij}}\Big{(}k^{a_{i}}\breve{p}_{\pm}^{a_{j}}-k^{a_{j}}\breve{p}_{\pm}^{a_{i}}\Big{)}\,.\qquad$ (11)
Since $k_{a}k^{a}=0$ and $k_{a}\breve{p}_{\pm}^{a}=0$ there are constraints $k_{a_{i}}(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)=0$ and $k_{A_{ij}}T_{(}T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)=0$. The fields $(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)$ with helicities $\lambda=\pm 1$ contain one basic vector $\breve{p}_{\pm}^{a}$ and consequently relation (8) is satisfy.
Both $k^{a}$ and $\Big{(}k^{a_{i}}\breve{p}_{\pm}^{a_{j}}-k^{a_{j}}\breve{p}_{\pm}^{a_{i}}\Big{)}$ are gauge invariant. Consequently fields $(T^{An})_{\pm 1(-)}^{a_{i}a_{j}A_{ij}}(k)$ are Lorentz invariant and we can use them as physical fields. Instead of two components $(T^{An}_{+1(-)})^{A}(k)$ and $(T^{An}_{-1(-)})^{A}(k)$, according to (11) we will introduce one antisymmetric tensor
$F^{A}(k)=\alpha(T^{An}_{+1(-)})^{A}(k)+\beta(T^{An}_{-1(-)})^{A}(k)\,,\qquad$ (12)
which is irreducible representation of Poincare group with space inversion. Similarly, from (12.2.1) we have
$F^{A}(k)=k^{A_{ij}}F^{a_{i}a_{j}}\,,\qquad F^{a_{i}a_{j}}=k^{a_{i}}A^{a_{j}}(k)-k^{a_{j}}A^{a_{i}}(k)\,,\qquad$
$A^{a}(k)=\alpha A_{+}^{a}(k)+\beta A_{-}^{a}(k)\,.\qquad$ (13)
Using scalar product of basic vectors we can conclude that there are two constraints on field $F^{ab}(k)$
$k_{a}F^{ab}(k)=0\,,\qquad\varepsilon_{abcd}k^{b}F^{cd}(k)=0\,.\qquad$ (14)
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, in the same way as before we get
$\partial_{a}F^{ab}(x)=0\,,\qquad\varepsilon_{abcd}\partial^{b}F^{cd}(x)=0\,,\qquad$ (15)
which are just vacuum Maxwell equations. As well as in the previous case the representation of the field $F^{ab}(x)$ is two dimensional and consequently it has two degrees of freedom. We could have expected that, since it describes Maxwell equations.
To express coupling of electromagnetic fields with matter (to a conserved current) we must solve second equation which produces
$F_{ab}(x)=\partial_{a}A_{b}(x)-\partial_{b}A_{a}(x)\,,\qquad$ (16)
which also follows from (12.2.1).
Note that $A_{a}(x)$ is gauge field since $F_{ab}(x)$ is invariant under transformation $\delta A_{a}(x)=\partial_{a}\Omega$. Then interaction with a matter can be express as in (33).
12.3 Symmetric tensor with helicities $\lambda=\pm 2$ and free Einstein equations
Similarlly as in the previous subsections we can consider fields with two projectors $(\pi_{\pm})^{a}{}_{b}$ and $n-2$ projectors $(\pi_{0-})^{a}{}_{b}$ (gauge invariant part of $(\pi_{0})^{a}{}_{b}$).
Using expressions for projectors we obtain
$(T^{Sy})_{\pm 2(-)}^{a_{i}a_{j}A_{ij}}(k)=\frac{1}{2^{n}}\Big{(}(\breve{p}_{\mp})_{b_{k}}(\breve{p}_{\mp})_{b_{q}}q_{B_{kq}}T^{b_{k}b_{q}B_{kq}}(k)\Big{)}\sum_{i,j=1}^{n}k^{A_{ij}}\breve{p}_{\pm}^{a_{i}}\breve{p}_{\pm}^{a_{j}}\equiv\mathbf{T}_{\mp n}\sum_{i,j=1}^{n}k^{A_{ij}}\breve{p}_{\pm}^{a_{i}}\breve{p}_{\pm}^{a_{j}}\,.\qquad$ (17)
It is proportional to the tensor (1)
$T_{\pm n}^{ab}(k)=\frac{1}{4}(\breve{p}_{\mp})_{c}(\breve{p}_{\mp})_{d}T^{cd}(k)\breve{p}_{\pm}^{a}\breve{p}_{\pm}^{b}\,,\qquad$ (18)
so that
$(T^{Sy})_{\pm 2(-)}^{a_{i}a_{j}A_{ij}}(k)\sim\sum_{i,j=1}^{n}k^{A_{ij}}T_{\pm 2}^{a_{i}a_{j}}(k)\,.\qquad$ (19)
If we go to arbitrary frame and coordinate dependent fields, $k^{a}\to p^{a}\to i\partial^{a}$, in the same way as before we get
$H^{a_{i}a_{j}A_{ij}}(x)\sim\sum_{i,j=1}^{n}\partial^{A_{ij}}h^{a_{i}a_{j}}(x)\,.\qquad$ (20)
Therefore, corresponding field is $n-2$ derivative of the field $h^{ab}$ which, as we showed earlier, describe Einstein equations in weak field approximation.
13 Conclusion
Our fundamental assumption is that all field equations for an arbitrary symmetry group can be derived from the corresponding Casimir eigenvalue equations. In this work, we explicitly confirm this statement for the case of the massless Poincaré group.
Through our systematic approach, we have verified Weinberg’s observation [4] that Lorentz invariance is violated by terms taking the form of local gauge transformations. The principle field equations provide a simpler confirmation of Weinberg’s approach by working directly with fields rather than creation and annihilation operators. Indeed, these equations allow for a more straightforward derivation of the corresponding actions, which play a crucial role in quantization. The origin of Weinberg’s observation lies in the non-commutativity of the Pauli-Lubanski vector components. We have rigorously verified these results for tensor fields of arbitrary rank.
The introduction of the projectors is crucial because their diagonalization of the characteristic polynomial allows us to solve the corresponding determinants of arbitrary order and determine the spectrum of all helicities.
We defined the basis vectors such that the helicity $\lambda$ takes a simple form in terms of these vectors. Specifically, we have $\lambda=n_{+}-n_{-}$, where $n_{\pm}$ denotes the number of basis vectors $\breve{p}_{\pm}^{a}$. We emphasize that helicity plays a fundamental role in determining both the field equations for irreducible components of a given field $\Psi^{A}$ and the number of their physical degrees of freedom. Importantly, the same physical theory can be described using different field representations. For instance, as demonstrated in Eq.(8) all representations with helicities $\lambda=\pm 1$ (including vector fields, symmetric and antisymmetric second-rank tensors and arbitrary rank tensors ) exhibit one additional vector $\breve{p}^{a}_{\pm}$ compared to $\breve{p}^{a}_{\mp}$ and consistently describe electromagnetic interactions.
Similarly, representations with helicities $\lambda=\pm 2$ (such as symmetric second-rank tensors and fourth-rank tensors with appropriate symmetry properties) possess two additional vectors $\breve{p}^{a}_{\pm}$ over $\breve{p}^{a}_{\pm}$ and naturally describe gravitation interaction.
These results challenge the conventional view that only the highest helicity of a given tensor field is physically significant. Instead, we demonstrate that all helicity states are physically relevant. While different field representations with the same helicity can describe identical physics, formulating the field equations in terms of certain representations may prove non-trivial in practice.
While the primary contribution of this work is the demonstration that the principle field equations (8) generate all known free field theory equations, our results also represent a significant advancement beyond previous studies of highest helicity fields [5, 6, 7, 8]. In all existing literature, only free relativistic fields with the highest helicity, based to Fierz-Pauli constraints, have been studied. This represents a particular case within present approach, with totally symmetric tensor fields. Unlike previous studies limited to highest-helicity fields, this approach also covers lower-helicity not totally symmetric fields. These fields admit a clear physical interpretation, as detailed in Sections 7, 8, and 12.
A crucial logical progression emerges in our framework: Covariant helicity definition $\to$ little group invariance $\to$ local gauge transformations $\to$ equations of motion $\to$ actions for massless fields.
This sequence reveals that the requirement of Lorentz invariance necessarily leads to the existence of local gauge transformations for massless fields.